A Bibliography of Publications in ACM SIGPLAN Notices, 1970–1979

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

13 January 2024 Version 3.20

Title word cross-reference

#5 [299].

+ [982]. 0 ÷ 0 [195]. = [982]. \ [25].

1 [906, 895, 996, 731, 444]. 1.5 [111]. 10 [761, 190]. 11 [657, 670, 805, 645]. 1100 [507]. 1130 [96, 131, 198, 25]. 1130/system [96]. 15th [94]. 16th [94]. 16th/71 [94]. 1970 [1003]. 1971 [1102, 197, 190]. 1973 [1103, 1106, 1110]. 1974 [1105, 419]. 1975 [1109, 1111]. 1976 [1113, 1114]. 1977 [1116, 1117, 1118]. 1978 [1119]. 1979 [1120].

2 [989]. 2R [344, 345, 346, 343].

3 [92]. 3000 [350]. 360 [85, 56, 507, 49, 129]. 360/370 [702, 701]. 36O [30]. 370 [320]. 6 [556, 265, 228, 280]. 60 [933, 924, 358, 677, 363, 926]. 635 [128]. 6500 [98]. 6600 [44]. 6600/7600 [44]. 67 [156]. 68 [236, 46, 784, 776, 775, 798, 783, 1028, 765, 777, 791, 788, 786, 7, 1031, 425, 790, 793, 789, 787]. 68-R [776]. 68C [785]. 68’R [796, 792]. 71 [94]. 72 [1068]. 7600 [44]. 77 [761, 1045, 850]. 8 [595]. 80 [290, 1014]. 84 [982]. 900 [433].

A-TABLE [1071]. A. [447]. AADC [331]. ABACUS [660]. ABACUS/X [660]. Abstract [635, 889, 794, 94, 630, 622, 323, 976, 181, 639, 261, 580, 809, 619, 382, 846, 627]. Abstraction [1112, 758, 757, 373, 625, 1034, 629]. Abstractions [811, 871, 624, 521, 768]. Abstracts [295]. accelerate [31]. access [89, 1100]. accessing [976, 78]. Ackermann [86]. ACM [1102, 1103, 1106, 1110, 1117, 929, 933, 937, 939, 1115, 946, 1104, 924, 931, 1108, 944, 948, 1107]. action [247]. actions [755]. activity [1094]. ad [133]. ADA [1090, 1024, 1025]. Adapting [664]. addenda [1075]. additions [247, 367]. addressed [470]. adequacy [1077, 976, 326]. adic [87]. advanced [356, 1099]. AEPL [144]. aerospace [349, 347]. After [415]. Afterword [950]. again [459]. against [233]. ahead [1054]. AI [818]. aid [702]. aided [676]. Aiding [729]. AIDS [1099, 525, 704, 528]. airborne [656]. AL [241, 248]. Algebra [923]. Algebra/Assertions [923]. algebraic [170, 747, 900, 1004, 383, 846, 567]. ALGOL [1118, 786, 236, 46, 784, 795, 23, 734, 785, 1010, 924, 783, 1028, 765, 773, 772, 477, 779, 1031, 926, 780, 425, 793, 143, 787, 807, 781, 796, 142, 489, 771, 774, 792, 776, 775, 798, 358, 777, 791, 788, 677, 363, 7, 925, 790, 789, 206]. Algol-60 [677]. algorithm [369, 3, 192, 20, 841]. algorithmic [171, 771]. Algorithms [199, 381, 90, 255, 477, 175, 185, 42, 58, 604, 403, 811]. ALICE [385]. alignment [60]. alive [887]. allocation [1037, 1059]. allow [20]. along [395]. Alpha [556]. Alpha-6 [556]. Alphard [757]. alternative [152, 857]. Alternatives [16]. ambiguity [601]. AMBIT [216]. AMBIT/L [216]. American [925]. among [574]. AMORD [824]. AMTRAN [139, 136, 137, 138]. analogy [812]. analysing [443]. Analysis [552, 32, 499, 551, 678, 102, 680, 277, 464, 337]. Analyzer [482, 1098]. analyzing [417]. and/or [31]. Angeles [1109, 1119]. Announcing [911]. ANS [616, 617, 1003, 726]. ANSI [132, 64, 59, 133]. ANSI/X3J1 [59]. answering [700]. anthology [294]. any [450]. APL [94, 85, 197, 927, 1017, 124, 90, 97, 54, 56, 88, 66, 23, 57, 21, 48, 92, 96, 131, 198, 928, 247, 128, 98, 55, 95, 194, 272, 82, 47, 89, 127, 25, 86, 584, 49, 65, 129, 195, 22, 196, 435]. APL-FIDOS [89]. APL-FIOS [89]. APL-like [435]. APL/1130 [131, 198]. APL/360 [85, 56, 49, 129]. APL/6500 [98]. APLISP [196]. apostate [1061]. appendix [95]. Application [1065, 166, 1030, 651, 716, 140, 430]. Applications [399, 652, 1076, 799, 194, 661, 654, 409, 224, 169, 348, 129, 138]. Applicative [401, 375, 385, 745, 403, 390]. approach [39, 697, 523, 1038, 598, 720, 152, 110, 321, 539, 500, 985, 1083, 768, 884, 612]. approaches [565]. appropriate [205]. April [1103, 1109, 1114, 1104, 1115]. APT [929, 930]. architectural [397]. architecture [964, 353, 349, 836, 527]. architectures [814]. Arden [1107]. argument [870]. argument-parameter [870]. arguments [148]. Arithmetic [860, 866, 735, 40]. Array [498, 483, 496, 875, 57, 855, 969, 501, 876, 43, 45, 488]. array-oriented [57]. arrays [736, 591, 716, 673, 611]. ARS [471]. art [220]. Artificial [1116]. artists [679]. ARTSPEAK [679]. ASC [497]. Aspects [745, 1031, 310, 311, 35, 50, 8, 541, 112]. Aspen [861]. assembler [792, 835]. assembly [702, 74, 241, 248]. assert [703]. assertion [520]. assertions [170, 531, 855, 810, 923]. assessment [516]. assignment [872, 873]. Assistant [222]. associate [73]. associated [282]. association [870]. Associative [80, 81, 488, 393]. ATOL [1001]. atomic [755]. attaining [544]. attribute [253, 1037, 789, 1000]. August [1116, 1120]. authors [13]. automate [666]. Automated [519, 528, 819, 525, 648, 614, 557]. Automatic [139, 810, 1043, 230, 380, 693, 378, 804, 529, 384, 104, 813, 790, 769, 700]. automatically [930]. automating [189]. Automation [620]. axiomatic [171]. axiomatization [903].

B5500 [670]. B6700 [70]. Babel [140, 106]. backtracking [821, 229]. bad [476]. balanced [644]. BALM [439]. barrier [459]. base [897, 469, 263, 8, 623, 1033, 629, 641]. Based [429, 825, 185, 455, 837, 208, 818, 998]. Basel [8]. bases [624]. basic [101, 362, 696, 653, 932, 931, 438]. basis [806, 293]. be [713, 996]. Beach [1114]. beaver [406]. beens [774]. Beginner [225]. behavior [898]. Bell [514]. Berlin [786]. best [896]. better [698, 276, 817]. between [570, 1030, 181, 417]. Bibliography [1081, 983, 987, 885]. binary [602, 728, 637, 397]. bindings [364]. biochemical [666]. bit [634, 132]. bits [953]. Bliss [78]. Block [715, 69, 215, 1049, 599, 353]. block-structured [1049]. BLUE [959]. BNF [276]. body [1058]. Book [426]. books [967, 905, 916]. Boolean [897, 977, 979]. Bootstrapping [507, 250]. bottom [151]. bottom-up [151]. Bounce [4]. Bounce-and-skip [4]. boundary [181]. box [583]. Breaking [459]. Brian [419]. Bridges [707]. Brief [113, 940]. BUILD [697]. Builder [826]. building [115, 247, 278]. bulletin [19, 93, 134, 91]. Bureau [654]. Business [620, 376, 409]. by-name [721].

C [725]. CA [1105, 1109, 1119]. caching [401]. Calculating [188]. Calculus [180, 512, 179]. Call [852]. calling [815, 234]. calls [835]. CAMIL [971]. Can [713, 476, 276]. candidates [476]. capability [1013]. capsule [219]. Carnaugh [1018]. case [420, 522, 232, 703, 1073, 233, 668, 453, 610, 597, 699]. causes [551, 187]. CDC [44]. cellular [393]. certain [559]. CF [385]. CFD [490]. changing [509]. Chapin [966]. chapter [895]. Character [132, 791, 21, 48, 65]. Characteristic [28]. Characteristics [908, 34, 1020, 1021, 766, 224, 945, 966]. chart [1017, 425]. charts [1038, 857]. check [195]. checker [129]. checking [743, 807, 1036, 564]. checkout [807]. chemistry [456]. Chi [542]. Chi/OS [542]. CII [290]. CIMS [1046]. circuits [417]. City [1112]. class [866]. classes [639]. classification [985]. CLEOPATRA [269]. clerical [762]. Closing [970]. CLU [758]. CO [1120]. COBOL [933, 436, 898, 701, 646, 503, 706, 1066, 488, 585, 934, 99, 105]. Code [1041, 1055, 39, 896, 897, 1040, 825, 500, 1056, 1065, 1048, 430, 105, 108]. code-generator [1040]. coded [220]. coefficients [214]. Cognitive [582]. collection [816, 778]. COLLOQUE [197]. coloring [1007]. commands [512]. Comment [904, 611, 718, 966, 1101]. Comments [770, 587, 726, 474, 717, 719, 920]. commercial [669]. committee [133, 1003, 94, 133]. common [33, 263, 996, 880, 601]. Communication [395, 671, 320, 322]. community [47]. Company [419]. Comparative [464]. Compared [823]. Comparison [1057, 808, 1017, 23, 53, 883]. compatible [449]. Compilation [36, 63, 952, 1049, 610]. compile [14, 509]. compile-time [509]. Compiler [1120, 253, 1041, 274, 1048, 307, 1046, 290, 1044, 30, 437, 897, 35, 50, 523, 34, 265, 250, 1045, 660, 271, 1047, 163, 272, 182, 278, 481, 1051, 492, 40, 878, 1050, 829, 1014, 18, 787, 807, 430]. compiler-compiler [18]. compiler-compilers [35, 50]. Compilers [1111, 1058, 854, 41, 30, 77, 35, 50, 102, 500, 51, 1060, 140]. Compiling [1095, 969]. complement [38, 977]. Complementary [565]. Complete [761, 110]. complex [260, 337, 1015]. complexity [975, 888, 849, 1088, 292, 418, 459, 1101]. composability [448]. composition [913]. comprehensive [269]. computable [169]. computation [762, 1054]. computational [176]. computations [174]. Computer [543, 677, 76, 290, 863, 350, 697, 653, 1018, 472, 267, 366, 689, 249, 345, 648, 220, 135, 666, 964, 458, 448, 201, 43, 45, 347, 348, 730, 70, 129, 343, 552, 676, 111, 393, 693, 537, 292, 418]. computer-aided [676]. computers [854, 316, 484, 278, 655, 356, 111]. computing [565, 845, 286]. concept [46, 57, 137, 422, 17]. concepts [822, 714, 964, 107, 520]. concerning [790]. concerns [574]. concrete [381]. concurrency [303, 832, 406]. concurrent [310, 311, 227, 548, 550, 911, 547]. condition [100]. conditional [1032]. conditions [207]. Conference [1109, 1112, 1117, 1118, 1119, 92, 929, 933, 937, 939, 946, 924, 931, 944, 948]. configuration [160]. connections [570]. Connector [600, 1069]. conserving [866]. considered [502, 770, 765, 902, 860, 674, 586, 243]. consisting [835]. consolidation [239]. construct [865, 608]. Constructing [550, 833]. Construction [1120, 1047, 638, 1005, 327, 461]. Constructive [402]. constructs [832, 833, 1082, 869, 479]. content [815, 721]. content-directed [815]. context [3, 148, 988, 1052, 470]. context-addressed [470]. context-free [3, 988]. Continuous [387, 694]. contour [69, 180]. contribution [877]. Control [32, 688, 814, 1088, 479, 1101, 412, 652, 917, 872, 308, 487, 873, 223, 608, 226, 1010, 227, 1039, 367, 245, 20, 687, 401, 833, 661, 570, 362, 584, 706, 589, 607, 1085, 4, 651, 133, 159, 336, 237, 859, 456, 244, 339, 819, 576, 824]. controlling [192, 675]. controversy [855, 416]. convenient [96]. conventions [234]. convergence [31]. Conversational [160]. Conversion [107, 112]. Copying [392]. Cornell [1074]. Coroutines [958, 907, 803, 901, 920]. Corporate [665]. correct [562, 185, 555, 550]. correction [187, 553]. Correctness [635, 574, 182, 173, 309, 523, 175, 577, 656, 590, 674]. corrigenda [1075]. cost [992]. Count [888]. counting [881]. counts [980]. creating [675]. creation [101]. criteria [669, 1015]. criterion [590]. critical [719]. cross [792, 648]. cross-system [648]. CS [725]. CSP [590]. CSSA [822]. CSSL [694]. CT [612]. culture [981]. Current [114, 356]. CWIC [30]. CWIC/36O [30]. cycles [505]. Cyclomatic [888, 1008, 975, 849]. Cylinders [83].

D [225]. DAHLIA [281]. dangling [359]. Data [1112, 76, 624, 572, 155, 577, 60, 1034, 629, 68, 75, 132, 621, 1001, 1077, 381, 635, 871, 157, 631, 256, 748, 897, 889, 469, 1071, 580, 1081, 794, 411, 759, 517, 1010, 569, 636, 678, 373, 630, 625, 640, 321, 322, 622, 1004, 82, 454, 324, 204, 561, 809, 619, 81, 382, 521, 846, 409, 706, 623, 330, 768, 628, 79, 618, 70, 282, 1033, 287, 1100, 828, 731, 732, 627, 386, 468, 839, 470, 903, 641, 83, 78, 80]. data-flow [409]. data-type [1071]. database [637]. databases [408]. dataflow [388]. Datasaab [431]. datasets [519]. DBTG [631]. DBTG-data [631]. DDA [678]. debate [415]. Debugging [1099, 540, 588, 529, 1083]. DEC [657]. decidability [181]. decimal [836]. Decision [100, 176, 101, 1018, 109, 602, 110, 102, 104, 107, 111, 106, 103, 105, 108, 112]. decision-logic-table [104]. decisions [465].

Decomposition [1015]. defense [674, 863, 880]. defined [584, 639]. defining [153, 757]. Definition [1112, 173, 1076, 313, 376, 321, 728, 442, 136, 29, 184, 623, 335, 618, 284, 234, 154]. definitional [714]. definitions [171, 174, 1004, 604, 403]. delitescent [964]. Denniston [966]. Denver [1120]. Department [880, 863]. dependencies [890]. Derivation [171, 512, 819]. Derivatives [775]. deriving [417]. descent [1050]. Description [781, 1062, 170, 940, 334, 137, 337, 789, 468, 457, 142]. descriptions [819]. Design [1117, 696, 256, 733, 844, 260, 34, 347, 524, 676, 208, 818, 564, 486, 697, 525, 558, 315, 559, 1002, 994, 387, 242, 269, 270, 271, 515, 485, 374, 581, 783, 317, 694, 1025, 273, 323, 428, 396, 737, 1030, 508, 1060, 742, 1048, 645, 814, 739, 1073, 954, 956, 955, 957, 582, 578, 658, 575]. designed [314]. designers [867]. designing [465]. desirability [800]. desirable [625]. Detab [99]. Detab/65 [99]. Detap [105]. Detap/55 [105]. determinacy [512]. determination [749, 750]. deterministic [151, 885]. developing [566]. Development [761, 258, 657, 286, 973, 650, 1027, 938, 630, 1096, 538, 185, 454, 59, 332, 655, 926, 947, 925, 285, 941, 1094, 513, 118]. device [683, 689]. devices [684]. diagrammatic [1038]. diagrams [315, 215, 508]. dialect [338]. dialogue [283]. DIAM [637]. dictionaries [77]. DIDYMUS [776]. differences [1030]. digital [348, 691, 692]. dilemma [615]. Dimensional [225, 212, 995, 210, 213, 224, 691, 692, 209, 710, 217]. Direct [437, 351, 347]. directed [801, 815, 42, 58, 213, 1074]. directing [4]. directions [200, 627]. director [656]. disc [470]. discipline [1093, 760]. disciplined [725, 342]. disciplines [976]. discover [46]. discussion [320]. disease [249]. display [678, 682]. displaying [1013]. displays [685, 686]. distributed [401, 666, 335, 389]. distribution [553]. DLP [1002, 994]. DMAP [488]. do [276, 588, 718, 896]. document [463]. documentation [596]. documenting [100]. DOD [895, 959, 918, 960, 962, 961, 996]. DOD-1 [895, 996]. DoD-I [918]. does [1019, 27]. domain [623]. DOMINO [194]. dora [583]. down [614, 885]. Draft [616, 617, 726]. drawing [215]. dual [469]. dumping [807]. duration [915]. during [518]. Dynamic [441, 1037, 429, 435, 695, 749, 898, 855, 750, 767, 591, 675, 716, 673, 781, 611]. DYNOSOR [767]. DYSTAL [989].

eager [406]. Earley [185]. Early

[756, 945, 934]. ECL

[157, 163, 159, 338, 145]. ECMA [59]. ECMA/TC1O [59]. economics [210]. editing [856]. educational [657, 129]. effect [559, 753]. effective [588, 492, 461]. Effectiveness [40, 777]. effects [1058, 475, 881]. Efficient [1054, 743, 979, 78, 713, 603, 827, 830, 953, 884, 1048, 1073, 550]. effort [996, 915, 880]. electronic [296]. elegance [390]. elegant [269]. elements [419]. elimination [33]. Eliza [62]. ELMOL [296]. else [406, 668, 586]. embedded [424, 423]. Embedding [621, 735, 910]. Empirical [53]. emulator [428]. encapsulation [641, 340]. enforcement [748]. enforcing [855]. Engineering [729, 762, 533, 291]. English [391, 923]. enhance [271]. enhancements [64]. enough [583]. entry [109, 107, 108]. Environment [1113, 532, 46, 359, 919, 535, 1074]. Environmental [259]. environments [509]. EPS [12]. equation [28]. equational [1077]. equivalence [186, 1063, 73]. equivalent [417]. Errata [190, 191]. Error [1053, 807, 983, 270, 42, 58, 187, 848, 552]. error-handling [983]. error-prone [270]. Errors [56, 992, 551, 1052, 553, 999, 1075]. essential [766]. Estimating [206]. estimation [537]. ETC [161, 235]. ETOS [659]. Euclid [890, 893, 889, 894, 1035, 740, 742, 892, 891]. EULER [359]. European [926]. Evaluation [977, 251, 866, 467, 385, 759, 406, 557, 979]. evaluators [1037]. Evans [36]. evening [306, 305]. events [886]. Evolution [842, 518, 928, 332]. evolutionary [398]. example [580, 185]. examples [613, 365, 402, 956, 957, 1090]. Exception [581, 752, 751, 1012, 1013]. exceptions [837]. Excerpt [783]. exchange [721, 95]. EXEC [595]. Executing [407, 806]. execution [652, 540, 437, 345, 347, 355, 206, 606]. Executive [438]. exercisers [878]. exhibition [202]. exist [1019]. existing [704]. Expects [761]. Experience [854, 597, 708, 645, 525, 756, 325, 542, 709, 514]. Experiences [646, 655, 1064, 676]. experiment [723, 711, 712]. Experimental [438, 20, 759, 753, 418]. experiments [474]. explicit [1039, 824]. EXPLOR [220]. expressing [1010, 1005, 878]. Expression [38, 872, 866, 873]. expressions [251, 460, 897, 875, 866, 74, 212, 876, 977, 40, 240, 494, 979, 710]. Extended [635, 181, 647, 735, 1047, 677, 607, 1000, 108]. extended-entry [108]. extending [589]. Extensibility [156, 578, 723, 14, 17]. Extensible [168, 165, 336, 1023, 41, 15, 708, 261, 678, 8, 155, 144, 215, 7, 16, 987, 146, 368, 5, 166, 162, 140, 18, 167, 164]. extension [11, 762, 413, 88, 9, 127, 160, 953, 449, 849, 675, 488, 989, 389]. Extensions [952, 279, 838, 1076, 141, 708, 799, 445, 446, 747, 158, 153, 489, 491]. extensive [702]. EXUS [357].

fabrication [281]. facilitate [589, 739]. facilities [1001, 821, 625, 9, 159]. facility [11, 8]. failed [167]. family [331]. Fast [26, 355]. fatal [249]. Fault [565, 884, 563]. fault-intolerance [565]. Fault-tolerance [565]. FCPU [431]. FEA [923]. feature [672]. features [754, 144, 89, 10]. Feb [94]. feedback [399]. fewer [509, 999, 1075]. FFP [397]. FFP’S [392]. FIDOS [89]. File [247, 54, 466, 745, 89]. final [246]. Finding [531]. Finite [538, 1077, 985, 903]. FIOS [89]. FIRST [355]. fixes [56]. FL [1114]. flight [656]. flight-director [656]. Floating [126, 836]. floating-point [836]. FLOE [122, 123]. Florida [1115]. flourishing [981]. Flow [504, 32, 227, 324, 409, 228, 4, 1088, 1101]. flow-of-control [227]. Flowblocks [882]. Flowchart [298, 724, 1003]. Flowcharting [874, 505, 727]. flowcharts [857]. Floyd [36]. Floyd-Evans [36]. FOL [466]. FORAL [637]. form [592, 829]. Formal [185, 716, 284, 39, 540, 29, 334, 335, 923]. formalised [878]. FORMAT [856]. formatting [968, 856, 1064, 804, 868]. formulas [1030, 915]. formulation [501]. FORTH [963]. FORTRAN [695, 616, 132, 936, 647, 592, 23, 726, 594, 935, 480, 220, 834, 607, 738, 279, 481, 883, 455, 717, 989, 93, 134, 958, 808, 617, 761, 982, 870, 763, 865, 441, 707, 96, 708, 1045, 843, 766, 767, 500, 64, 449, 589, 790, 615, 739, 1014, 859, 850, 839, 490, 497]. Fortran-80 [1014]. FORTRAN-based [455]. FORTRAN-coded [220]. Fortran-Like [490]. Fortranner [726]. found [369]. foundations [810, 626]. four [915]. Fourier [26]. FP [411, 373, 377, 379]. FP-like [377]. Frame [252]. Free [592, 3, 588, 988]. Freiburghouse [904]. FROST [122, 123]. fuel [416]. Full [1003, 474]. fully [105]. fully-nested [105]. Function [371, 31, 194, 24, 86]. Functional [429, 408, 407, 399]. functions [174, 158, 584, 600, 169, 87]. functors [794]. fundamental [107]. Fundamentals [216]. Further [719, 920]. future [1100].

GA [1103]. Gaithersburg [1110, 1108]. game [511]. garbage [816, 778]. Gaussian [27]. general [621, 6, 218, 1005, 989, 440, 293, 938, 1098]. general-purpose [440, 293]. generalization [57]. Generalized [815, 79, 864, 749, 750, 74, 720, 24]. Generalizing [88]. generating [602, 105]. Generation [1041, 729, 896, 253, 648, 500, 296, 519, 1048, 99, 693, 430, 108]. generative [781, 293]. generator [1040, 150, 825, 1051, 1065, 724]. generators [757]. geographical [680]. Geometric [795]. Georgia [1104]. get [476]. Glasgow [1118]. Global [33, 1052, 243, 855, 363, 1060]. GLOSS [352, 319]. Goal [801]. going [168]. Gopalakrishna [1101]. GOTO [415, 416, 588, 232, 231, 600, 233, 1089]. GOTO-free [588]. GPL [6]. GPSS [937, 938]. graded [987]. graduated [1090]. grammar [781, 1000, 294]. grammars [253, 268, 995, 988, 1037, 988]. Grammatical [153]. Graphic [1114, 1115, 222, 684]. Graphical [212]. graphics [76, 683, 688, 690, 694, 218, 220, 685, 686, 676, 434, 693, 682, 679, 75, 677]. graphics-oriented [694]. GRAPHOS [213]. GREEN [960, 1012]. Grenoble [94]. grok [322]. growth [66]. GSA [1098]. Guarded [512]. guide [485]. Gypsy [741].

Habermann [447]. HAL [224]. Halstead [669]. handling [983, 581, 752, 751, 1012, 1013, 837]. hardware [773, 527]. hardware-software [527]. harmful [502, 770, 765, 902, 860, 674, 586, 243, 1089]. Harriman [1106, 1107]. HARVEY [258]. Hashed [1039]. heterogeneous [458]. Hierarchical [313, 1073, 862, 360, 640, 306, 549, 840]. hierarchy [119, 244]. High [1105, 344, 378, 1027, 268, 623, 863, 635, 351, 437, 256, 733, 844, 200, 909, 266, 242, 636, 391, 1029, 352, 640, 273, 827, 661, 996, 680, 609, 396, 119, 826, 675, 836, 355, 356, 663, 440]. High-level [344, 1027, 351, 437, 256, 200, 640, 827, 661, 680, 609, 826, 836, 440]. High-speed [268]. higher [147, 429, 230, 676, 71, 880]. higher-level [147, 676]. highlights [446]. highly [725]. highly-disciplined [725]. Hill [419]. historical [1027]. History [1119, 936, 943, 929, 933, 937, 939, 895, 946, 924, 949, 931, 945, 944, 934, 948]. hobbits [66]. hoc [133]. HOL [347]. Honeywell [128]. horizontally [432]. House [1107, 247]. HP [350]. hypergeometric [24].

I/O [85, 1093]. IBM [701, 507]. IBM/360 [507]. ICL1900 [792]. Icon [1009]. ID [389]. ideal [746]. ideas [636]. identification [389]. identifiers [1011]. If [668, 406, 586]. If-then-else [406, 586]. Ignorance [765]. II [365, 936, 95, 637]. III [936, 967]. ILLIAC [481, 490]. illustrations [779]. IMP [9]. IMP72 [413]. impact [515]. impacts [996]. Implanting [397]. Implementation [290, 792, 556, 1045, 1042, 353, 169, 662, 823, 1102, 741, 906, 46, 157, 467, 387, 743, 316, 788, 321, 194, 273, 786, 135, 207, 595, 396, 1034, 455, 780, 229, 858, 645, 1087, 113, 1073, 676, 208, 838, 114, 817, 143, 991, 202, 67]. implementations [369, 827, 401, 879]. implemented [111]. implementing [30, 466, 632, 1035, 1049, 116]. implicit [494]. importance [46]. impossibility [721]. imposter [359]. improvement [877]. Improving [412]. Impurities [369]. inconsistencies [1087]. increasing [707]. incremental [816, 660, 160]. indention [1091]. Independent [1041, 1055, 647, 683, 689, 722, 705, 1056, 835, 1059, 782]. Indeterminate [195]. Index [894]. indicate [869]. induce [31]. Inductive [172]. industrial [1094]. inference [153]. Inferred [125]. influence [554]. infological [637]. Information [462, 1110, 631, 466, 467, 458, 465, 73, 332, 473, 1108, 474, 858, 461, 93, 134]. inhibitor [406]. input [993, 417, 850, 684, 1070]. input-output [850]. input/output [417]. installment [365]. instances [1037, 778]. instantiations [811]. Instruction [349, 439, 953]. Instruments [497]. integers [986]. Integrated [317, 649, 254, 1033]. integration [199, 163, 27, 665]. Intel [1014]. Intelligence [1116]. Intent [355]. interaction [754]. Interactive [993, 568, 1099, 845, 444, 1093, 267, 690, 678, 376, 318, 471, 135, 458, 464, 465, 286, 222, 676, 193, 1070, 612, 217, 75, 1042, 506]. Interface [1103, 1113, 1104, 1108, 1107, 1106, 1110, 463, 161, 235, 458, 427, 217]. interim [365]. Intermediate [1055, 1057, 437, 805, 970]. Internal [834, 354]. International [1115, 1109]. interpretation [404, 720]. Interpreter [163, 162]. Interpreter/compiler [163]. interpreters [780]. Interpretive [652, 690]. interpretive/interactive [690]. interprocess [320]. interrupts [976]. Intgrl [27]. intolerance [565]. Introduction [72, 410, 5, 346, 1090, 400]. introductory [249, 883]. intuitive [667]. Invariant [255, 748, 810, 492, 531]. inversion [194]. investigation [881, 753]. invocation [815, 315]. involving [227]. IRIS [290]. IRONMAN [863, 954, 956, 955, 957, 1012]. ISMs [312]. Isolation [890]. issue [139, 190, 191]. Issues [575, 581]. iteration [757]. iterative [604, 699]. ITRAN [279]. IV [481, 490]. Ivtran [482].

J [365, 419]. January [761]. Job [245]. JOSS [939]. JOVIAL [940, 941]. July [863]. June [1106, 1119, 1107].

KAIL [608]. Kernel [332, 545]. Kernighan [419]. key [422, 634]. keyword [671]. keywords [970, 922]. know [555].

Knowledge [522, 825]. knowledge-based [825].

L [903, 216]. label [359]. labels [246]. laboratory [189, 666, 658]. Labs [514]. Lafayette [1102]. Lake [1112]. LALR [1054, 1080, 848]. Lambda [180, 179]. Lament [726]. lamentably [887]. Language [1117, 754, 118, 1055, 271, 445, 821, 465, 1100, 954, 956, 490, 823, 808, 621, 741, 344, 295, 927, 929, 933, 937, 872, 173, 1001, 939, 1062, 487, 702, 486, 973, 413, 252, 351, 313, 122, 437, 698, 256, 141, 168, 733, 844, 258, 314, 405, 466, 873, 260, 719, 263, 1019, 223, 959, 960, 962, 961, 1040, 123, 250, 1002, 994, 242, 74, 205, 946, 269, 316, 366, 515, 6, 485, 446, 1010, 374, 1097, 924, 1009, 783, 678, 241, 248, 439, 358, 1096, 391, 746, 365, 820, 640, 321, 1025, 273, 394, 622, 8, 20]. language [126, 827, 805, 687, 144, 714, 661, 996, 324, 680, 740, 431, 447, 421, 275, 428, 277, 396, 931, 327, 737, 59, 1030, 7, 409, 296, 224, 331, 203, 280, 978, 1013, 146, 455, 433, 780, 675, 971, 1033, 858, 835, 137, 286, 867, 346, 869, 944, 930, 724, 645, 837, 1087, 116, 222, 230, 676, 355, 356, 637, 948, 13, 955, 957, 691, 582, 1068, 357, 289, 1094, 10, 167, 578, 663, 457, 840, 818, 998, 440, 462, 692, 679, 575, 981, 641, 880, 861, 78, 819, 516, 216, 711, 164, 712, 294, 117, 130, 400]. language [620, 771, 822, 120, 935, 662, 427, 1023]. language-oriented [439]. language-related [295]. Languages [1103, 1105, 1111, 1114, 1116, 1119, 1057, 2, 1102, 1106, 1110, 171, 744, 1, 863, 929, 933, 937, 862, 800, 308, 939, 360, 697, 1115, 635, 1063, 762, 147, 15, 259, 467, 385, 1020, 1026, 200, 909, 165, 266, 387, 226, 1080, 3, 1027, 946, 304, 270, 643, 1104, 1039, 636, 924, 949, 245, 1021, 671, 1029, 694, 319, 820, 984, 377, 995, 155, 401, 609, 1049, 372, 29, 931, 119, 16, 329, 353, 987, 628, 1108, 1012, 883, 947, 5, 133, 284, 229, 579, 432, 286, 836, 944, 429, 166, 113, 452, 972, 948, 114, 18, 693, 682, 390]. languages [71, 68, 183, 479, 209, 1107, 339, 293, 202, 576, 435]. languages-information [1110]. languages-Microprogramming [1106]. Languages-Operating [1103]. Large [514, 525, 526, 476, 392, 557, 664]. Large-scale [514, 525]. last [926]. laws [192]. layout [1011]. learner [793]. learner-programmers [793]. Least [992]. Least-cost [992]. legal [474]. Level [1105, 344, 371, 635, 351, 147, 437, 256, 733, 844, 200, 909, 266, 378, 242, 1027, 636, 391, 352, 640, 321, 562, 273, 827, 661, 680, 609, 396, 826, 623, 675, 836, 429, 230, 676, 356, 637, 663, 440, 71]. Lexical [499]. library [463]. lifetime [401]. Like [490, 820, 377, 500, 435]. limited [109, 65, 107]. limited-character [65]. limited-entry [109, 107]. Limits [846, 900]. line [463, 129]. Linear [724]. lines [1069]. lingual [659]. linguistic [340]. linking [61]. Linus [711, 712]. LIS [321]. Lisp [1072, 19, 1071, 1019, 207, 943, 942, 182, 111, 823]. list [797, 196, 375, 216, 595]. list-processing [216]. literary [277, 396]. Little [662]. liveness [997, 1006]. LL [779]. Local [37]. logic [897, 404, 1010, 104, 169, 129]. logical [798, 100, 343, 567]. Logo [685, 686]. longer [450]. look [1063, 1054, 453]. look-ahead [1054]. Lookahead [848]. loop [865, 1079, 1083, 842]. loops [896, 1084, 492, 494, 505]. LR [52, 268, 1053, 53, 779]. LRLTRAN [122, 123, 491].

M [1010]. M.O.L. [597]. Macaroni [817]. Machine [1041, 436, 1056, 1059, 890, 360, 647, 124, 96, 385, 499, 3, 245, 352, 614, 428, 705, 283, 395, 835, 357, 440, 354, 130, 782, 438, 180]. Machine-Independent [1041, 1056, 1059, 647, 782]. machine-oriented [3]. Machines [1111, 486, 261, 720, 323, 977, 432, 664]. Macro [12, 152, 213, 987, 768, 835, 149, 1097]. macroprocessing [1066]. Macros [147, 702, 115, 148]. Madcap [223, 79, 228, 280]. made [465]. main [144]. Maintenance [543]. Majster [1022]. make [820]. making [899]. man [283, 534]. man-machine [283]. Man-Month [534]. management [463, 784, 897, 388, 321, 435]. Managing [513, 653]. manipulation [831, 640, 82, 830, 618, 781]. Manual [1024, 942, 954, 955]. many [537]. many-state [537]. map [1007]. mapping [680, 60]. maps [1018, 634]. March [1105, 1111, 1113, 1112, 1117, 1118, 761]. Mark [97]. marking [811]. Markov [536, 537]. MARY [237, 240, 246]. Maryland [1108]. MASK [856]. master [476]. matching [909, 791, 60, 851]. Mathematical [139, 478, 371, 587, 212, 655, 618, 710, 264, 810]. Mathematician [222]. mathematics [129]. matrices [651, 1050]. matrix [28, 910, 194]. maxicomputer [650]. May [1106, 1107]. McGraw [419]. McGraw-Hill [419]. MD [1110]. means [1090]. measure [975, 414, 849]. Measurement [888, 510, 769, 164]. measurements [881]. measures [705]. Measuring [669]. mechanism [721]. mechanisms [367, 758, 154]. medium [95]. Meeting [1103, 1113, 761, 1104, 1108, 1107, 1106, 1110, 94]. meets [47]. MEMBERS [438]. memories [81]. Memory [599, 429, 852, 914, 767, 343, 435]. Memory-Based [429]. MENELAUS [164]. merits [769]. Mesa [756]. Messages [984]. Meta [285, 965]. Meta-optimizer [285]. META/REDUCE [965]. METAPI [141]. method [855, 881, 588, 900, 158, 61, 337]. methodology [822, 448, 283, 418, 530]. Methods [682, 173, 1077, 306, 172, 201]. metric [520]. Miami [1114]. Michigan [98]. micro [648, 942, 651, 430, 1023]. micro-code [430]. micro-manual [942]. micro-programming [651]. micro/mini [648]. microcomputers [963]. microprocessor [649, 661]. Microprocessors [429, 1014]. microprogrammable [124]. microprogrammed [437, 432, 434, 438]. Microprogramming [1106, 431, 433, 658, 1107, 427]. might [774]. might-have-beens [774]. mini [854, 366, 648, 29, 666]. mini-computer [366, 666]. mini-computers [854]. mini-languages [29]. Minicomputer [650, 643, 647, 660, 646, 656]. minicomputers [275, 654, 663, 440]. minimally [270]. minimum [602]. MIRFAC [211]. missing [236, 25]. MLP [433]. MLP-900 [433]. mode [236, 687, 1031]. Model [826, 180, 360, 522, 318, 319, 1035, 69, 536, 443, 618, 146, 70, 537, 55, 622]. modeling [124, 218]. Modelling [71, 795, 538]. models [46, 186, 359, 267, 73, 176, 188, 68]. Modern [442, 1020, 1045, 1021]. modest [489]. Modification [198, 681]. Modifications [96, 131, 619]. Modula [893]. modular [559, 318]. module [672]. modules [1035]. molecular [217]. Monica [1105]. monitoring [204]. Month [534]. most [187, 246]. Motivation [15]. Movement [494]. moving [646]. MU5 [498]. multi [385, 1016, 1047, 1013, 1037, 87, 659, 796]. multi-adic [87]. multi-lingual [659]. multi-pass [1047, 1037]. multi-path [1013]. multi-processor [385]. multi-task [1016]. multilevel [89]. Multiple [730, 199, 634]. multiprocessor [349]. MUMPS [644]. MUMS [835]. music [296]. Mythical [534].

N [447, 595, 143, 142]. name [721, 853]. names [869]. naming [681]. NASA [47]. National [654]. ’natural [362, 192, 820, 819]. natural-language-like [820]. nature [820]. Navy [331]. NC [1117]. necessary [46]. need [29, 509, 339]. nested [492, 105, 293]. nested-loops [492]. nested-sequential [293]. network [666, 458, 495]. News [886]. Newsletter [761, 299]. next [628]. no [594]. nodes [602]. non [800, 512, 151, 699]. non-desirability [800]. non-determinacy [512]. non-deterministic [151]. non-iterative [699]. nondeterministic [917]. nonprocedural [372]. normal [1008, 829]. notation [221, 281, 283]. Note [451, 1091, 871, 1078, 1079, 900, 986, 591, 473, 851, 601, 610, 611]. Notes [870, 201, 742, 696, 109, 956, 957]. novel [651]. November [1110, 190, 1108]. Number [761, 888, 1008, 975, 199, 476]. Numerical [654, 919, 175]. NY [1106, 1111, 1116].

O [85, 1093]. object [401]. objects [371]. obliterate [339]. Observations [1030]. obtaining [705]. October [1102]. offs [443]. Old [808, 129]. old-line [129]. OLIVER [897]. Omitted [248]. on-conditions [207]. on-line [463]. One [1092, 1063, 977, 610, 798, 114]. One-Pass [1092, 610]. online [535]. onto [646]. operands [392]. operate [561]. Operating [1103, 307, 122, 314, 96, 123, 1027, 269, 1104, 320, 324, 326, 329, 334, 335, 129, 133, 544, 549]. Operation [240, 332]. Operational [183, 264]. operations [498, 378, 910, 496, 230, 470]. Operator [888, 985]. Operators [383, 856, 38, 473, 356, 375, 641]. OPSs [312]. Optimal [251, 393, 897, 866, 810, 107, 355]. optimisation [832]. Optimization [41, 35, 50, 52, 969, 1055, 42, 58, 43, 45, 39, 1058, 598, 1043, 1079, 743, 38, 914, 599, 493, 1056, 282, 451, 1048, 380, 44]. optimizations [37]. Optimized [647, 108, 36]. optimizer [377, 1060, 285]. optimizing [40]. order [863, 251, 247, 1029, 996, 119, 355, 880].

Ordered [864]. Ordering [706]. Oregano [72]. organization [767]. organizations [466]. organized [1094]. Organizing [518]. orientated [245]. oriented [57, 733, 844, 3, 439, 694, 403]. Origins [930]. Orleans [1113]. Orsay [787]. orthogonality [1026]. other [574]. outline [618]. output [417, 850]. overkill [587, 478]. overlays [653]. Overview [150, 681, 1027, 764, 1009, 766, 372, 1068, 1094, 575, 145].

P [419]. PACOL [487]. pair [888]. Panel [13, 17]. paper [248, 447, 904]. papers [1115, 583]. paragraph [105]. Parallel [1111, 404, 714, 151, 51, 487, 359, 405, 385, 485, 500, 484, 428, 181, 590, 230, 494, 111]. Parallelism [872, 873, 482]. Paralyzer [482]. Parameter [912, 870, 721, 671, 1066, 389]. Parameterisation [919]. parameters [1029, 834, 922, 682, 537]. parametric [875, 876]. Paris [197, 787]. Paris-XI [787]. parser [1052]. parsers [53, 848]. parsing [499, 1080, 268, 779, 42, 58, 151, 614, 991, 885]. part [131, 198, 64, 798]. partial [174]. particularly [719]. PASCAL [762, 547, 993, 719, 743, 804, 715, 736, 952, 806, 591, 705, 921, 1057, 883, 1056, 716, 673, 818, 1070, 516, 611, 290, 854, 1091, 1093, 855, 968, 265, 798, 1011, 1064, 548, 1047, 877, 447, 421, 868, 1086, 1067, 892, 1000]. PASCAL-based [818]. Pass [1092, 1047, 492, 1037, 610]. passing [922, 1066]. past [886, 1100]. Patching [887]. path [1013]. paths [395]. Pattern [909, 791, 851]. PDEL [389]. PDEL-ID [389]. PDP [657, 670, 805, 645]. PDP-11 [657, 670, 805, 645]. PDP11 [792]. PDP11/ICL1900 [792]. PDP8 [659]. PEARL [193]. pedagogic [242, 273]. performance [688, 537]. Permanent [282]. phase [926]. physics [722]. PICASSO [218]. Picture [683, 681, 687, 691, 692]. pictures [675]. PL [1046, 1061, 1076, 906, 341, 23, 725, 669, 913, 799, 249, 121, 60, 59, 328, 752, 14, 61, 945, 944, 645, 731, 208, 732, 357, 125, 506, 612, 444, 84, 424, 423]. PL-11 [645]. PL/1 [906, 731, 444]. PL/C [725]. PL/CS [725]. PL/CT [612]. PL/EXUS [357]. PL/I’ [506, 1046, 1061, 1076, 341, 23, 669, 913, 799, 249, 121, 60, 59, 328, 752, 14, 61, 945, 944, 208, 732, 612, 84, 424, 423]. PL/S [125]. placing [247]. plan [281]. Planning [133]. Plauger [419]. plotting [689]. PLUS [54]. point [165, 126, 836, 468]. pointer [802, 633]. Pointers [635]. Polish [358, 841]. politics [210]. polyadic [264]. polygons [214]. polymorphic [10]. POLYP [709]. portability [291]. portable [708, 1045, 780]. Porting [1044]. portions [248]. Positional [671]. potential [165]. power [1032]. powerful [1001]. pp [761]. PPL [10, 167]. Practical [1053, 408, 1064]. Practically [110]. practice [493]. pre [1002, 994]. pre-processor [1002, 994]. precedence [53]. precompiler [455]. precompilers [834]. predicate [1010]. Predicting [1058]. Prediction [915, 537]. prefix [970]. prefix-style [970]. Preliminary [1062, 745, 1024, 687]. Prelude [783]. preparation [193]. Preprocessing [710, 711, 712]. preprocessor [982, 763, 594, 208, 103, 105]. preprocessor-based [208]. preprocessors [605]. present [246, 1100]. presentation [144]. presentations [92]. presented [1115]. preserved [178]. Prettyprinter [1092]. Prettyprinting [921, 1069]. primitive [697, 194, 20]. primitives [683, 259, 161, 235, 239, 25]. principle [485]. principles [559, 387]. priority [976]. Problem [501, 1007, 517, 392, 234]. Problems [121, 951, 41, 359, 472, 1013, 176, 790, 612]. Procedural [340, 1029]. procedure [1058, 1078, 815, 834, 922, 980]. procedureless [374]. procedures [797]. Proceedings [1112, 1104, 1108, 1107, 94, 95, 1102, 1103, 1106, 1110, 1117, 1118, 1120]. Process [755, 754, 322, 661]. processes [816, 310, 311, 227, 69, 552]. processing [1038, 109, 483, 501, 485, 152, 12, 409, 495, 797, 691, 692, 216]. Processor [1113, 351, 385, 1002, 994, 483, 150, 632, 358, 104, 355, 196, 149, 595]. Processor-N [595]. processors [407, 733, 844, 488, 1094]. produce [583]. Product [56]. Production [276, 532, 515, 830, 814]. productions [36]. products [368]. profile [898, 585, 606]. profiling [776]. Program [173, 56, 568, 345, 728, 830, 970, 493, 826, 651, 1074, 812, 1036, 101, 1058, 973, 908, 831, 974, 598, 832, 975, 177, 414, 888, 538, 95, 218, 804, 379, 539, 384, 619, 577, 573, 753, 737, 504, 656, 201, 847, 181, 1060, 228, 849, 417, 368, 1006, 1048, 494, 674, 375, 206, 819, 700, 264, 966]. programmable [821]. programmed [751, 930]. Programmer [401]. programmers [806, 793]. Programming [1103, 1113, 1116, 1119, 221, 862, 381, 526, 225, 1042, 633, 231, 382, 329, 333, 402, 2, 980, 1023, 1090, 294, 1015, 620, 1106, 1110, 502, 613, 951, 171, 621, 744, 1, 295, 863, 929, 933, 937, 695, 800, 532, 939, 702, 360, 647, 762, 413, 252, 822, 1093, 254, 698, 556, 408, 189, 258, 801, 259, 405, 719, 1020, 450, 580, 1026, 223, 832, 378, 242, 701, 226, 205, 593, 946, 301, 304, 366, 745, 270, 759, 476, 1016, 374, 560, 643, 1104, 1097, 924, 949, 1009, 882, 670, 1021, 671, 391, 319, 746, 365, 820, 249]. programming [603, 121, 1025, 995, 102, 273, 622, 370, 144, 399, 714, 661, 442, 996, 324, 680, 160, 484, 740, 609, 976, 325, 447, 421, 275, 384, 277, 931, 51, 119, 410, 1030, 503, 589, 279, 331, 704, 628, 203, 280, 988, 709, 1108, 298, 883, 730, 651, 342, 419, 884, 971, 284, 229, 858, 793, 835, 286, 346, 944, 1086, 645, 739, 1087, 452, 972, 422, 116, 222, 948, 288, 582, 1068, 553, 902, 10, 578, 1074, 682, 338, 375, 575, 145, 68, 183, 479, 209, 1107, 78, 293, 516, 760, 583, 117, 576, 1111, 427]. Programming-in-the [526]. programming-in-the-small [526]. programs [741, 308, 371, 540, 776, 441, 170, 566, 407, 531, 653, 898, 1018, 899, 404, 968, 1008, 262, 713, 512, 669, 913, 551, 267, 588, 411, 522, 227, 1011, 1028, 1064, 777, 767, 562, 529, 833, 1083, 1084, 561, 172, 813, 584, 705, 448, 826, 443, 555, 921, 4, 406, 868, 1069, 510, 769, 585, 590, 188, 550, 193, 1088, 567, 238, 292, 418, 1101, 530]. Progress [211, 203]. project [365, 322, 915, 120, 202]. Prolog [823]. prone [270]. Proof [403, 129]. Proof-oriented [403]. proofs [674, 811]. Proper [227]. Properties [178, 748, 255, 749, 750, 625, 172, 188, 71]. PROPHET [257]. proposal [132, 221, 865, 727, 736, 953, 397, 489]. Proposals [839]. Proposed [617, 799, 607, 808, 616, 1076, 726, 414, 127, 111]. protection [744, 364, 326, 305, 672, 546, 340]. Proteus [11, 164]. Provable [560]. provide [96]. Proving [562, 326, 656, 531, 172]. proximity [473]. PS440 [116]. PSI [826]. Psychological [292, 418]. psychology [582]. Pun [583]. Pun-dora [583]. Purpose [938, 621, 547, 6, 548, 989, 440, 293]. purposes [676].

Quad [22]. Quantifying [777].

Quantitative [541]. queries [460]. query [462]. quest [273, 242]. question [700]. question-answering [700]. Quickscan [851]. QUILL [277]. Quote [22]. Quote-Quad [22].

R [776]. R. [904]. Radar [796]. radix [184]. Raleigh [1117]. RAMBAM [283]. range [110, 570]. raster [690, 685, 686]. Rationale [1025, 818]. RDL [1096]. Readability [891, 970, 1084, 448]. readable [899, 584]. real [652, 1076, 799, 887, 687, 296, 535, 203, 760, 438]. real-time [652, 1076, 799, 887, 687, 203, 760, 438]. realism [613]. reason [485]. reasoning [824]. recall [848]. recogniser [185]. recognition [3, 230, 355]. recognizers [36]. Recovery [1053, 42, 58, 755, 848, 1052]. recursed [49]. recursion [177, 178]. Recursive [174, 604, 566, 813, 406, 1050, 264]. recursive-descent [1050]. RED [962]. redefinition [831]. REDL [1012]. REDUCE [965]. reduction [385, 567]. redundant [251]. REFAL [998]. Reference [1024, 762, 359, 852, 496, 731, 732, 954, 955]. references [707, 969, 517]. refinement [874, 333]. Reflections [314, 117]. Register [953, 1059]. regular [74]. related [295]. relation [417]. relational [621, 460, 405, 410, 623, 83]. relations [728]. relationships [100, 847]. relative [769]. Reliability [542, 912, 707, 571, 511, 271, 759, 572, 536, 751, 535, 554]. Reliable [1109, 1117, 527, 565, 525, 533, 515, 327, 1066, 528, 544, 514, 513]. Remarks [85, 472, 809, 778]. reminder [867]. remote [1065]. Removal [492, 178]. repair [992]. repertoire [953]. repetition [1005]. repetitive [1082]. replacing [843]. Reply [447, 506]. Report [303, 300, 301, 304, 740, 306, 305, 133, 302, 1003, 734, 894, 773, 365, 220, 286, 356, 771, 1072]. reporting [471]. Representation [82, 21, 48, 773]. representations [635, 828, 470]. required [361]. requirement [954, 956, 955, 957]. Requirements [906, 133, 863, 631, 814]. Research [627, 295, 593, 286, 1094]. resented [574]. reserved [1080]. Resource [388, 976, 343]. Restricted [748, 363, 852]. restrictions [509]. results [392]. retention [361]. Retrieval [470, 1110, 463, 466, 467, 458, 464, 465, 473, 1108, 474, 858, 461, 634, 462]. reverse [841]. Review [200, 365, 419, 2, 524, 209]. reviews [426]. Revised [863, 761, 425, 771, 734]. revisions [367]. right [450]. robot [1016]. Rochester [1116]. role [1029, 751]. rooster [879]. Rosetta [202]. Roster [452, 972]. rough [705]. routines [441]. rules [424, 423]. Run [776, 359, 743]. Run-time [776, 743].

S [990, 125]. S-SNOBOL [990]. safe [906]. SAGES [729]. Salt [1112]. Santa [1105]. satellite [434]. Savannah [1103, 1104]. SBA [620]. scalar [88]. scale [525, 278, 514]. SCAMP [776]. scan [20]. scanning [986]. Scarcely [1086]. schemata [179, 181, 264]. scheme [641]. schemes [177, 379]. scholars [396]. Schwartz [365]. science [762, 881, 129]. Sciences [543]. scientific [286]. Scope [866, 1067, 1039, 869, 422]. Scott [169]. SCRATCHPAD [219, 394]. searching [393]. SEAS [94, 95]. Secondary [475]. section [782]. secure [544]. Security [545]. seems [850]. SELECT [540]. Selection [95, 828, 569, 827, 1015]. selective [1082]. selector [608]. Self [831, 520, 564]. self-checking [564]. self-metric [520]. Self-redefinition [831]. Semantic [310, 311, 1038, 267, 319, 158, 333, 390]. Semantics [287, 360, 300, 1085, 432, 451, 684, 183, 125, 264]. semiautomatic [380]. semicolons [1011]. seminar [2]. semiotic [1]. sense [97]. sentences [1030]. separate [952, 1049]. September [197, 191]. sequence [31, 847]. sequential [381, 73, 356, 293]. serial [111]. session [303, 300, 301, 304, 306, 305, 302]. Set [402, 702, 55, 632, 439, 767, 362, 289]. SETL [365, 1073, 380]. sets [1054]. several [1020, 1021]. SFOR [455]. Shaded [677]. Shape [214]. Sharing [659, 122, 123, 401, 116]. shifting [860]. short [320, 144, 986]. shortest [395]. sic [245]. side [926, 925]. SIGGRAPH [1115]. SIGIR [1110, 1108]. SIGMA [265]. SIGMICRO [1106, 1107]. SIGMIN [1113]. SIGMIN/SIGPLAN [1113]. signed [986]. SIGOPS [1104]. SIGPLAN [1102, 1103, 1105, 1106, 1110, 1113, 1119, 1120, 929, 933, 937, 939, 1115, 946, 1104, 924, 931, 1108, 944, 948, 1107]. SIGPLAN-SIGIR [1110, 1108]. SIGPLAN-SIGMICRO [1106, 1107]. SIGPLAN-SIGOPS [1104]. SIGPLAN-SIGTOPS [1103]. SIGTOPS [1103]. SIMPL [335]. simple [1001, 896, 588, 377, 705, 196]. simplify [619]. SIMULA [946, 723, 947, 368, 672, 156]. Simulation [938, 170, 387, 391, 694, 796]. single [872, 873, 492, 105]. single-paragraph [105]. SITBOL [446]. situation [699]. Sixty [21, 48]. Sixty-character [21, 48]. size [736]. skip [4]. SL5 [764]. Small [1113, 653, 526, 316, 135, 278, 655, 664]. small-scale [278]. Smalltalk [1068]. Smalltalk-72 [1068]. SNOBAT [838]. SNOBOL [446, 367, 949, 948, 91, 990, 606]. SNOBOL4 [412, 250, 62, 735, 910, 445, 61, 510, 63, 838, 851, 990]. sociology [210]. Software [1109, 1117, 558, 315, 751, 535, 291, 912, 650, 525, 649, 571, 518, 657, 533, 511, 881, 638, 919, 515, 1096, 648, 323, 327, 536, 722, 665, 655, 554, 557, 563, 528, 337, 544, 541, 915, 552, 520, 537, 514, 513, 527, 564, 84, 729]. solution [392]. Some [525, 566, 109, 446, 636, 1028, 1064, 747, 625, 8, 1004, 809, 964, 584, 1013, 790, 138, 717, 1087, 18, 10, 778, 112, 200, 900, 356, 402]. sorting [184]. Source [1055]. SP [911, 988]. SP-grammars [988]. SP/k [911]. space [321, 1037]. spaceflight [224]. spaghetti [817]. SPARC [133]. sparse [910]. Speakeasy [398, 297, 299]. Special [139, 762, 89]. specialized [166]. specific [188]. Specification [308, 411, 521, 731, 732, 741, 973, 1077, 748, 974, 262, 263, 900, 976, 846, 228, 283, 432, 549, 545, 400]. Specifications [1022, 747, 1004, 1032, 861]. Specifying [460, 847, 409, 757, 682]. SPECL [338]. spectrum [973]. speed [268, 206]. SPIL [327]. sponsored [1115]. Srinivasan [1101]. stack [1022, 837, 817, 999, 1075]. Standard [1003, 64, 808, 1061, 341, 726, 773, 328, 783, 1072]. Standards [654, 133]. Star [766]. starvation [976]. State [98, 538, 985, 537]. statement [864, 597, 703, 980, 453, 699, 601, 610]. statements [715, 1085, 492, 842]. Static [749, 898, 750, 585, 1034]. statistics [1028]. step [687, 628]. stepwise [874]. Stickiness [997, 1006]. Sticky [914]. Stone [202]. Storage [784, 237, 441, 866, 466, 467, 1043, 443, 858, 71, 75]. strategies [866]. strategy [831, 1034, 1052]. Strathclyde [1118]. STRCMACS [702]. stream [495]. String [64, 791, 126, 358]. strong [373]. strongly [1063]. structural [913]. Structure [1112, 351, 358, 857, 412, 862, 77, 414, 192, 728, 570, 753, 737, 60, 1031, 73, 847, 709, 978, 70, 554, 478, 159, 563, 343, 68, 83, 819, 142]. Structured [502, 532, 596, 763, 968, 832, 701, 594, 421, 384, 503, 579, 717, 386, 902, 545, 576, 800, 702, 592, 254, 698, 908, 469, 899, 713, 450, 593, 301, 882, 241, 248, 480, 562, 995, 69, 370, 1049, 325, 857, 536, 834, 504, 706, 353, 449, 589, 443, 279, 988, 298, 455, 1069, 793, 615, 869, 724, 739, 422, 549, 288, 193, 990, 238, 583, 966]. structuredness [1088, 1101]. structures [76, 917, 381, 631, 256, 23, 688, 226, 411, 517, 1010, 630, 640, 322, 383, 82, 454, 561, 599, 81, 362, 589, 607, 768, 79, 282, 336, 287, 828, 731, 732, 627, 456, 386, 839, 244, 80, 75]. Structuring [653, 859, 1001, 755, 628]. student [769]. studies [217]. study [744, 189, 913, 522, 1073, 430]. studying [267]. style [970, 419, 375]. sub [126]. sub-language [126]. subcommittee [132]. subexpression [33]. subexpressions [251]. sublanguage [621, 772]. submitted [959, 960, 962, 961]. Subroutine [161, 235, 690, 234]. subroutines [767]. subscripted [969]. subset [725, 798, 182, 923, 63, 789]. subsets [84]. subsystem [54]. SUE [120, 320]. SUE/360/370 [320]. sufficient [485]. SUGAR [680]. Suggested [367]. Suggestions [974]. suitable [443]. Summary [1020, 1021, 927, 929, 933, 937, 939, 92, 946, 924, 931, 944, 948]. summing [918]. super [783]. super-language [783]. supercompiler [998]. superset [268]. supplement [734]. support [973, 649, 279, 836]. Supporting [981, 648]. Survey [704, 115, 226, 113]. switching [841]. symbol [344, 1039, 55, 345, 346, 343]. symbol-2R [343]. symbolic [540]. symbolism [65]. symbols [21, 48]. Symposium [1105, 1116, 1120, 200, 1102, 1115, 201, 1114]. synchronization [872, 873, 239, 755]. synonyms [715]. SYNPROC [150]. Syntactic [424, 154, 423, 36]. Syntax [413, 148, 1098, 992, 150, 367, 42, 58, 213, 1052, 425, 432, 1074, 125, 149]. syntax-directed [42, 58, 1074]. synthesis [810, 384, 813, 826, 812, 700]. Synthesizer [482, 1074]. SYSL [457]. System [309, 543, 568, 1099, 965, 563, 497, 695, 350, 30, 540, 649, 254, 556, 408, 897, 653, 257, 258, 314, 469, 96, 131, 198, 398, 533, 261, 262, 551, 267, 387, 205, 269, 690, 1016, 689, 376, 318, 694, 320, 121, 321, 648, 830, 471, 135, 160, 215, 609, 213, 274, 685, 686, 326, 327, 278, 465, 623, 332, 709, 335, 70, 133, 546, 544, 814, 116, 549, 288, 193, 434, 470, 389, 457, 840, 998, 338, 145, 700, 164, 217, 202, 424, 120, 938, 1095, 225, 1042, 729, 332, 659, 620, 355]. systematic [115]. Systems [1103, 1113, 1102, 307, 463, 683, 122, 466, 1071, 260, 719, 467, 123, 115, 638, 1027, 304, 316, 366, 745, 212, 1104, 670, 603, 648, 805, 210, 324, 160, 275, 276, 119, 458, 464, 329, 665, 987, 203, 334, 1013, 146, 858, 557, 337, 113, 208, 114, 664, 117, 118, 438].

T [365]. table [101, 102, 104, 103, 105, 112, 1071]. tables [1018, 109, 1039, 110, 100, 107, 111, 106, 991, 108]. tablet [221]. Tabsol [103]. tabular [598, 80]. tactical [472]. Tactics [254]. tagged [353]. taming [802]. Target [1055]. TARTAN [955, 957]. TARTAN-language [955, 957]. task [1016]. Tasking [328, 186]. TC1O [59]. Teaching [249, 670, 220, 421, 342]. tears [775]. technical [863]. technique [896, 667, 974, 899, 882, 632, 204, 184, 4, 878, 340]. techniques [52, 499, 501, 387, 51, 521, 298, 282, 524, 18, 430, 339, 112]. Teleprocessing [543]. Teletypes [55]. tells [820]. TELOS [818]. templates [1035]. Ten [29]. tense [730]. term [450]. terminal [85, 223]. termination [227]. test [569, 777, 722, 201, 553]. testability [559]. testcase [519]. Testing [571, 448, 557, 540, 539, 570, 829, 674]. tests [518]. Texas [497]. text [437, 277, 474]. their [697, 174, 551, 1030, 716]. theorem [478]. theory [1, 569, 976, 722, 493, 289, 402]. things [574]. third [99]. third-generation [99]. thriving [887]. time [652, 1076, 361, 122, 776, 359, 123, 743, 799, 887, 687, 296, 443, 535, 203, 14, 509, 116, 760, 659, 438]. time-sharing [122, 123, 116]. time/storage [443]. times [553]. timing [1078]. timings [980]. TINY [696]. Today [2]. tolerance [565, 563]. tolerant [884]. TOMAL [661]. tool [707, 558, 100, 705, 508, 368, 116]. Tools [510, 1045, 1100, 930, 980]. top [614, 885]. top-down [614, 885]. total [89]. touch [850]. Trac [20]. trace [588, 504]. Traces [718]. tracing [807]. Tracking [796]. trade [443]. trade-offs [443]. Transaction [508]. transfer [897]. Transferability [266, 330, 325, 302]. Transform [26]. transformation [375]. Transformations [11, 379, 566]. transient [282]. Transition [158, 651, 1050]. Translating [177, 458]. TRANslation [139, 1018, 344, 667, 897, 148, 614, 213, 330, 790]. Translator [965, 101, 592, 269, 480, 424]. translators [697]. transmission [721]. transportable [919, 1047, 663]. transportation [785]. transput [782]. Traversable [999, 1075, 1022]. Tree [238, 407, 395]. Tree-structured [238]. trees [602, 107, 397]. truth [942]. TSO [299]. TSO-speakeasy [299]. turtle [685, 686]. TV [685, 686]. Two [951, 359, 225, 995, 224, 212, 321, 210, 213, 691, 692, 209, 710, 612, 217, 225]. TWO-D [225]. Two-Dimensional [225, 995, 224, 212, 210, 213, 691, 692, 209, 710, 217]. two-level [321]. Type [1063, 1077, 749, 1071, 750, 1035, 1004, 953, 626, 903, 641]. typed [1063, 984, 710]. Types [553, 864, 132, 157, 748, 889, 580, 1081, 759, 636, 630, 622, 155, 642, 809, 619, 382, 577, 846, 639, 716]. typing [373].

UCC [2]. UK [1118]. UMTA [400]. unary [38]. undecidability [181]. understandable [270]. understanding [753, 626]. Unfurling [341]. unified [608, 836]. Uniform [837, 517, 731, 732]. unions [749, 750]. UNIVAC [507]. Universal [1095]. University [365, 787, 1115, 98, 67]. unstructured [370]. until [420]. USA [1102, 1103, 1105, 1106, 1109, 1110, 1111, 1113, 1114, 1116, 1117, 1119, 1120]. Use [124, 1018, 597, 975, 115, 660, 270, 694, 365, 220, 619, 922, 1050, 869, 645, 106, 18, 634]. used [122, 123, 881, 113]. user [800, 96, 165, 20, 584, 114]. user-defined [584]. Users [92]. USF [225]. Using [666, 797, 1066, 650, 776, 657, 52, 669, 501, 1045, 74, 38, 249, 728, 599, 755, 781, 796]. Utah [1112]. Utilization [101]. utilizing [641]. Utopia [982].

validation [912, 558, 184, 541, 193, 520]. valuations [188, 403]. value [356]. values [781]. variable [736, 570, 422, 243]. variabled [1086]. Variables [364, 707, 748, 914, 633, 363, 639, 1006]. Vector [1111, 491, 486, 499, 230, 494, 489]. Veloso [1022]. verifiable [741, 638, 546]. Verification [568, 788, 170, 573, 524, 757]. Verifying [561, 567, 530]. Version [745]. versus [526, 671, 779, 392]. Very [1105, 242, 391, 273, 396, 826]. VI [79]. via [102, 830, 651]. view [1061, 907, 165, 219, 803, 901, 370, 920, 573, 600, 114, 468, 644]. virtual [407, 603, 357]. Visibility [642]. vocabularies [461]. Volume [761].

W [419]. Washington [67]. WATFOR [34]. where [168]. which [561]. while [420, 465, 896]. while-until [420]. whole [942]. wide [973]. will [996]. wishbook [307]. within [597]. without [763, 775, 1080, 148, 633, 505]. word [246]. words [1080]. Working [94, 1076]. Workshop [92]. world [887]. wrap [583]. writability [891]. write [314, 555]. writer [307]. Writing [477, 965, 30, 274, 84, 424].

X [660]. X3.5 [1003]. X3.5-1970 [1003]. X3J1 [59]. X3J3 [132, 761]. XDX [27]. Xerox [265]. XI [787]. XL [978]. XPL [507, 879].

YELLOW [961]. York [1111, 365, 1107].

Zed [1044]. ZENO [1062].

References

Anonymous:1968:TST

[1] Anonymous. Toward a semiotic theory of programming languages. ACM SIG-PLAN Notices, 3(5):24–27, May 1968. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1969:RUS

[2] C. J. Shaw. Review of UCC seminar on “Programming Languages Today”. ACM SIGPLAN Notices, 4(2): 24–25, February 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Floyd:1969:MOR

[3] Robert W. Floyd. A machine-oriented recognition algorithm for context-free languages. ACM SIGPLAN Notices, 4(5):28–29, May 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nievergelt:1969:BST

[4] J. Nievergelt and M. I. Irland. Bounceand-skip: a technique for directing the flow of control in programs. ACM SIG-PLAN Notices, 4(5):29–31, May 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perlis:1969:IEL

[5] Alan J. Perlis. Introduction to extensible languages. ACM SIGPLAN Notices, 4(8):3–5, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garwick:1969:GGP

[6] Jan V. Garwick. GPL, a general purpose language. ACM SIGPLAN Notices, 4(8):6–8, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mailloux:1969:AEL

[7] Barry J. Mailloux and John E. L. Peck. Algol 68 as an extensible language. ACM SIGPLAN Notices, 4 (8):9–13, August 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jorrand:1969:SAB

[8] Philippe Jorrand. Some aspects of Basel, the base language for an extensible language facility. ACM SIG-PLAN Notices, 4(8):14–17, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Irons:1969:EFI

[9] Edgar T. Irons. The extension facilities of IMP. ACM SIGPLAN Notices, 4 (8):18–19, August 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Standish:1969:SFP

[10] Thomas A. Standish. Some features of PPL, a polymorphic programming language. ACM SIGPLAN Notices, 4 (8):20–26, August 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bell:1969:TEF

[11] James R. Bell. Transformations: the extension facility of Proteus. ACM SIGPLAN Notices, 4(8):27–31, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLaren:1969:MPE

[12] M. Donald MacLaren. Macro processing in EPS. ACM SIGPLAN Notices, 4 (8):32–36, August 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1969:PLA

[13] Christopher J. Shaw. Panel of language authors. ACM SIGPLAN Notices, 4 (8):37–39, August 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nicholls:1969:PCT

[14] John Nicholls. PL/I compile time extensibility. ACM SIGPLAN Notices, 4 (8):40–44, August 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheatham:1969:MEL

[15] Thomas E. Cheatham, Jr. Motivation for extensible languages. ACM SIGPLAN Notices, 4(8):45–49, August 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McIlroy:1969:AEL

[16] M. D. McIlroy. Alternatives to extensible languages. ACM SIGPLAN Notices, 4(8):50–52, August 1969. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1969:PCE

[17] Peter Wegner. Panel on the concept of extensibility. ACM SIGPLAN Notices, 4(8):53–54, August 1969. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Standish:1969:SCC

[18] Thomas A. Standish. Some compilercompiler techniques for use in extensible languages. ACM SIGPLAN Notices, 4(8):55–62, August 1969. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bobrow:1969:LB

[19] D. G. Bobrow. LISP bulletin. ACM SIGPLAN Notices, 4(9):17–57, September 1969. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kagan:1970:ETL

[20] Claude A. R. Kagan. Experimental Trac language primitive to allow user control of the scan algorithm. ACM SIGPLAN Notices, 5(1–2):27–28, 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Charmonman:1970:SCRa

[21] Sak Charmonman. Sixty-character representation of APL symbols. ACM SIGPLAN Notices, 5(1–2):28–31, 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1970:AQQ

[22] Christopher J. Shaw. APL Quote-Quad. ACM SIGPLAN Notices, 5 (3):21, March 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Charmonman:1970:CSA

[23] S. Charmonman. A comparison of the structures of APL, FORTRAN, AL-GOL and PL/I. ACM SIGPLAN Notices, 5(3):22–24, March 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1970:GHF

[24] Alan L. Jones. The generalized hypergeometric function. ACM SIGPLAN Notices, 5(3):26–27, March 1970. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lauer:1970:AMP

[25] John E. Lauer. APL\1130 missing primitives. ACM SIGPLAN Notices, 5(3):28, March 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1970:FFT

[26] Alan R. Jones. Fast Fourier Transform. ACM SIGPLAN Notices, 5 (3):29, March 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lauer:1970:IXD

[27] John E. Lauer. Intgrl XDX does Gaussian integration. ACM SIG-PLAN Notices, 5(3):30, March 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bickart:1970:CEM

[28] T. A. Bickart. Characteristic equation of a matrix. ACM SIGPLAN Notices, 5(3):31, March 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ledgard:1970:TML

[29] Henry F. Ledgard. Ten mini-languages in need of formal definition. ACM SIGPLAN Notices, 5(4–5):14–37, 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Book:1970:CSC

[30] Erwin Book, Dewey Val Shorre, and Steven J. Sherman. The CWIC/36O system, a compiler for writing and implementing compilers. ACM SIG-PLAN Notices, 5(6):11–29, June 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bickart:1970:FAI

[31] Theodore A. Bickart. Function to accelerate and/or induce sequence convergence. ACM SIGPLAN Notices, 5 (6):38–41, June 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:1970:CFA

[32] Frances E. Allen. Control flow analysis. ACM SIGPLAN Notices, 5(7):1– 19, July 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cocke:1970:GCS

[33] John Cocke. Global common subexpression elimination. ACM SIG-PLAN Notices, 5(7):20–24, July 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cowan:1970:DCW

[34] D. D. Cowan and J. W. Graham. Design characteristics of the WATFOR compiler. ACM SIGPLAN Notices, 5 (7):25–36, July 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheatham:1970:OACa

[35] T. E. Cheatham, Jr. and Thomas A. Standish. Optimization aspects of compiler-compilers. ACM SIGPLAN Notices, 5(7):37, July 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haynes:1970:COS

[36] H. R. Haynes and L. J. Schutte. Compilation of optimized syntactic recognizers from Floyd-Evans productions. ACM SIGPLAN Notices, 5(7):38–51, July 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bagwell:1970:LO

[37] John T. Bagwell, Jr. Local optimizations. ACM SIGPLAN Notices, 5 (7):52–66, July 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frailey:1970:EOU

[38] Dennis J. Frailey. Expression optimization using unary complement operators. ACM SIGPLAN Notices, 5(7):67– 85, July 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aho:1970:FAC

[39] Alfred V. Aho, Ravi Sethi, and J. D. Ullman. A formal approach to code optimization. ACM SIGPLAN Notices, 5 (7):86–100, July 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Painter:1970:EOC

[40] James A. Painter. Effectiveness of an optimizing compiler for arithmetic expressions. ACM SIGPLAN Notices, 5 (7):101–126, July 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bolas:1970:OPE

[41] Bruce J. Bolas. Optimization problems in extensible compilers. ACM SIGPLAN Notices, 5(7):127, July 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LaFrance:1970:OERa

[42] Jacques E. LaFrance. Optimization of error recovery in syntax-directed parsing algorithms. ACM SIGPLAN Notices, 5(7):128, July 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Millstein:1970:OACa

[43] Robert E. Millstein. Optimization for an array computer. ACM SIG-PLAN Notices, 5(7):129, July 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zwakenberg:1970:CO

[44] Richard G. Zwakenberg. CDC 6600/ 7600 optimization. ACM SIGPLAN Notices, 5(7):130, July 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Millstein:1970:OACb

[45] Robert E. Millstein, Edward P. Krugman, and David Goldberg. Optimization for an array computer. ACM SIG-PLAN Notices, 5(9):8–13, September 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1970:IIM

[46] Daniel M. Berry. The importance of implementation models in ALGOL 68: or how to discover the concept of necessary environment. ACM SIGPLAN Notices, 5(9):14–24, September 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Korn:1970:ACM

[47] Karl Korn. APL community meets at NASA. ACM SIGPLAN Notices, 5(9): 27–28, September 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Charmonman:1970:SCRb

[48] S. Charmonman. Sixty-character representation of APL symbols. ACM SIG-PLAN Notices, 5(9):29–34, September 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McEwan:1970:AR

[49] Alan T. McEwan and Don W. A. Watson. APL/360 recursed! ACM SIG-PLAN Notices, 5(9):35–40, September 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheatham:1970:OACb

[50] Thomas E. Cheatham and Thomas A. Standish. Optimization aspects of compiler-compilers. ACM SIGPLAN Notices, 5(10):10–17, October 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lincoln:1970:PPT

[51] Neil Lincoln. Parallel programming techniques for compilers. ACM SIG-PLAN Notices, 5(10):18–31, October 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeRemer:1970:OUL

[52] Franklin L. DeRemer. Optimization using LR(k) techniques. ACM SIG-PLAN Notices, 5(10):32–34, October 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Horning:1970:ECL

[53] J. J. Horning and W. R. Lalonde. Empirical comparison of LR(k) and precedence parsers. ACM SIGPLAN Notices, 5(11):10–24, November 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Breed:1970:APF

[54] Lawrence M. Breed. The APL PLUS file subsystem. ACM SIGPLAN Notices, 5(11):29–31, November 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hagerty:1970:ASS

[55] P. E. Hagerty. An APL symbol set for Model 35 Teletypes. ACM SIG-PLAN Notices, 5(11):31–33, November 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Breed:1970:EFA

[56] Lawrence M. Breed and Roger D. Moore. Errors and fixes in the APL/ 360 Program Product. ACM SIGPLAN Notices, 5(11):33–38, November 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Charmonman:1970:GAA

[57] S. Charmonman. A generalization of APL array-oriented concept. ACM SIGPLAN Notices, 5(11):38– 42, November 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LaFrance:1970:OERb

[58] Jacques LaFrance. Optimization of error recovery in syntax-directed parsing algorithms. ACM SIGPLAN Notices, 5(12):2–17, December 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLaren:1970:PLD

[59] M. Donald MacLaren. PL/I language development in ANSI/X3J1 and ECMA/TC1O. ACM SIGPLAN Notices, 5(12):18–29, December 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLaren:1970:DMD

[60] M. Donald MacLaren. Data matching, data alignment, and structure mapping in PL/I. ACM SIGPLAN Notices, 5(12):30–43, December 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pagan:1970:MLP

[61] F. G. Pagan. A method of linking PL/I to SNOBOL4. ACM SIGPLAN Notices, 5(12):46–49, December 1970. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Duquet:1970:ES

[62] Robert T. Duquet. Eliza in SNOBOL4. ACM SIGPLAN Notices, 5(12):50– 60, December 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Santos:1970:CSS

[63] Paul J. Santos and W. Douglas Maurer. Compilation of a subset of SNOBOL4. ACM SIGPLAN Notices, 5(12):60– 68, December 1970. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lambird:1971:SEA

[64] Robert J. Lambird. String enhancements to ANSI Standard Fortran (part 1). ACM SIGPLAN Notices, 6(1):5–12, January 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McMurchie:1971:LCA

[65] Tom McMurchie. A limited-character APL symbolism. ACM SIGPLAN Notices, 6(1):15–16, January 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bunderson:1971:HGA

[66] C. Victor Bunderson. On hobbits and the growth of APL. ACM SIG-PLAN Notices, 6(1):17, January 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zosel:1971:UWI

[67] Mary Zosel. University of Washington implementation. ACM SIG-PLAN Notices, 6(1):18–30, January 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1971:DSM

[68] Peter Wegner. Data structure models for programming languages. ACM SIGPLAN Notices, 6(2):1–54, February 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnston:1971:CMB

[69] John B. Johnston. The contour model of block structured processes. ACM SIGPLAN Notices, 6(2):55–82, February 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Organick:1971:DSM

[70] E. I. Organick and J. G. Cleary. A data structure model of the B6700 computer system. ACM SIGPLAN Notices, 6(2): 83–145, February 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walk:1971:MSP

[71] Kurt Walk. Modelling of storage properties of higher level languages. ACM SIGPLAN Notices, 6(2):146– 170, February 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1971:IO

[72] Daniel M. Berry. Introduction to Oregano. ACM SIGPLAN Notices, 6 (2):171–190, February 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McGowan:1971:ESA

[73] Clement McGowan and Peter Wegner. The equivalence of sequential and associate information structure models. ACM SIGPLAN Notices, 6(2): 191–216, February 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferguson:1971:GAL

[74] W. A. Ferguson and G. J. Lipovski. A generalized assembly language using regular expressions. ACM SIG-PLAN Notices, 6(2):217–236, February 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanDam:1971:DSS

[75] Andries van Dam. Data and storage structures for interactive graphics. ACM SIGPLAN Notices, 6(2): 237–267, February 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrams:1971:DSC

[76] Marshall D. Abrams. Data structures for computer graphics. ACM SIG-PLAN Notices, 6(2):268–286, February 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Busam:1971:SDC

[77] Vincent A. Busam. On the structure of dictionaries for compilers. ACM SIG-PLAN Notices, 6(2):287–305, February 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wile:1971:EDA

[78] David S. Wile and C. M. Geschke. Efficient data accessing in the programming language Bliss. ACM SIGPLAN Notices, 6(2):306–320, February 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morris:1971:GDS

[79] James B. Morris, Jr. and Mark B. Wells. Generalized data structures in Madcap VI. ACM SIGPLAN Notices, 6(2):321–336, February 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilkes:1971:ATD

[80] M. V. Wilkes. Associative tabular data structures. ACM SIGPLAN Notices, 6(2):337–345, February 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lipovski:1971:DSA

[81] G. J. Lipovski. On data structures in associative memories. ACM SIG-PLAN Notices, 6(2):346–365, February 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Katzan:1971:RMD

[82] Harry Katzan, Jr. Representation and manipulation of data structures in APL. ACM SIGPLAN Notices, 6(2): 366–397, February 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weston:1971:CRD

[83] Paul E. Weston and Stephen M. Taylor. Cylinders: a relational data structure. ACM SIGPLAN Notices, 6(2): 398–416, February 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vPeschke:1971:PSS

[84] Joachim v. Peschke. PL/I subsets for software writing. ACM SIG-PLAN Notices, 6(4):16–22, May 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrams:1971:RAT

[85] Philip S. Abrams. Remarks about terminal I/O in APL/360. ACM SIG-PLAN Notices, 6(4):33, May 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Linden:1971:AFA

[86] Theodore A. Linden. The Ackermann function in APL. ACM SIG-PLAN Notices, 6(4):34–35, May 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rose:1971:MAM

[87] A. J. Rose. More about multi-adic functions. ACM SIGPLAN Notices, 6 (5):10–11, July 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Breed:1971:GAS

[88] Larry Breed. Generalizing APL scalar extension. ACM SIGPLAN Notices, 6 (5):12–14, July 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Korn:1971:AFA

[89] Karl Korn and John Rose. APL-FIOS; APL-FIDOS: multilevel total file access and special features. ACM SIG-PLAN Notices, 6(5):14–18, July 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bonyun:1971:AA

[90] D. A. Bonyun. APL algorithms. ACM SIGPLAN Notices, 6(5):21–31, July 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1971:SB

[91] Christopher J. Shaw. SNOBOL bulletin. ACM SIGPLAN Notices, 6(6):3– 5, July 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1971:SPA

[92] John R. Clark, Thomas R. Dickson, William H. Marshall, and Angeloa Segalla. A summary of the presentations at APL Users Conference Workshop 3. ACM SIGPLAN Notices, 6 (6):3–13, July 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1971:FIBa

[93] Christopher J. Shaw. FORTRAN information bulletin. ACM SIGPLAN Notices, 6(6):6–22, July 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gellert:1971:APS

[94] Niels Gellert. Abstract of the proceedings at the SEAS APL Working Committee’s meeting Grenoble, Feb. 15th, 16th/71. ACM SIGPLAN Notices, 6 (6):14–15, July 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hendricks:1971:SMP

[95] Hendricks. Selection of a medium for program exchange (appendix II to SEAS APL proceedings). ACM SIG-PLAN Notices, 6(6):15–16, July 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clementi:1971:MASa

[96] J. F. Clementi and A. P. Fletcher. Modifications to the APL 1130/system to provide more convenient operating on a Fortran user’s machine. ACM SIG-PLAN Notices, 6(6):16–18, July 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bonyun:1971:MSA

[97] David A. Bonyun. Mark sense APL. ACM SIGPLAN Notices, 6(6):18–19, July 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greenberg:1971:AMS

[98] Lew Greenberg. APL/6500 at Michigan State University. ACM SIG-PLAN Notices, 6(6):20–21, July 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silberg:1971:DTG

[99] Bruce Silberg. Detab/65 in thirdgeneration Cobol. ACM SIGPLAN Notices, 6(8):4–8, September 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LaFleur:1971:DTT

[100] Thomas G. LaFleur. Decision tables: a tool for documenting logical condition relationships. ACM SIGPLAN Notices, 6(8):9–12, September 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arnold:1971:UDT

[101] Henry O. Arnold. Utilization of a decision table translator for basic program creation. ACM SIGPLAN Notices, 6(8):12–19, September 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jarvis:1971:APD

[102] J. M. Jarvis. An analysis of programming via decision table compilers. ACM SIGPLAN Notices, 6(8): 20–32, September 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sterbenz:1971:TDT

[103] Robert F. Sterbenz. Tabsol decision table preprocessor. ACM SIGPLAN Notices, 6(8):33–40, September 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lemoine:1971:ADL

[104] Donald J. Lemoine. An automatic decision-logic-table processor. ACM SIGPLAN Notices, 6(8):40–45, September 1971. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weisbard:1971:DDT

[105] Michael F. Weisbard. Detap/55: a decision table preprocessor for generating single-paragraph, fully-nested Cobol code. ACM SIGPLAN Notices, 6(8): 45–53, September 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scowen:1971:UDT

[106] R. Scowen. The use of decision tables in Babel. ACM SIGPLAN Notices, 6(8): 54–68, September 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rabin:1971:CLE

[107] Jonas Rabin. Conversion of limitedentry decision tables into optimal decision trees: fundamental concepts. ACM SIGPLAN Notices, 6(8):68–74, September 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Woods:1971:OCG

[108] C. Geoffrey Woods and Mary K. Hawes. Optimized code generation from extended-entry decision tables. ACM SIGPLAN Notices, 6(8):74–80, September 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dailey:1971:SNP

[109] W. H. Dailey. Some notes on processing limited-entry decision tables. ACM SIGPLAN Notices, 6(8):81–89, September 1971. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1971:PCD

[110] William J. Harrison. Practically complete decision tables: a range approach. ACM SIGPLAN Notices, 6(8): 89–93, September 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schwartz:1971:LDT

[111] Bennett M. Schwartz. Lisp 1.5 decision tables implemented for a serial computer and proposed for parallel computers. ACM SIGPLAN Notices, 6(8): 93–103, September 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yasui:1971:SAD

[112] Toshio Yasui. Some aspects of decision table conversion techniques. ACM SIG-PLAN Notices, 6(8):104–111, September 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sammet:1971:BSL

[113] Jean E. Sammet. Brief survey of languages used for systems implementation. ACM SIGPLAN Notices, 6(9): 1–19, October 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Slimick:1971:CSI

[114] John Slimick. Current systems implementation languages: One user’s view. ACM SIGPLAN Notices, 6(9): 20–28, October 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Farber:1971:SSU

[115] David J. Farber. A survey of the systematic use of macros in systems building. ACM SIGPLAN Notices, 6(9): 29–36, October 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sapper:1971:PLP

[116] Gerd R. Sapper. The programming language PS440 as a tool for implementing a time-sharing system. ACM SIGPLAN Notices, 6(9):37–41, October 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wulf:1971:RSP

[117] W. Wulf, C. Geschke, D. Wile, and J. Apperson. Reflections on a systems programming language. ACM SIGPLAN Notices, 6(9):42–49, October 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergeron:1971:LSD

[118] R. Daniel Bergeron, John D. Gannon, and Andries van Dam. Language for Systems Development. ACM SIGPLAN Notices, 6(9):50–72, October 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lyle:1971:HHO

[119] Don M. Lyle. A hierarchy of high order languages for systems programming. ACM SIGPLAN Notices, 6(9): 73–78, October 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1971:SLP

[120] B. L. Clark and J. J. Horning. The System Language for Project SUE. ACM SIGPLAN Notices, 6(9):79–88, October 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hopkins:1971:PPS

[121] Martin Hopkins. Problems of PL/I for system programming. ACM SIG-PLAN Notices, 6(9):89–91, October 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bois:1971:LLU

[122] Pierre J. Du Bois, Jr. and Jeanne T. Martin. The LRLTRAN language as used in the FROST and FLOE timesharing operating systems. ACM SIG-PLAN Notices, 6(9):92–104, October 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DuBois:1971:LLU

[123] P. J. Du Bois, Jr. and J. T. Martin. The LRLTRAN language as used in the FROST and FLOE time-sharing operating systems. ACM SIGPLAN Notices, 6(9):92–104, October 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bingham:1971:UAM

[124] Harvey W. Bingham. Use of APL in microprogrammable machine modeling. ACM SIGPLAN Notices, 6(9): 105–109, October 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiederhold:1971:ISS

[125] Gio Wiederhold and John Ehrman. Inferred syntax and semantics of PL/S. ACM SIGPLAN Notices, 6(9):110– 121, October 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kagan:1971:FPS

[126] Claude A. R. Kagan. Floating point sub-language of a string language. ACM SIGPLAN Notices, 6 (10):20–22, November 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1317449.

Korsan:1971:PAE

[127] Robert J. Korsan. A proposed APL extension. ACM SIGPLAN Notices, 6 (10):21–23, November 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1317454.

Glick:1971:AH

[128] Norman Glick and Richard Schrader. APL on the Honeywell 635. ACM SIG-PLAN Notices, 6(10):23–30, November 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.

acm.org/citation.cfm?id=1317455.

Page:1971:OLP

[129] P. D. Page. An old-line proof checker operating under APL/360, with educational applications in logic, mathematics, and computer science. ACM SIG-PLAN Notices, 6(10):30–34, November 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.

acm.org/citation.cfm?id=1317456.

Zaks:1971:LM

[130] Rodnay Zaks. A language machine. ACM SIGPLAN Notices, 6 (10):34–39, November 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL https://dl.acm.org/
citation.cfm?id=1317457.

Clementi:1971:MASb

[131] J. F. Clementi, B. G. James, and R. P. Fletcher. Modifications to the APL/ 1130 system part 2. ACM SIGPLAN Notices, 6(10):40–46, November 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1317458.

Ahl:1971:CBD

[132] Kenneth Ahl, John Barrington, John Hillier, Evelyn Mack, and Walter Whipple. Character and BIT data types for FORTRAN: a proposal to ANSI subcommittee X3J3. ACM SIG-PLAN Notices, 6(10):22–40, November 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.

acm.org/citation.cfm?id=1317450.

Perstein:1971:RSS

[133] Millard H. Perstein. Report to SPARC (the Standards Planning and Requirements Committee of ANSI) from the ad hoc committee on operating system control languages. ACM SIG-PLAN Notices, 6(10):41–49, November 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.

acm.org/citation.cfm?id=1317451.

Shaw:1971:FIBb

[134] Christopher J. Shaw. FORTRAN information bulletin. ACM SIGPLAN Notices, 6(10):50–65, November 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL https://dl.acm.

org/citation.cfm?id=1317452.

Kopetz:1971:IIS

[135] Hermann Kopetz. On the implementation of an interactive system on a small computer. ACM SIGPLAN Notices, 6(11):2–4, November 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kratky:1971:DA

[136] Gerhard Kratky. The definition of AM-TRAN. ACM SIGPLAN Notices, 6 (11):5–31, November 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reinfelds:1971:CCD

[137] Juris Reinfelds. A concept by concept description of the AMTRAN language. ACM SIGPLAN Notices, 6(11): 32–59, November 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reinfelds:1971:SAA

[138] Juris Reinfelds. Some applications of AMTRAN. ACM SIGPLAN Notices, 6(11):60–90, November 1971.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1971:SIA

[139] Anonymous. Special issue on AM-TRAN (Automatic Mathematical TRANslation). ACM SIGPLAN Notices, 6(11):??, November 1971. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Scowen:1971:BAE

[140] R. S. Scowen. Babel, an application of extensible compilers. ACM SIG-PLAN Notices, 6(12):1–7, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chandler:1971:MLE

[141] Gerald D. Chandler. METAPI: a language for extensions. ACM SIG-PLAN Notices, 6(12):8–9, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yoneda:1971:DSA

[142] N. Yoneda. The description and the structure of ALGOL N. ACM SIG-PLAN Notices, 6(12):10–14, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Suzuki:1971:IA

[143] N. Suzuki, T. Fujisaki, K. Kakehi, E. Wada, and T. Simauti. The implementation of ALGOL N. ACM SIG-PLAN Notices, 6(12):15–22, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Katzenelson:1971:SPM

[144] J. Katzenelson and E. Milgrom. A short presentation of the main features of AEPL — an extensible programming language. ACM SIGPLAN Notices, 6(12):23–25, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegbreit:1971:OEP

[145] Ben Wegbreit. An overview of the ECL programming system. ACM SIG-PLAN Notices, 6(12):26–28, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Notley:1971:MEL

[146] M. G. Notley. A model of extensible language systems. ACM SIGPLAN Notices, 6(12):29–38, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bowlden:1971:MHL

[147] Henry J. Bowlden. Macros in higherlevel languages. ACM SIGPLAN Notices, 6(12):39–44, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grant:1971:STC

[148] Charles A. Grant. Syntax translation with context macros or macros without arguments. ACM SIGPLAN Notices, 6(12):45–47, December 1971.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wodon:1971:SMP

[149] P. L. Wodon. A syntax macro processor. ACM SIGPLAN Notices, 6(12): 48–50, December 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fernandez:1971:OSP

[150] A. Fernandez, J. C. Heliard, and J. D. Ichbiah. Overview of the syntax processor generator SYNPROC. ACM SIG-PLAN Notices, 6(12):51–55, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lang:1971:PND

[151] Bernard Lang. Parallel non-deterministic bottom-up parsing. ACM SIGPLAN Notices, 6(12):56–57, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hammer:1971:AAM

[152] Michael Hammer. An alternative approach to macro processing. ACM SIG-PLAN Notices, 6(12):58–64, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reghizzi:1971:GID

[153] S. Crespi Reghizzi. Grammatical inference for defining extensions. ACM SIG-PLAN Notices, 6(12):65–66, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanGils:1971:SDM

[154] Teun van Gils. Syntactic definition mechanisms. ACM SIGPLAN Notices, 6(12):67–74, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jorrand:1971:DTE

[155] Philippe Jorrand. Data types and extensible languages. ACM SIGPLAN Notices, 6(12):75–83, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ichbiah:1971:ES

[156] Jean D. Ichbiah. Extensibility in Simula 67. ACM SIGPLAN Notices, 6(12): 84–86, December 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brosgol:1971:IED

[157] Benjamin M. Brosgol. An implementation of ECL data types. ACM SIG-PLAN Notices, 6(12):87–95, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leavenworth:1971:TFM

[158] B. M. Leavenworth. Transition functions: A method for semantic extensions. ACM SIGPLAN Notices, 6(12): 96–103, December 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prenner:1971:CSF

[159] Charles J. Prenner. The control structure facilities of ECL. ACM SIG-PLAN Notices, 6(12):104–112, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krutar:1971:CSP

[160] Rudolph A. Krutar. Conversational systems programming by incremental extension of system configuration. ACM SIGPLAN Notices, 6(12):113– 116, December 1971. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dickman:1971:SIP

[161] B. N. Dickman. Subroutine interface primitives for ETC. ACM SIG-PLAN Notices, 6(12):117–119, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schuman:1971:EI

[162] Stephen A. Schuman. An extensible interpreter. ACM SIGPLAN Notices, 6(12):120–128, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holloway:1971:ICI

[163] Glenn H. Holloway. Interpreter/ compiler integration in ECL. ACM SIGPLAN Notices, 6(12):129–134, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Woolley:1971:MSM

[164] John D. Woolley. MENELAUS: A system for measurement of the extensible language Proteus. ACM SIGPLAN Notices, 6(12):135–136, December 1971.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Duby:1971:ELP

[165] J. J. Duby. Extensible languages: A potential user’s point of view. ACM SIGPLAN Notices, 6(12):137–140, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sammet:1971:AEL

[166] Jean E. Sammet. Application of extensible languages to specialized application languages. ACM SIGPLAN Notices, 6(12):141–143, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Standish:1971:PEL

[167] Thomas A. Standish. PPL: an extensible language that failed. ACM SIG-PLAN Notices, 6(12):144–145, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheatham:1971:ELW

[168] T. E. Cheatham, Jr. Extensible language — where are we going. ACM SIGPLAN Notices, 6(12):146–147, December 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Milner:1972:IAS

[169] Robin Milner. Implementation and applications of Scott’s logic for computable functions. ACM SIGPLAN Notices, 7(1):1–6, January 1972. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burstall:1972:ADP

[170] Rod M. Burstall. An algebraic description of programs with assertions, verification and simulation. ACM SIGPLAN Notices, 7(1):7–14, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allen:1972:DAD

[171] C. David Allen. Derivation of axiomatic definitions of programming languages from algorithmic definitions. ACM SIGPLAN Notices, 7(1):15–26, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Manna:1972:IMP

[172] Zohar Manna, Stephen Ness, and Jean Vuillemin. Inductive methods for proving properties of programs. ACM SIG-PLAN Notices, 7(1):27–50, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ashcroft:1972:PCM

[173] E. A. Ashcroft. Program correctness methods and language definition. ACM SIGPLAN Notices, 7(1):51–57, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cadiou:1972:RDP

[174] J. M. Cadiou and Zohar Manna. Recursive definitions of partial functions and their computations. ACM SIG-PLAN Notices, 7(1):58–65, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hull:1972:CNA

[175] T. E. Hull, W. H. Enright, and A. E. Sedgwick. The correctness of numerical algorithms. ACM SIGPLAN Notices, 7 (1):66–73, January 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Paterson:1972:DPC

[176] Michael S. Paterson. Decision problems in computational models. ACM SIGPLAN Notices, 7(1):74–82, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garland:1972:TRS

[177] Stephen J. Garland and David C. Luckham. Translating recursion schemes into program schemes. ACM SIG-PLAN Notices, 7(1):83–96, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Strong:1972:PPU

[178] H. R. Strong and S. A. Walker. Properties preserved under recursion removal. ACM SIGPLAN Notices, 7(1):97–103, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fischer:1972:LCS

[179] Michael J. Fischer. Lambda calculus schemata. ACM SIGPLAN Notices, 7(1):104–109, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McGowan:1972:CML

[180] Clement McGowan. The Contour Model Lambda Calculus Machine. ACM SIGPLAN Notices, 7(1):110–115, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1972:BBD

[181] Raymond E. Miller. A boundary between decidability and undecidability for parallel program schemata (extended abstract). ACM SIGPLAN Notices, 7(1):116–120, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

London:1972:CCL

[182] Ralph L. London. Correctness of a compiler for a Lisp subset. ACM SIG-PLAN Notices, 7(1):121–127, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1972:OSP

[183] Peter Wegner. Operational semantics of programming languages. ACM SIG-PLAN Notices, 7(1):128–141, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1972:DVR

[184] John A. N. Lee. The definition and validation of the radix sorting technique. ACM SIGPLAN Notices, 7(1):142–149, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1972:FDC

[185] C. B. Jones. Formal development of correct algorithms: An example based on Earley’s recogniser. ACM SIG-PLAN Notices, 7(1):150–169, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1972:EMT

[186] Daniel M. Berry. The equivalence of models of tasking. ACM SIGPLAN Notices, 7(1):170–190, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McGowan:1972:MRE

[187] Clement L. McGowan. The “most recent” error: Its causes and correction. ACM SIGPLAN Notices, 7(1):191–202, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sintzoff:1972:CPP

[188] Michel Sintzoff. Calculating properties of programs by valuations on specific models. ACM SIGPLAN Notices, 7(1): 203–207, January 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheatham:1972:LSA

[189] T. E. Cheatham, Jr. and Ben Wegbreit. On a laboratory for the study of automating programming. ACM SIG-PLAN Notices, 7(1):208–211, January 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1972:ENI

[190] Christopher J. Shaw. Errata to 1971 November 10 issue. ACM SIG-PLAN Notices, 7(2):18, February 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1972:ESI

[191] Christopher J. Shaw. Errata to 1971 September issue. ACM SIG-PLAN Notices, 7(2):18, February 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Halstead:1972:NLC

[192] M. H. Halstead. Natural laws controlling algorithm structure? ACM SIG-PLAN Notices, 7(2):19–26, February 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snowdon:1972:PIS

[193] R. A. Snowdon. PEARL: an interactive system for the preparation and validation of structured programs. ACM SIGPLAN Notices, 7(3):9–26, March 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jenkins:1972:DAP

[194] M. A. Jenkins. DOMINO: an APL primitive function for matrix inversion — its implementation and applications. ACM SIGPLAN Notices, 7(4):29–40, April 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pettus:1972:IDC

[195] C. Pettus. Indeterminate 0 ÷ 0 check in APL. ACM SIGPLAN Notices, 7 (4):40–41, April 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smillie:1972:ASL

[196] K. W. Smillie. APLISP: a simple list processor in APL. ACM SIG-PLAN Notices, 7(4):41–45, April 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrams:1972:CAP

[197] Phillip S. Abrams and Lawrence M. Breed. COLLOQUE APL: Paris, September 1971. ACM SIGPLAN Notices, 7(4):45–47, April 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clementi:1972:MAS

[198] J. F. Clementi, R. P. Fletcher, and B. G. James. Modification to the APL/1130 system: part 3. ACM SIG-PLAN Notices, 7(4):47–49, April 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fletcher:1972:ANM

[199] Robert P. Fletcher. Algorithms: number 77 multiple integration. ACM SIG-PLAN Notices, 7(4):55–64, April 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dickman:1972:RSS

[200] B. N. Dickman. Review of “Symposium on some directions in high-level languages”. ACM SIGPLAN Notices, 7(5):4–8, May 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McGonagle:1972:NCP

[201] J. David McGonagle. Notes on the computer program test methods symposium. ACM SIGPLAN Notices, 7 (5):8–12, May 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wulf:1972:PRS

[202] W. A. Wulf. Project Rosetta Stone: an exhibition of languages for system implementation. ACM SIGPLAN Notices, 7(5):13, May 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mor:1972:PPL

[203] Leonard Mor, Joseph H. Yott, and Hans Zima. Progress: a programming language for real-time systems. ACM SIGPLAN Notices, 7(6):4–24, June 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lang:1972:NTD

[204] Bernard Lang. A new technique for data monitoring. ACM SIG-PLAN Notices, 7(6):41–42, June 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fletcher:1972:ALS

[205] J. G. Fletcher, C. S. Badger, G. L. Boer, and G. G. Marshall. On the appropriate language for system programming. ACM SIGPLAN Notices, 7(7):28–30, July 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wichmann:1972:EES

[206] B. A. Wichmann. Estimating the execution speed of an Algol program. ACM SIGPLAN Notices, 7(8):24–44, August 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krall:1972:ICL

[207] Edward J. Krall. An implementation of on-conditions in LISP. ACM SIGPLAN Notices, 7(9):13–20, 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schwartz:1972:DIP

[208] B. M. Schwartz. Design and implementation of PL/I preprocessor-based systems. ACM SIGPLAN Notices, 7(9): 21–36, 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:1972:RTD

[209] Mark B. Wells. A review of twodimensional programming languages. ACM SIGPLAN Notices, 7(10):1–10, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Klerer:1972:EPS

[210] Melvin Klerer. The economics, politics, and sociology of two-dimensional systems. ACM SIGPLAN Notices, 7(10): 11–22, October 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gawlik:1972:PM

[211] H. J. Gawlik. Progress in MIR-FAC. ACM SIGPLAN Notices, 7(10): 23–29, October 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

George:1972:GST

[212] James E. George. Graphical systems and two-dimensional mathematical expressions. ACM SIGPLAN Notices, 7(10):30–34, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lazik:1972:GMD

[213] George L. Lazik. GRAPHOS: A macro directed system for two-dimensional syntax translation. ACM SIGPLAN Notices, 7(10):35–47, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morrison:1972:SCP

[214] D. R. Morrison and R. C. Allen, Jr. Shape coefficients for polygons. ACM SIGPLAN Notices, 7(10):48–56, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kulsrud:1972:ESD

[215] H. E. Kulsrud. An extensible system for drawing block diagrams. ACM SIG-PLAN Notices, 7(10):57–65, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wolfberg:1972:FAL

[216] Michael S. Wolfberg. Fundamentals of the AMBIT/L list-processing language. ACM SIGPLAN Notices, 7(10): 66–75, October 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wright:1972:TDI

[217] W. V. Wright. The two-dimensional interface of an interactive system for molecular studies. ACM SIGPLAN Notices, 7(10):76–84, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holmes:1972:PGG

[218] H. H. Holmes and D. M. Austin. PI-CASSO a general graphics modeling program. ACM SIGPLAN Notices, 7 (10):85–92, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griesmer:1972:SCV

[219] James H. Griesmer and Richard D. Jenks. SCRATCHPAD: A capsule view. ACM SIGPLAN Notices, 7(10): 93–102, October 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Proceedings of the symposium on Twodimensional man-machine communication, Mark B. Wells and James B. Morris (eds.).

Knowlton:1972:RUF

[220] Ken Knowlton. A report on the use of FORTRAN-coded EXPLOR for the teaching of computer graphics and computer art. ACM SIGPLAN Notices, 7(10):103–112, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderson:1972:PTP

[221] Robert H. Anderson. Programming on a tablet: A proposal for a new notation. ACM SIGPLAN Notices, 7(10): 113–123, October 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saylor:1972:AMI

[222] Jean Saylor. The Assistant Mathematician: An interactive graphic programming language. ACM SIGPLAN Notices, 7(10):124–129, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Devaney:1972:TCL

[223] Marjorie J. Devaney and Jeanne Hudgins. The terminal control language for the Madcap programming language. ACM SIGPLAN Notices, 7(10):130– 136, October 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1972:TDC

[224] James S. Miller. Two-dimensional characteristics of HAL, a language for spaceflight applications. ACM SIG-PLAN Notices, 7(10):137–141, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haag:1972:TDU

[225] James N. Haag, Michael A. Kelly, and Paul F. Sherman. TWO-D: The USF Beginner’s Two-Dimensional Programming System. ACM SIGPLAN Notices, 7(10):142–152, October 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fisher:1972:SCS

[226] David A. Fisher. A survey of control structures in programming languages. ACM SIGPLAN Notices, 7 (11):1–13, November 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gostelow:1972:PTF

[227] Kim Gostelow, Vincent G. Cerf, Gerald Estrin, and Saul Volansky. Proper termination of flow-of-control in programs involving concurrent processes. ACM SIGPLAN Notices, 7(11):15– 27, November 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morris:1972:SPF

[228] James B. Morris and Mark B. Wells. The specification of program flow in Madcap 6. ACM SIGPLAN Notices, 7 (11):28–35, November 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prenner:1972:IBP

[229] Charles J. Prenner, Jay M. Spitzen, and Ben Wegbreit. An implementation of backtracking for programming languages. ACM SIGPLAN Notices, 7(11):36–44, November 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneck:1972:ARV

[230] Paul B. Schneck. Automatic recognition of vector and parallel operations in a higher level language. ACM SIG-PLAN Notices, 7(11):45–52, November 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leavenworth:1972:PG

[231] B. M. Leavenworth. Programming with(out) the GOTO. ACM SIGPLAN Notices, 7(11):54–58, November 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hopkins:1972:CG

[232] Martin E. Hopkins. A case for the GOTO. ACM SIGPLAN Notices, 7 (11):59–62, November 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wulf:1972:CAG

[233] William A. Wulf. A case against the GOTO. ACM SIGPLAN Notices, 7(11):63–69, November 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wulf:1972:PDS

[234] William A. Wulf. The problem of the definition of subroutine calling conventions. ACM SIGPLAN Notices, 7(12): 3–8, December 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dickman:1972:SIP

[235] B. N. Dickman. Subroutine interface primitives for ETC. ACM SIGPLAN Notices, 7(12):9–19, December 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baecker:1972:MMA

[236] H. D. Baecker. On a missing mode in ALGOL 68. ACM SIGPLAN Notices, 7(12):20–30, December 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rain:1972:SCM

[237] Mark Rain. Storage control in MARY. ACM SIGPLAN Notices, 7(12):31– 39, December 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1972:TSP

[238] Eberhard Wegner. Tree-structured programs. ACM SIGPLAN Notices, 7(12): 40–42, December 1972. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holager:1972:CSP

[239] Per Holager. A consolidation of synchronization primitives. ACM SIG-PLAN Notices, 7(12):42–52, December 1972. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rain:1973:OEM

[240] Mark Rain. Operation expressions in MARY. ACM SIGPLAN Notices, 8 (1):7–14, January 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haines:1973:ASA

[241] Edward C. Haines. AL: a structured assembly language. ACM SIG-PLAN Notices, 8(1):15–20, January 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fenner:1973:QDV

[242] T. I. Fenner, M. A. Jenkins, and R. D. Tennent. QUEST: the design of a very high level, pedagogic programming language. ACM SIGPLAN Notices, 8(2): 3–27, February 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wulf:1973:GVC

[243] W. Wulf and Mary Shaw. Global variable considered harmful. ACM SIG-PLAN Notices, 8(2):28–34, February 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1973:HCS

[244] Eberhard Wegner. A hierarchy of control structures. ACM SIGPLAN Notices, 8(3):11–17, March 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gurski:1973:JCL

[245] Aaron Gurski. Job control languages as machine orientated [sic] languages. ACM SIGPLAN Notices, 8(3):18–23, March 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rain:1973:PMR

[246] Mark Rain and Per Holager. The present most recent final word about labels in MARY. ACM SIGPLAN Notices, 8(3):24–32, March 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foster:1973:FAA

[247] Garth H. Foster. File action in APL: building new additions and placing the house in order. ACM SIG-PLAN Notices, 8(4):6–15, April 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haines:1973:OPP

[248] Edward C. Haines. Omitted portions of paper: AL: a structured assembly language. ACM SIGPLAN Notices, 8 (4):16–21, April 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holt:1973:TFD

[249] Richard C. Holt. Teaching the fatal disease: (or) introductory computer programming using PL/I. ACM SIG-PLAN Notices, 8(5):8–23, May 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dunn:1973:SLB

[250] Richard Dunn. SNOBOL4 as a language for bootstrapping a compiler. ACM SIGPLAN Notices, 8(5):28–32, May 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Apperson:1973:OEO

[251] Jerry L. Apperson. Optimal evaluation order for expressions with redundant subexpressions. ACM SIG-PLAN Notices, 8(6):1–3, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blois:1973:FPL

[252] M. S. Blois, R. R. Henley, and P. LeBeux. Frame programming language. ACM SIGPLAN Notices, 8 (6):4–6, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bochmann:1973:CGA

[253] G. V. Bochmann. Compiler generation for attribute grammars. ACM SIGPLAN Notices, 8(6):7, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brooks:1973:TIS

[254] Nancy B. Brooks. Tactics, an integrated system for structured programming. ACM SIGPLAN Notices, 8 (6):8–11, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bulut:1973:IPA

[255] Necdet Bulut and M. H. Halstead. Invariant properties of algorithms. ACM SIGPLAN Notices, 8(6):12–13, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buckle:1973:DLH

[256] N. Buckle. Design of a language for high-level data structures. ACM SIG-PLAN Notices, 8(6):14–15, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Castleman:1973:PS

[257] Paul A. Castleman. PROPHET system. ACM SIGPLAN Notices, 8(6): 16–18, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Christopher:1973:DSP

[258] Thomas W. Christopher. Development of the system programming language: HARVEY. ACM SIGPLAN Notices, 8(6):19–21, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1973:EPP

[259] R. O. Clark. Environmental primitives for programming languages. ACM SIG-PLAN Notices, 8(6):22–23, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Conklin:1973:DLC

[260] James W. Conklin. Design of a language for complex systems. ACM SIG-PLAN Notices, 8(6):24–25, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Daquin:1973:ESA

[261] C. Daquin, C. Giraud, B. De La Passardiere, and B. Quement. An extensible system for abstract machines. ACM SIGPLAN Notices, 8(6):25–27, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Davis:1973:SSP

[262] Thomas M. Davis. A system for the specification of programs. ACM SIG-PLAN Notices, 8(6):27–29, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dennis:1973:TSC

[263] J. B. Dennis. Toward the specification of a common base language. ACM SIG-PLAN Notices, 8(6):29–31, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

deRoever:1973:OMS

[264] W. P. de Roever. Operational and mathematical semantics for recursive polyadic program schemata. ACM SIGPLAN Notices, 8(6):31–34, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Desjardins:1973:PCX

[265] P. Desjardins. A Pascal compiler for the Xerox SIGMA 6. ACM SIG-PLAN Notices, 8(6):34–36, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dupras:1973:THL

[266] M. Dupras, M. Gauthier, and C. Girault. Transferability of high level languages. ACM SIGPLAN Notices, 8 (6):36–37, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fairley:1973:ISS

[267] Richard L. Fairley. An interactive system for studying semantic models of computer programs. ACM SIG-PLAN Notices, 8(6):37–39, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraley:1973:HSP

[268] Robert A. Fraley. High-speed parsing of a superset of LR(k) grammars. ACM SIGPLAN Notices, 8(6):39–41, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Friedman:1973:CCL

[269] H. George Friedman and Axel-Tobias Schreiner. CLEOPATRA: comprehensive language for elegant operating system and translator design. ACM SIG-PLAN Notices, 8(6):41–43, June 1973.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Funk:1973:DUU

[270] E. J. Funk. The design and use of understandable, minimally error-prone programming languages. ACM SIG-PLAN Notices, 8(6):43–46, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gannon:1973:LCD

[271] John D. Gannon. Language and compiler design to enhance reliability. ACM SIGPLAN Notices, 8(6):47–49, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jenkins:1973:CA

[272] M. A. Jenkins. A compiler for APL. ACM SIGPLAN Notices, 8(6):49–50, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jenkins:1973:DIQ

[273] M. A. Jenkins and R. D. Tennent. The design and implementation of quest, a very high level pedagogic programming language. ACM SIGPLAN Notices, 8 (6):50–51, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lecarme:1973:CWS

[274] Olivier Lecarme. Compiler writing system. ACM SIGPLAN Notices, 8 (6):51–54, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lechner:1973:SPL

[275] Robert Lechner and William Stallings. A systems programming language for minicomputers. ACM SIGPLAN Notices, 8(6):54–56, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ledgard:1973:PSC

[276] Henry F. Ledgard. Production systems — or can we do better than BNF? ACM SIGPLAN Notices, 8(6):56–58, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levison:1973:QPL

[277] M. Levison. QUILL: a programming language for literary text analysis. ACM SIGPLAN Notices, 8(6):58–60, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Madrigal:1973:CBS

[278] Orlando S. Madrigal. A compiler building system for small-scale computers. ACM SIGPLAN Notices, 8(6):60–62, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1973:EFS

[279] Edward F. Miller. Extensions to FOR-TRAN to support structured programming (ITRAN). ACM SIGPLAN Notices, 8(6):63–64, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morris:1973:MPL

[280] James B. Morris and Mark B. Wells. The Madcap 6 programming language. ACM SIGPLAN Notices, 8(6):64–66, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pfeffer:1973:DNF

[281] Lawrence Pfeffer. DAHLIA: a notation fabrication plan. ACM SIG-PLAN Notices, 8(6):66–70, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pfeffer:1973:PTD

[282] Lawrence Pfeffer. Permanent and transient data structures and associated optimization techniques. ACM SIG-PLAN Notices, 8(6):70–72, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pfeffer:1973:RNM

[283] Lawrence Pfeffer. RAMBAM: a notation and methodology for man-machine dialogue specification. ACM SIG-PLAN Notices, 8(6):72–74, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pratt:1973:FDP

[284] Terrence Pratt. Formal definition of programming languages. ACM SIG-PLAN Notices, 8(6):74–76, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Press:1973:MOD

[285] B. Press. Meta-optimizer development. ACM SIGPLAN Notices, 8(6):77–80, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reinfelds:1973:DIL

[286] Juris Reinfelds. Development of interactive languages for scientific computing: programming language research report. ACM SIGPLAN Notices, 8 (6):80–81, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Robinet:1973:SDS

[287] Bernard J. Robinet. Semantics of data structures. ACM SIGPLAN Notices, 8 (6):82–83, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1973:SSP

[288] Mary Shaw. A system for structured programming. ACM SIGPLAN Notices, 8(6):84–87, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1973:NST

[289] Ian D. Smith. A new set theory language. ACM SIGPLAN Notices, 8 (6):87–88, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1973:IPC

[290] Anonymous. Implementation of a Pascal compiler for the CII IRIS 80 computer. ACM SIGPLAN Notices, 8 (6):89–90, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Waite:1973:SEP

[291] W. M. Waite. Software engineering: portability. ACM SIGPLAN Notices, 8(6):90–92, June 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weissman:1973:PCC

[292] Larry Weissman. Psychological complexity of computer programs. ACM SIGPLAN Notices, 8(6):92–95, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wile:1973:GNS

[293] David S. Wile. A generative nestedsequential basis for general-purpose programming languages. ACM SIG-PLAN Notices, 8(6):96–98, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wortman:1973:PLG

[294] D. B. Wortman. Programming language grammar anthology. ACM SIG-PLAN Notices, 8(6):99–100, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1973:APL

[295] Anonymous. Abstracts in programming language-related research. ACM SIGPLAN Notices, 8(6):??, June 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marti:1973:ELR

[296] Jed B. Marti. ELMOL: a language for the real time generation of electronic music. ACM SIGPLAN Notices, 8(7):23–30, July 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1973:S

[297] S. Cohen. Speakeasy. ACM SIG-PLAN Notices, 8(7):31–44, July 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nassi:1973:FTS

[298] I. Nassi and B. Shneiderman. Flowchart techniques for structured programming. ACM SIGPLAN Notices, 8(8): 12–26, August 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1973:TSN

[299] S. Cohen. TSO-speakeasy newsletter #5. ACM SIGPLAN Notices, 8(8): 27–32, August 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dennis:1973:RSS

[300] Jack B. Dennis. Report of session on semantics. ACM SIGPLAN Notices, 8 (9):2–4, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freiburghouse:1973:RSSa

[301] Robert Freiburghouse and Barbara Liskov. Report of session on structured programming. ACM SIGPLAN Notices, 8(9):5–10, September 1973.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weller:1973:RST

[302] Maria F. Weller. Report of session on transferability. ACM SIGPLAN Notices, 8(9):11–16, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dennis:1973:RSC

[303] Jack Dennis and Michael D. Schroeder. Report of session on concurrency. ACM SIGPLAN Notices, 8(9):17–21, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freiburghouse:1973:RSSb

[304] Robert A. Freiburghouse and Robert M. Graham. Report of session on systems programming languages. ACM SIG-PLAN Notices, 8(9):22–26, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neumann:1973:RES

[305] Peter G. Neumann. Report of evening session protection. ACM SIGPLAN Notices, 8(9):27, September 1973. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levitt:1973:RES

[306] Karl Levitt. Report of evening session on hierarchical methods. ACM SIG-PLAN Notices, 8(9):28–29, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahams:1973:CWW

[307] Paul Abrahams. A compiler writer’s wishbook for operating systems. ACM SIGPLAN Notices, 8(9):30–33, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Austin:1973:SLC

[308] Joseph H. Austin, Jr. Specification languages for control programs. ACM SIGPLAN Notices, 8(9):34–37, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ballard:1973:SC

[309] Alan Ballard and Dennis Tsichritzis. System correctness. ACM SIGPLAN Notices, 8(9):38–41, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Balpaire:1973:SAC

[310] G. Balpaire and J. P. Wilmotte. Semantic aspects of concurrent processes. ACM SIGPLAN Notices, 8(9): 42–45, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Belpaire:1973:SAC

[311] Gérald Belpaire and Jean Pierre Wilmotte. Semantic aspects of concurrent processes. ACM SIGPLAN Notices, 8(9):42–45, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1973:TIO

[312] Daniel M. Berry. Towards ISMs for OPSs. ACM SIGPLAN Notices, 8(9): 46–49, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bochmann:1973:HLD

[313] Gregor V. Bochmann. Hierarchical language definition. ACM SIGPLAN Notices, 8(9):50–51, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1973:RLD

[314] B. L. Clark and J. J. Horning. Reflections on a language designed to write an operating system. ACM SIGPLAN Notices, 8(9):52–56, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Denil:1973:SDI

[315] N. J. Denil. Software design with invocation diagrams. ACM SIGPLAN Notices, 8(9):57–59, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Friedman:1973:SIL

[316] Frank L. Friedman and Victor B. Schneider. A systems implementation language for small computers. ACM SIGPLAN Notices, 8(9): 60–63, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Habermann:1973:ID

[317] A. N. Habermann. Integrated design. ACM SIGPLAN Notices, 8(9): 64–66, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henderson:1973:MMI

[318] D. Austin Henderson, Jr. A model of a modular interactive system. ACM SIG-PLAN Notices, 8(9):67–69, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herriot:1973:GSM

[319] Robert G. Herriot. GLOSS: A semantic model of programming languages. ACM SIGPLAN Notices, 8(9): 70–73, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holt:1973:SDI

[320] Richard C. Holt and Marc S. Grushcow. A short discussion of interprocess communication in the SUE/360/ 370 operating system. ACM SIGPLAN Notices, 8(9):74–78, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ichbiah:1973:TLA

[321] J. D. Ichbiah, J. P. Rissen, and J. C. Heliard. The two-level approach to data definition and space management in the LIS system implementation language. ACM SIGPLAN Notices, 8(9): 79–81, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jensen:1973:GPD

[322] Peter Jensen. The grok project data structures and process communication. ACM SIGPLAN Notices, 8(9): 82–85, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

King:1973:AMS

[323] J. C. King. Abstract machines and software design. ACM SIGPLAN Notices, 8(9):86–88, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kosinski:1973:DFL

[324] Paul R. Kosinski. A data flow language for operating systems programming. ACM SIGPLAN Notices, 8(9): 89–94, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lecarme:1973:ESP

[325] Olivier Lecarme. An experience in structured programming and transferability. ACM SIGPLAN Notices, 8(9): 95–96, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Linden:1973:PAP

[326] Theodore A. Linden. Proving the adequacy of protection in an operating system. ACM SIGPLAN Notices, 8(9): 97–99, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liskov:1973:SLC

[327] Barbara H. Liskov and Leroy A. Smith. SPIL: A language for construction of reliable system software. ACM SIG-PLAN Notices, 8(9):100–103, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLaren:1973:TSP

[328] M. Donald MacLaren. Tasking in standard PL/I. ACM SIGPLAN Notices, 8(9):104–108, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McKeag:1973:PLO

[329] R. M. McKeag. Programming languages for operating systems. ACM SIGPLAN Notices, 8(9):109–111, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Merten:1973:TTD

[330] Alan G. Merten and Edgar H. Sibley. Transferability and translation of data. ACM SIGPLAN Notices, 8(9):112– 114, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1973:PLF

[331] James S. Miller. A programming language family for the Navy AADC. ACM SIGPLAN Notices, 8(9):115– 116, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mills:1973:KSI

[332] Harlan D. Mills and Max L. Wilson. A Kernel System for information system development, evolution, and operation. ACM SIGPLAN Notices, 8(9):117–119, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morris:1973:PSR

[333] James B. Morris. Programming by semantic refinement. ACM SIGPLAN Notices, 8(9):120–122, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neuhold:1973:TFD

[334] E. J. Neuhold. Towards the formal description of operating systems. ACM SIGPLAN Notices, 8(9):123– 126, September 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Noonan:1973:SDO

[335] R. Noonan, V. Basili, R. Hamlet, M. Lay, D. Mills, J. Turner, and M. Zelkowitz. A SIMPL distributed operating system and its formal definition. ACM SIGPLAN Notices, 8 (9):127–128, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prenner:1973:ECS

[336] Charles J. Prenner. Extensible control structures. ACM SIGPLAN Notices, 8(9):129–132, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Riddle:1973:MDA

[337] William E. Riddle. A method for the description and analysis of complex software systems. ACM SIGPLAN Notices, 8(9):133–136, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Udin:1973:SSP

[338] David Udin. SPECL: A system programming dialect of ECL. ACM SIG-PLAN Notices, 8(9):137–139, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wiederhold:1973:NTO

[339] Gio C. M. Wiederhold. The need and techniques to obliterate control languages. ACM SIGPLAN Notices, 8 (9):140–141, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zilles:1973:PEL

[340] Stephen N. Zilles. Procedural encapsulation: A linguistic protection technique. ACM SIGPLAN Notices, 8(9): 142–146, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beech:1973:UPS

[341] David Beech and Michael Marcotty. Unfurling the PL/I standard. ACM SIGPLAN Notices, 8(10):12–43, October 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pagan:1973:TDP

[342] F. G. Pagan. On the teaching of disciplined programming. ACM SIG-PLAN Notices, 8(10):44–48, October 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Richards:1973:LSM

[343] Hamilton Richards and Roy J. Zingg. The logical structure of the memory resource in the symbol-2R computer. ACM SIGPLAN Notices, 8(11): 1–10, November 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anderberg:1973:HLL

[344] J. W. Anderberg and C. L. Smith. High-level language translation in symbol 2R. ACM SIGPLAN Notices, 8 (11):11–19, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hutchison:1973:PES

[345] P. C. Hutchison and K. Ethington. Program execution in the symbol 2R computer. ACM SIGPLAN Notices, 8(11): 20–26, November 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Richards:1973:ISP

[346] Hamilton Richards and Charles Wright. Introduction to the symbol 2R programming language. ACM SIGPLAN Notices, 8(11):27–33, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nielsen:1973:DAC

[347] William C. Nielsen. Design of an aerospace computer for direct HOL execution. ACM SIGPLAN Notices, 8 (11):34–42, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nissen:1973:AAD

[348] Stanley M. Nissen and Steven J. Wallach. The all applications digital computer. ACM SIGPLAN Notices, 8(11): 43–51, November 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1973:IAA

[349] James S. Miller and Woodrow H. Vandever. Instruction architecture of an aerospace multiprocessor. ACM SIG-PLAN Notices, 8(11):52–60, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bartlett:1973:HCS

[350] Joel F. Bartlett. The HP 3000 computer system. ACM SIGPLAN Notices, 8(11):61–69, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bloom:1973:SDH

[351] Howard M. Bloom. Structure of a direct high-level language processor. ACM SIGPLAN Notices, 8(11): 70–80, November 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herriot:1973:GHL

[352] Robert G. Herriot. GLOSS: a high level machine. ACM SIGPLAN Notices, 8(11):81–90, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McMahan:1973:ITA

[353] L. N. McMahan and E. A. Feustel. Implementation of a tagged architecture for block structured languages. ACM SIGPLAN Notices, 8(11):91– 100, November 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Welin:1973:IM

[354] Andrew M. Welin. The internal machine. ACM SIGPLAN Notices, 8(11): 101–108, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schroeder:1973:HOL

[355] S. C. Schroeder and L. E. Vaughn. A high order language optimal execution processor: \underline{F}ast \underline{I}ntent \underline{R}ecognition \underline{S}ys\underline{t}em (FIRST). ACM SIGPLAN Notices, 8(11):109–116, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schutte:1973:RVS

[356] Lawrence J. Schutte. A report on the value of some advanced high level language operators on current sequential computers. ACM SIGPLAN Notices, 8(11):117–123, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sitton:1973:PEL

[357] Gary A. Sitton, Thomas A. Kendrick, and A. Gil Carrick. The PL/EXUS language and virtual machine. ACM SIG-PLAN Notices, 8(11):124–130, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Haynes:1973:SPS

[358] Leonard S. Haynes. Structure of a Polish String language for an Algol 60 language processor. ACM SIGPLAN Notices, 8(11):131–140, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chirica:1973:TPE

[359] L. M. Chirica, T. A. Dreisbach, D. F. Martin, J. G. Peetz, and A. Sorkin. Two parallel EULER run time models: the dangling reference, imposter environment, and label problems. ACM SIGPLAN Notices, 8(11):141–151, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Basili:1973:HMM

[360] Victor R. Basili and Albert J. Turner. A hierarchical machine model for the semantics of programming languages. ACM SIGPLAN Notices, 8(11):152– 164, November 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1973:TRR

[361] D. M. Berry, L. Chirica, J. B. Johnston, D. F. Martin, and A. Sorkin. On the time required for retention. ACM SIG-PLAN Notices, 8(11):165–178, November 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martin:1973:NSB

[362] Johannes J. Martin. The ’natural’ set of basic control structures. ACM SIG-PLAN Notices, 8(12):5–14, December 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krutar:1973:RGV

[363] R. A. Krutar. Restricted global variables in Algol 60. ACM SIGPLAN Notices, 8(12):15–17, December 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

George:1973:VBP

[364] James E. George and Gary R. Sager. Variables: bindings and protection. ACM SIGPLAN Notices, 8(12):18– 29, December 1973. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ho:1974:RPI

[365] Thomas I. M. Ho. Review of “On programming: an interim report on the SETL project, installment II: the SETL language and examples of its use by J. T. Schwartz. New York University, 1973”. ACM SIGPLAN Notices, 9(1):6–7, January 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Friedman:1974:PLM

[366] Frank L. Friedman and Victor B. Schneider. A programming language for mini-computer systems. ACM SIGPLAN Notices, 9(1):8–32, January 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griswold:1974:SRA

[367] Ralph E. Griswold. Suggested revisions and additions to the syntax and control mechanisms of SNOBOL 4. ACM SIGPLAN Notices, 9(2):7–23, February 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Palme:1974:STE

[368] Jacob Palme. SIMULA as a tool for extensible program products. ACM SIG-PLAN Notices, 9(2):24–40, February 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bulut:1974:IFA

[369] Neodet Bulut and Maurice H. Halstead. Impurities found in algorithm implementations. ACM SIG-PLAN Notices, 9(3):9–12, March 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karpinski:1974:UVS

[370] Richard H. Karpinski. An unstructured view of structured programming. ACM SIGPLAN Notices, 9(3):12–20, March 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Backus:1974:FLP

[371] John Backus. Function level programs as mathematical objects. ACM SIG-PLAN Notices, 9(4):1–10, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leavenworth:1974:ONL

[372] B. M. Leavenworth and J. E. Sammet. An overview of nonprocedural languages. ACM SIGPLAN Notices, 9(4):1–12, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guttag:1974:FDA

[373] John Guttag, James Horning, and John Williams. FP with data abstraction and strong typing. ACM SIG-PLAN Notices, 9(4):11–24, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goldsmith:1974:DPP

[374] C. W. Goldsmith. The design of a procedureless programming language. ACM SIGPLAN Notices, 9(4):13–24, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wadler:1974:ASP

[375] Philip Wadler. Applicative style programming, program transformation, and list operators. ACM SIGPLAN Notices, 9(4):25–32, April 1974. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hammer:1974:IBD

[376] M. M. Hammer, W. G. Howe, and I. Wladawsky. An interactive business definition system. ACM SIG-PLAN Notices, 9(4):25–33, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Islam:1974:SOF

[377] N. Islam, T. J. Myers, and P. Broome. A simple optimizer for FP-like languages. ACM SIGPLAN Notices, 9 (4):33–40, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Earley:1974:HLO

[378] J. Earley. High level operations in automatic programming. ACM SIG-PLAN Notices, 9(4):34–42, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kieburtz:1974:TFP

[379] Richard B. Kieburtz and Jonathan Shultis. Transformations of FP program schemes. ACM SIGPLAN Notices, 9(4):41–48, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schwartz:1974:ASO

[380] J. T. Schwartz. Automatic and semiautomatic optimization of SETL. ACM SIGPLAN Notices, 9(4):43–49, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1974:PCD

[381] G. Berry. Programming with concrete data structures and sequential algorithms. ACM SIGPLAN Notices, 9 (4):49–58, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liskov:1974:PAD

[382] Barbara Liskov and Stephen Zilles. Programming with abstract data types. ACM SIGPLAN Notices, 9(4):50–59, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kapur:1974:OAS

[383] D. Kapur, D. R. Musser, and A. A. Stepanov. Operators and algebraic structures. ACM SIGPLAN Notices, 9 (4):59–64, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lee:1974:SPA

[384] R. C. T. Lee and S. K. Chang. Structured programming and automatic program synthesis. ACM SIGPLAN Notices, 9(4):60–70, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Darlington:1974:AMP

[385] John Darlington and Mike Reeve. AL-ICE: a multi-processor reduction machine for the parallel evaluation CF applicative languages. ACM SIG-PLAN Notices, 9(4):65–76, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shneiderman:1974:SDS

[386] B. Shneiderman and P. Scheuermann. Structured data structures. ACM SIG-PLAN Notices, 9(4):71–72, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fairley:1974:CSS

[387] R. E. Fairley. Continuous system simulation languages: design principles and implementation techniques. ACM SIG-PLAN Notices, 9(4):73–81, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Catto:1974:RMD

[388] A. J. Catto and J. R. Gurd. Resource management in dataflow. ACM SIG-PLAN Notices, 9(4):77–84, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tam:1974:PIE

[389] Wing Cheung Tam and W. J. Karplus. PDEL-ID: an extension of PDEL for distributed parameter system identification. ACM SIGPLAN Notices, 9 (4):82–90, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turner:1974:SEA

[390] D. A. Turner. The semantic elegance of applicative languages. ACM SIG-PLAN Notices, 9(4):85–92, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heidorn:1974:EVH

[391] G. E. Heidorn. English as a very high level language for simulation programming. ACM SIGPLAN Notices, 9(4): 91–100, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mago:1974:COV

[392] Gyula Magó. Copying operands versus copying results: A solution to the problem of large operands in FFP’S. ACM SIGPLAN Notices, 9(4):93–98, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stanat:1974:OAS

[393] Donald F. Stanat and E. Hollins Williams, Jr. Optimal associative searching on a cellular computer. ACM SIGPLAN Notices, 9(4):99–106, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jenks:1974:SL

[394] R. D. Jenks. The SCRATCHPAD language. ACM SIGPLAN Notices, 9(4): 101–111, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Presnell:1974:CAS

[395] H. A. Presnell and R. P. Pargas. Communication along shortest paths in a tree machine. ACM SIGPLAN Notices, 9(4):107–114, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levison:1974:DIV

[396] M. Levison. The design and implementation of a very high level language for literary scholars. ACM SIGPLAN Notices, 9(4):112–117, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tolle:1974:IFT

[397] Donald MacDavid Tolle. Implanting FFP trees in binary trees: An architectural proposal. ACM SIG-PLAN Notices, 9(4):115–122, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1974:SES

[398] S. Cohen. Speakeasy an evolutionary system. ACM SIGPLAN Notices, 9 (4):118–126, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Keller:1974:AFF

[399] Robert M. Keller and Gary Lindstrom. Applications of feedback in functional programming. ACM SIG-PLAN Notices, 9(4):123–130, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ziegler:1974:IUS

[400] E. W. Ziegler. An introduction to the UMTA specification language. ACM SIGPLAN Notices, 9(4):127–132, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Keller:1974:ACP

[401] Robert M. Keller and M. Ronan Sleep. Applicative caching: Programmer control of object sharing and lifetime in. distributed implementations of applicative languages. ACM SIGPLAN Notices, 9(4):131–140, April 1974. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nordstrom:1974:PCS

[402] Bengt Nordström. Programming in Constructive Set Theory: Some examples. ACM SIGPLAN Notices, 9(4): 141–154, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sintzoff:1974:POA

[403] Michel Sintzoff. Proof-oriented and applicative valuations in definitions of algorithms. ACM SIGPLAN Notices, 9 (4):155–162, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Conery:1974:PIL

[404] John S. Conery and Dennis F. Kibler. Parallel interpretation of logic programs. ACM SIGPLAN Notices, 9 (4):163–170, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1974:RLP

[405] Keith L. Clark and Steve Gregory. A relational language for parallel programming. ACM SIGPLAN Notices, 9 (4):171–178, April 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Page:1974:ITE

[406] R. L. Page, M. G. Conant, and D. H. Grit. If-then-else as a concurrency inhibitor in eager beaver evaluation of recursive programs. ACM SIG-PLAN Notices, 9(4):179–186, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burton:1974:EFP

[407] F. Warren Burton and M. Ronan Sleep. Executing functional programs on a virtual tree of processors. ACM SIG-PLAN Notices, 9(4):187–194, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buneman:1974:PFP

[408] Peter Buneman, Rishiyur Nikhil, and Robert Frankel. A practical functional programming system for databases. ACM SIGPLAN Notices, 9(4):195–202, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malhotra:1974:DFL

[409] V. M. Malhotra and V. Rajaraman. A data-flow language for specifying business data processing applications. ACM SIGPLAN Notices, 9(4):203–212, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLennan:1974:IRP

[410] B. J. MacLennan. Introduction to relational programming. ACM SIG-PLAN Notices, 9(4):213–220, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frank:1974:SDS

[411] Geoffrey A. Frank. Specification of data structures for FP programs. ACM SIGPLAN Notices, 9(4):221–228, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahams:1974:ICS

[412] Paul W. Abrahams. Improving the control structure of SNOBOL4. ACM SIG-PLAN Notices, 9(5):10–12, May 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bilofsky:1974:SEI

[413] Walter Bilofsky. Syntax extension and the IMP72 programming language. ACM SIGPLAN Notices, 9(5):13–30, May 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gileadi:1974:PMP

[414] Amos N. Gileadi and Henry F. Ledgard. On a proposed measure of program structure. ACM SIG-PLAN Notices, 9(5):31–36, May 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Besel:1974:AGD

[415] Hilmer W. Besel. After the GOTO debate. ACM SIGPLAN Notices, 9 (6):13–16, June 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dickman:1974:MFG

[416] B. N. Dickman. More fuel for the GOTO controversy. ACM SIGPLAN Notices, 9(6):17, June 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ostapko:1974:DRB

[417] Daniel L. Ostapko. On deriving a relation between circuits and input/output by analyzing an equivalent program. ACM SIGPLAN Notices, 9(6):18–24, June 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weissman:1974:PCC

[418] Larry Weissman. Psychological complexity of computer programs: an experimental methodology. ACM SIG-PLAN Notices, 9(6):25–36, June 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Peterson:1974:REP

[419] James Peterson. Review of The elements of programming style by Brian W. Kernighan and P. J. Plauger. McGraw-Hill Company, 1974. ACM SIGPLAN Notices, 9(7):6, July 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Friedman:1974:CWU

[420] Daniel P. Friedman and Stuart C. Shapiro. A case for while-until. ACM SIGPLAN Notices, 9(7):7–14, July 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lecarme:1974:SPP

[421] Olivier Lecarme. Structured programming, programming teaching and the language Pascal. ACM SIG-PLAN Notices, 9(7):15–21, July 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sanfield:1974:SVC

[422] Stuart H. Sanfield. The scope of variable concept: the key to structured programming? ACM SIGPLAN Notices, 9(7):22–29, July 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vonPeschke:1974:SRE

[423] Joachim von Peschke. Syntactic rules embedded in PL/I. ACM SIGPLAN Notices, 9(7):30–38, July 1974. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vPeschke:1974:SRE

[424] J. v. Peschke. Syntactic rules embedded in PL/I (translator writing system). ACM SIGPLAN Notices, 9(7):30–38, July 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Peck:1974:RAS

[425] J. E. L. Peck, J. M. Watt, and M. Sintzoff. Revised ALGOL 68 syntax chart. ACM SIGPLAN Notices, 9(7):39, July 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1974:BR

[426] Anonymous. Book reviews. ACM SIGPLAN Notices, 9(7):??, July 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reigel:1974:PLM

[427] E. W. Reigel and H. W. Lawson. At the Programming Language — Microprogramming interface. ACM SIGPLAN Notices, 9(8):2–22, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lesser:1974:DEP

[428] Victor R. Lesser. The design of an emulator for a parallel machine language. ACM SIGPLAN Notices, 9(8):23–36, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rossmann:1974:FMB

[429] George E. Rossmann and Louise H. Jones. Functional memory-based dynamic microprocessors for higher level languages. ACM SIGPLAN Notices, 9 (8):37–65, August 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tirrell:1974:SAC

[430] A. Keith Tirrell. A study of the application of compiler techniques to the generation of micro-code. ACM SIGPLAN Notices, 9(8):67–85, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lawson:1974:DFM

[431] Harold W. Lawson, Jr. and Lars Blomberg. The Datasaab FCPU microprogramming language. ACM SIG-PLAN Notices, 9(8):86–97, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rauscher:1974:TSS

[432] Tomlinson G. Rauscher. Towards a specification of syntax and semantics for languages for horizontally microprogrammed machines. ACM SIG-PLAN Notices, 9(8):98–111, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oestreicher:1974:MLM

[433] Donald R. Oestreicher. A microprogramming language for the MLP-900. ACM SIGPLAN Notices, 9(8):113–120, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stabler:1974:MSG

[434] George M. Stabler, Ingrid Carlbom, and Ken Magel. A microprogrammed satellite graphics system. ACM SIG-PLAN Notices, 9(8):121–129, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zaks:1974:DMM

[435] Rodnay Zaks. Dynamic memory management for APL-like languages. ACM SIGPLAN Notices, 9(8):130–138, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chevance:1974:CM

[436] R. J. Chevance. A COBOL machine. ACM SIGPLAN Notices, 9(8):139–144, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Broca:1974:DME

[437] François Robert Broca and Richard E. Merwin. Direct microprogrammed execution of the intermediate text from a high-level language compiler. ACM SIGPLAN Notices, 9(8):145–153, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Broadbent:1974:MME

[438] J. K. Broadbent and G. F. Coulouris. MEMBERS: a Microprogrammed Experimental Machine with a Basic Executive for Real-time Systems. ACM SIGPLAN Notices, 9(8):154–160, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Harrison:1974:LOI

[439] Malcolm C. Harrison. A languageoriented instruction set for the BALM language. ACM SIGPLAN Notices, 9 (8):161–168, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wade:1974:GPH

[440] Bradford W. Wade and Victor B. Schneider. A general-purpose highlevel language machine for minicomputers. ACM SIGPLAN Notices, 9(8):

169–177, August 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bray:1974:DSR

[441] David W. Bray. Dynamic storage routines for Fortran programs. ACM SIG-PLAN Notices, 9(9):2–13, September 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kleine:1974:MPD

[442] Henry Kleine and Richard V. Morris. Modern programming: A definition. ACM SIGPLAN Notices, 9(9): 14–17, September 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Middleton:1974:SMP

[443] Anthony G. Middleton. A structured model of programs suitable for analysing time/storage trade-offs. ACM SIGPLAN Notices, 9(9):18–28, September 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zelkowitz:1974:IP

[444] Marvin V. Zelkowitz. Interactive PL/1. ACM SIGPLAN Notices, 9(9): 29–31, September 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gimpel:1974:LES

[445] James F. Gimpel. Language extensions to SNOBOL4. ACM SIGPLAN Notices, 9(10):11–20, October 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gimpel:1974:SHS

[446] J. F. Gimpel. Some highlights of the SITBOL language extensions of SNOBOL 4. ACM SIGPLAN Notices, 9(10):11–20, October 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lecarme:1974:RPH

[447] Olivier Lecarme and Pierre Desjardins. Reply to a paper by A. N. Habermann on the programming language Pascal. ACM SIGPLAN Notices, 9(10): 21–27, October 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCoy:1974:TRC

[448] Kevin C. McCoy. Testing the readability and composability of computer programs: methodology. ACM SIG-PLAN Notices, 9(10):28, October 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meissner:1974:CSE

[449] Loren P. Meissner. A compatible “structured” extension to Fortran. ACM SIGPLAN Notices, 9(10): 29–36, October 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Denning:1974:SPL

[450] Peter J. Denning. Is ‘structured programming’ any longer the right term? ACM SIGPLAN Notices, 9(11): 4–6, November 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosen:1974:NSO

[451] Barry K. Rosen. Note on semantics and optimization. ACM SIGPLAN Notices, 9(11):6–10, November 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sammet:1974:RPL

[452] Jean E. Sammet. Roster of programming languages for 1973. ACM SIG-PLAN Notices, 9(11):18–31, November 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vaughan:1974:ALC

[453] W. C. M. Vaughan. Another look at the case statement. ACM SIGPLAN Notices, 9(11):32–36, November 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Korfhage:1974:DDS

[454] Robert R. Korfhage. On the development of data structures. ACM SIG-PLAN Notices, 9(12):14–22, December 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ONeill:1974:SPI

[455] Dennis M. O’Neill. SFOR — a precompiler for the implementation of a FORTRAN-based structured language. ACM SIGPLAN Notices, 9(12):22– 29, December 1974. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shneiderman:1974:CCS

[456] Ben Shneiderman. The chemistry of control structures. ACM SIGPLAN Notices, 9(12):29–34, December 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Terashima:1974:SSD

[457] Nobuyoshi Terashima. SYSL: system description language. ACM SIGPLAN Notices, 9(12):35–44, December 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marcus:1975:TCI

[458] Richard S. Marcus. A translating computer interface for a network of heterogeneous interactive information retrieval systems. ACM SIGPLAN Notices, 10(1):2–12, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Winograd:1975:BCB

[459] Terry Winograd. Breaking the complexity barrier again. ACM SIGPLAN Notices, 10(1):13–30, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boyce:1975:SQR

[460] R. F. Boyce, D. D. Chamberlin, M. M. Hammer, and W. F. King. Specifying queries as relational expressions. ACM SIGPLAN Notices, 10(1):31–47, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salton:1975:CEV

[461] G. Salton and Clement T. Yu. On the construction of effective vocabularies for information retrieval. ACM SIG-PLAN Notices, 10(1):48–60, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Waksman:1975:IRQ

[462] Abraham Waksman. Information retrieval and the query language. ACM SIGPLAN Notices, 10(1):61–70, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Artandi:1975:LIL

[463] Susan Artandi. The on-line interface of library management and document retrieval systems. ACM SIG-PLAN Notices, 10(1):71–74, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martin:1975:CAI

[464] Thomas H. Martin and Edwin B. Parker. Comparative analysis of interactive retrieval systems. ACM SIG-PLAN Notices, 10(1):75–85, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Martin:1975:LDM

[465] Thomas H. Martin and Richard L. Guertin. Language decisions made while designing an interactive information retrieval system. ACM SIGPLAN Notices, 10(1):86–100, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Claybrook:1975:FLI

[466] Billy G. Claybrook. FOL: a language for implementing file organizations for information storage and retrieval systems. ACM SIGPLAN Notices, 10(1): 101–112, January 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crouch:1975:ELI

[467] C. J. Crouch. An evaluation of languages for the implementation of information storage and retrieval systems. ACM SIGPLAN Notices, 10(1): 113–117, January 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1975:DDP

[468] Diane C. Pirog Smith. From a data description point of view. ACM SIG-PLAN Notices, 10(1):118–130, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1975:DSD

[469] W. O. Clark and L. L. Shannon. A dual structured data base system. ACM SIGPLAN Notices, 10(1):131–143, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Su:1975:ROD

[470] Stanley Y. W. Su, George P. Copeland, Jr., and G. Jack Lipovski. Retrieval operations and data representations in a context-addressed disc system. ACM SIGPLAN Notices, 10(1):144–160, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

King:1975:AIR

[471] P. F. King and J. E. Shemer. ARS: an interactive reporting system. ACM SIGPLAN Notices, 10(1):161–175, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dockery:1975:RTC

[472] John T. Dockery. Remarks on tactical computer problems. ACM SIG-PLAN Notices, 10(1):176–177, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mitchell:1975:NAP

[473] Patrick C. Mitchell. A note about the proximity operators in information retrieval. ACM SIGPLAN Notices, 10(1): 177–180, January 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Price:1975:CER

[474] Norman Price. Comments on experiments with retrieval of legal information in full text. ACM SIGPLAN Notices, 10(1):180–181, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beckman:1975:SE

[475] Anders Beckman. Secondary effects. ACM SIGPLAN Notices, 10(2):10– 11, February 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Geller:1975:HBC

[476] Dennis P. Geller. How bad can it get? (programming) by a large number of master’s candidates. ACM SIG-PLAN Notices, 10(2):12–13, February 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hill:1975:WAA

[477] I. D. Hill, R. S. Scowen, and B. A. Wichmann. Writing algorithms in AL-GOL 60. ACM SIGPLAN Notices, 10(2):14–31, February 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Plum:1975:MOS

[478] Thomas W.-S. S. Plum. Mathematical overkill and the structure theorem. ACM SIGPLAN Notices, 10(2): 32–33, February 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1975:CCP

[479] Eberhard Wegner. Control constructs for programming languages. ACM SIG-PLAN Notices, 10(2):34–41, February 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Higgins:1975:SFT

[480] Donald S. Higgins. A structured FOR-TRAN translator. ACM SIGPLAN Notices, 10(2):42–48, February 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Millstein:1975:IIF

[481] Robert E. Millstein and Charles A. Muntz. The ILLIAC IV FORTRAN compiler. ACM SIGPLAN Notices, 10 (3):1–8, March 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Presberg:1975:PIP

[482] David L. Presberg and Neil W. Johnson. The Paralyzer: Ivtran’s parallelism analyzer and synthesizer. ACM SIGPLAN Notices, 10(3):9–16, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Erickson:1975:APA

[483] David B. Erickson. Array processing on an array processor. ACM SIGPLAN Notices, 10(3):17–24, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lamport:1975:PPC

[484] Leslie Lamport. On programming parallel computers. ACM SIGPLAN Notices, 10(3):25–33, March 1975. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Geller:1975:PSR

[485] Dennis P. Geller and Gerald M. Weinberg. The principle of sufficient reason: A guide to language design for parallel processing. ACM SIG-PLAN Notices, 10(3):34–38, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Basili:1975:LDV

[486] V. R. Basili and J. C. Knight. A language design for vector machines. ACM SIGPLAN Notices, 10(3):39–43, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barta:1975:PPC

[487] Ben Zion Barta. PACOL: A parallel control language. ACM SIG-PLAN Notices, 10(3):44–53, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Resnick:1975:DCE

[488] Harold K. Resnick and Arvid G. Larson. DMAP: a COBOL extension for associative array processors. ACM SIGPLAN Notices, 10(3):54–61, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zosel:1975:MPV

[489] Mary E. Zosel. A modest proposal for vector extensions to ALGOL. ACM SIGPLAN Notices, 10(3):62–71, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stevens:1975:CFL

[490] K. G. Stevens, Jr. CFD — A Fortranlike language for the Illiac IV. ACM SIGPLAN Notices, 10(3):72–76, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zwakenberg:1975:VEL

[491] R. G. Zwakenberg. Vector extensions to LRLTRAN. ACM SIGPLAN Notices, 10(3):77–86, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Neel:1975:RIS

[492] D. Neel and M. Amirchahy. Removal of invariant statements from nested-loops in a single effective compiler pass. ACM SIGPLAN Notices, 10(3):87–96, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Loveman:1975:POT

[493] David B. Loveman and Ross A. Faneuf. Program optimization — theory and practice. ACM SIGPLAN Notices, 10 (3):97–102, March 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneck:1975:MIP

[494] Paul B. Schneck. Movement of implicit parallel and vector expressions out of program loops. ACM SIG-PLAN Notices, 10(3):103–106, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morris:1975:SPN

[495] D. Morris and P. C. Treleaven. A stream processing network. ACM SIG-PLAN Notices, 10(3):107–112, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hext:1975:ARO

[496] J. B. Hext. Array reference operations. ACM SIGPLAN Notices, 10(3): 113–118, March 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wedel:1975:FTI

[497] Dorothy Wedel. Fortran for the Texas Instruments ASC system. ACM SIG-PLAN Notices, 10(3):119–132, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Capon:1975:AOM

[498] P. C. Capon and R. N. Ibbett. Array operations in MU5. ACM SIG-PLAN Notices, 10(3):133–137, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Donegan:1975:LAP

[499] Michael K. Donegan and Stuart W. Katzke. Lexical analysis and parsing techniques for a vector machine. ACM SIGPLAN Notices, 10(3):138– 145, March 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Krohn:1975:PAC

[500] Howard E. Krohn. A parallel approach to code generation for Fortran like compilers. ACM SIGPLAN Notices, 10(3): 146–152, March 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Evans:1975:PFU

[501] D. J. Evans. Problem formulation using array processing techniques. ACM SIGPLAN Notices, 10(3):153– 163, March 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahams:1975:SPC

[502] Paul Abrahams. ‘structured programming’ considered harmful. ACM SIGPLAN Notices, 10(4):13–24, April 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McClure:1975:SPC

[503] Carma L. McClure. Structured programming in COBOL. ACM SIGPLAN Notices, 10(4):25–33, April 1975. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mathis:1975:FTS

[504] Robert F. Mathis. Flow trace of a structured program. ACM SIGPLAN Notices, 10(4):33–37, April 1975. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weiderman:1975:FLC

[505] N. H. Weiderman and B. M. Rawson. Flowcharting loops without cycles. ACM SIGPLAN Notices, 10(4):37–46, April 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wolman:1975:RIP

[506] Barry L. Wolman. Reply to ‘Interactive PL/I’. ACM SIGPLAN Notices, 10(4):46–48, April 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kamnitzer:1975:BXI

[507] Stephen H. Kamnitzer. Bootstrapping XPL from IBM/360 to UNIVAC 1100. ACM SIGPLAN Notices, 10(5):14–20, May 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McQuillan:1975:TDD

[508] David McQuillan. Transaction diagrams — a design tool. ACM SIG-PLAN Notices, 10(5):21–26, May 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parnas:1975:NFR

[509] David L. Parnas, John E. Shore, and W. David Elliott. On the need for fewer restrictions in changing compile-time environments. ACM SIGPLAN Notices, 10(5):29–36, May 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ripley:1975:TMS

[510] G. David Ripley and Ralph E. Griswold. Tools for the measurement of SNOBOL4 programs. ACM SIG-PLAN Notices, 10(5):36–52, May 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Davis:1975:SRG

[511] Ruth M. Davis. The software reliability game. ACM SIGPLAN Notices, 10(6): 1, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1975:GCN

[512] Edsger W. Dijkstra. Guarded commands, non-determinacy and a calculus for the derivation of programs. ACM SIGPLAN Notices, 10(6):2–2.13, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Williams:1975:MDR

[513] R. D. Williams. Managing the development of reliable software. ACM SIG-PLAN Notices, 10(6):3–8, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vyssotsky:1975:LSR

[514] Victor Vyssotsky. Large-scale reliable software: Recent experience at Bell Labs. ACM SIGPLAN Notices, 10(6): 9, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gannon:1975:ILD

[515] J. D. Gannon and J. J. Horning. The impact of language design on the production of reliable software. ACM SIG-PLAN Notices, 10(6):10–22, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirth:1975:APL

[516] Niklaus Wirth. An assessment of the programming language PASCAL. ACM SIGPLAN Notices, 10(6):23–30, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Geschke:1975:PUR

[517] C. M. Geschke and J. G. Mitchell. On the problem of uniform references to data structures. ACM SIG-PLAN Notices, 10(6):31–42, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cicu:1975:OTD

[518] A. Cicu, M. Maiocchi, R. Polillo, and A. Sardoni. Organizing tests during software evolution. ACM SIG-PLAN Notices, 10(6):43–50, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1975:AGT

[519] E. F. Miller, Jr. and R. A. Melton. Automated generation of testcase datasets. ACM SIGPLAN Notices, 10 (6):51–58, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stucki:1975:NAC

[520] Leon G. Stucki and Gary L. Foshee. New assertion concepts for selfmetric software validation. ACM SIG-PLAN Notices, 10(6):59–71, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liskov:1975:STD

[521] Barbara Liskov and Stephen Zilles. Specification techniques for data abstractions. ACM SIGPLAN Notices, 10 (6):72–87, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gerhart:1975:KAP

[522] Susan L. Gerhart. Knowledge about programs: A model and case study. ACM SIGPLAN Notices, 10(6):88–95, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chirica:1975:ACC

[523] Laurian M. Chirica and David F. Martin. An approach to compiler correctness. ACM SIGPLAN Notices, 10 (6):96–103, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rault:1975:DVT

[524] J. C. Rault. Design verification techniques — a review. ACM SIG-PLAN Notices, 10(6):104, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boehm:1975:SEA

[525] B. W. Boehm, R. K. McClean, and D. B. Urfrig. Some experience with automated aids to the design of largescale reliable software. ACM SIG-PLAN Notices, 10(6):105–113, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeRemer:1975:PLV

[526] Frank DeRemer and Hans Kron. Programming-in-the large versus programmingin-the-small. ACM SIGPLAN Notices, 10(6):114–121, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wulf:1975:RHS

[527] William A. Wulf. Reliable hardwaresoftware architecture. ACM SIG-PLAN Notices, 10(6):122–130, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reifer:1975:AAR

[528] Donald J. Reifer. Automated aids for reliable software. ACM SIG-PLAN Notices, 10(6):131–142, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Katz:1975:TAD

[529] Shmuel Katz and Zohar Manna. Towards automatic debugging of programs. ACM SIGPLAN Notices, 10 (6):143–155, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vonHenke:1975:MVP

[530] F. W. von Henke and D. C. Luckham. A methodology for verifying programs. ACM SIGPLAN Notices, 10 (6):156–164, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Caplain:1975:FIA

[531] Michel Caplain. Finding Invariant assertions for proving programs. ACM SIGPLAN Notices, 10(6):165– 171, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1975:SPP

[532] F. T. Baker. Structured programming in a production programming environment. ACM SIGPLAN Notices, 10(6): 172–185, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culpepper:1975:SRE

[533] L. M. Culpepper. A system for reliable engineering software. ACM SIG-PLAN Notices, 10(6):186–192, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brooks:1975:MMM

[534] Fred P. Brooks, Jr. The Mythical Man-Month. ACM SIGPLAN Notices, 10(6):193, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miyamoto:1975:SRO

[535] Isao Miyamoto. “software reliability in online real time environment”. ACM SIGPLAN Notices, 10(6):194– 203, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Littlewood:1975:RMM

[536] B. Littlewood. A reliability model for Markov structured software. ACM SIGPLAN Notices, 10(6):204–207, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Trivedi:1975:MSM

[537] Ashok K. Trivedi and Martin L. Shooman. A many-state Markov model for the estimation and prediction of computer software performance parameters. ACM SIGPLAN Notices, 10(6): 208–220, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Henderson:1975:FSM

[538] Peter Henderson. Finite state modelling in program development. ACM SIGPLAN Notices, 10(6):221–227, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

King:1975:NAP

[539] James C. King. A new approach to program testing. ACM SIGPLAN Notices, 10(6):228–233, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boyer:1975:SFS

[540] Robert S. Boyer, Bernard Elspas, and Karl N. Levitt. SELECT: a formal system for testing and debugging programs by symbolic execution. ACM SIGPLAN Notices, 10(6): 234–245, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rubey:1975:QAS

[541] Raymond J. Rubey. Quantitative aspects of software validation. ACM SIG-PLAN Notices, 10(6):246–251, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lynch:1975:REC

[542] W. C. Lynch, J. W. Langner, and M. S. Schwartz. Reliability experience with Chi/OS. ACM SIGPLAN Notices, 10 (6):252–259, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bucher:1975:MCS

[543] D. E. W. Bucher. Maintenance of the computer sciences teleprocessing system. ACM SIGPLAN Notices, 10(6): 260–266, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Robinson:1975:ARS

[544] Lawrence Robinson, Karl N. Levitt, Peter G. Neumann, and Ashok R. Saxena. On attaining reliable software for a secure operating system. ACM SIG-PLAN Notices, 10(6):267–284, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Walter:1975:SSS

[545] K. G. Walter, S. I. Schaen, W. F. Ogden, W. C. Rounds, D. G. Shumway, D. D. Schaeffer, K. J. Biba, F. T. Bradshaw, S. R. Ames, and J. M. Gilligan. Structured specification of a Security Kernel. ACM SIGPLAN Notices, 10 (6):285–293, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Popek:1975:VPS

[546] Gerald J. Popek and Charles S. Kline. A verifiable protection system. ACM SIGPLAN Notices, 10(6): 294–304, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

BrinchHansen:1975:PCP

[547] P. Brinch Hansen. The purpose of Concurrent PASCAL. ACM SIG-PLAN Notices, 10(6):305–309, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1975:PCP

[548] Per Brinch Hansen. The purpose of concurrent Pascal. ACM SIG-PLAN Notices, 10(6):305–309, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Saxena:1975:SSH

[549] Ashok R. Saxena and Thomas H. Bredt. A structured specification of a hierarchical operating system. ACM SIGPLAN Notices, 10(6):310– 318, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sintzoff:1975:CCE

[550] Michel Sintzoff and Axel Van Lamsweerde. Constructing correct and efficient concurrent programs. ACM SIG-PLAN Notices, 10(6):319–326, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Endres:1975:AET

[551] Albert Endres. An analysis of errors and their causes in system programs. ACM SIGPLAN Notices, 10 (6):327–336, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneidewind:1975:AEP

[552] Norman F. Schneidewind. Analysis of error processes in computer software. ACM SIGPLAN Notices, 10(6): 337–346, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shooman:1975:TDT

[553] M. L. Shooman and M. I. Bolsky. Types, distribution, and test and correction times for programming errors. ACM SIGPLAN Notices, 10(6): 347–357, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parnas:1975:ISS

[554] D. L. Parnas. The influence of software structure on reliability. ACM SIG-PLAN Notices, 10(6):358–362, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mills:1975:HWC

[555] Harlan D. Mills. How to write correct programs and know it. ACM SIG-PLAN Notices, 10(6):363–370, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buda:1975:IAP

[556] A. O. Buda, A. A. Granovsky, and A. P. Ershov. Implementation of the Alpha-6 programming system. ACM SIGPLAN Notices, 10(6):371– 381, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramamoorthy:1975:TLS

[557] C. V. Ramamoorthy and S. F. Ho. Testing large software with automated software evaluation systems. ACM SIGPLAN Notices, 10(6):382– 394, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Carpenter:1975:SDV

[558] Loren C. Carpenter and Leonard L. Tripp. Software design validation tool. ACM SIGPLAN Notices, 10(6): 395–400, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edwards:1975:ECM

[559] N. P. Edwards. The effect of certain modular design principles on testability. ACM SIGPLAN Notices, 10(6): 401–410, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Good:1975:PP

[560] Donald I. Good. Provable programming. ACM SIGPLAN Notices, 10(6): 411–419, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laventhal:1975:VPW

[561] Mark S. Laventhal. Verifying programs which operate on data structures. ACM SIGPLAN Notices, 10(6): 420–426, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Infante:1975:PSP

[562] R. Infante and U. Montanari. Proving structured programs correct, level by level. ACM SIGPLAN Notices, 10 (6):427–436, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Randell:1975:SSS

[563] B. Randell. System structure for software fault tolerance. ACM SIG-PLAN Notices, 10(6):437–449, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yau:1975:DSC

[564] S. S. Yau and R. C. Cheung. Design of self-checking software. ACM SIG-PLAN Notices, 10(6):450–457, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Avizienis:1975:FTF

[565] Algirdas Avižienis. Fault-tolerance and fault-intolerance: Complementary approaches to reliable computing. ACM SIGPLAN Notices, 10(6): 458–464, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burstall:1975:STD

[566] R. M. Burstall and John Darlington. Some transformations for developing recursive programs. ACM SIG-PLAN Notices, 10(6):465–472, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Suzuki:1975:VPA

[567] Norihisa Suzuki. Verifying programs by algebraic and logical reduction. ACM SIGPLAN Notices, 10(6):473– 481, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Good:1975:IPV

[568] Donald I. Good, Ralph L. London, and W. W. Bledsoe. An interactive program verification system. ACM SIG-PLAN Notices, 10(6):482–492, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goodenough:1975:TTT

[569] John B. Goodenough and Susan L. Gerhart. Toward a theory of test data selection. ACM SIGPLAN Notices, 10 (6):493–510, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kopetz:1975:CBR

[570] Hermann Kopetz. On the connections between range of variable and control structure testing. ACM SIG-PLAN Notices, 10(6):511–517, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brown:1975:TSR

[571] J. R. Brown and M. Lipow. Testing for software reliability. ACM SIG-PLAN Notices, 10(6):518–527, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoare:1975:DR

[572] C. A. R. Hoare. Data reliability. ACM SIGPLAN Notices, 10(6): 528–533, June 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

London:1975:VPV

[573] Ralph L. London. A view of program verification. ACM SIGPLAN Notices, 10(6):534–545, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1975:CCA

[574] Edsger W. Dijkstra. Correctness concerns and, among other things, why they are resented. ACM SIGPLAN Notices, 10(6):546–550, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wasserman:1975:IPL

[575] Anthony I. Wasserman. Issues in programming language design: an overview. ACM SIGPLAN Notices, 10 (7):10–12, July 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zahn:1975:SCP

[576] Charles T. Zahn, Jr. Structured control in programming languages. ACM SIGPLAN Notices, 10(7):13–15, July 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liskov:1975:DTP

[577] Barbara H. Liskov. Data types and program correctness. ACM SIG-PLAN Notices, 10(7):16–17, July 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Standish:1975:EPL

[578] Thomas A. Standish. Extensibility in programming language design. ACM SIGPLAN Notices, 10(7):18–21, July 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Presser:1975:SL

[579] Leon Presser. Structured languages. ACM SIGPLAN Notices, 10(7):22–24, July 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dennis:1975:EPA

[580] Jack B. Dennis. An example of programming with abstract data types. ACM SIGPLAN Notices, 10(7):25–29, July 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goodenough:1975:EHD

[581] John B. Goodenough. Exception handling design issues. ACM SIG-PLAN Notices, 10(7):41–45, July 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shneiderman:1975:CPP

[582] Ben Shneiderman. Cognitive psychology and programming language design. ACM SIGPLAN Notices, 10(7):46–47, July 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Witt:1975:PDB

[583] H. Witt. Pun-dora’s box or how to produce enough papers to wrap up structured programming. ACM SIG-PLAN Notices, 10(8):10–11, August 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mason:1975:SUD

[584] James A. Mason. Some user-defined control functions for more readable APL programs. ACM SIGPLAN Notices, 10(8):11–19, August 1975. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salvadori:1975:SPC

[585] A. Salvadori, J. Gordon, and C. Capstick. Static profile of COBOL programs. ACM SIGPLAN Notices, 10 (8):20–33, August 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Weinberg:1975:ITE

[586] Gerald M. Weinberg, Dennis P. Geller, and Thomas W. S. Plum. IF-THEN-ELSE considered harmful. ACM SIG-PLAN Notices, 10(8):34–44, August 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Denning:1975:CMO

[587] Peter J. Denning. Comments on mathematical overkill. ACM SIGPLAN Notices, 10(9):10–11, September 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foulk:1975:DTS

[588] Clinton R. Foulk. The DO trace: a simple and effective method for debugging GOTO-free programs. ACM SIG-PLAN Notices, 10(9):11–18, September 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meissner:1975:EFC

[589] Loren P. Meissner. On extending Fortran control structures to facilitate structured programming. ACM SIG-PLAN Notices, 10(9):19–30, September 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sayward:1975:CCC

[590] Frederick G. Sayward. A correctness criterion for CSP parallel programs. ACM SIGPLAN Notices, 10(9): 30–38, September 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLennan:1975:NDA

[591] B. J. MacLennan. A note on dynamic arrays in PASCAL. ACM SIGPLAN Notices, 10(9):39–40, September 1975.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:1975:FFS

[592] Reford Bond. Free form structured FORTRAN translator. ACM SIG-PLAN Notices, 10(10):12–15, October 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flon:1975:RSP

[593] Lawrence Flon. On research in structured programming. ACM SIGPLAN Notices, 10(10):16–17, October 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gales:1975:SFN

[594] Lawrence E. Gales. Structured FOR-TRAN with no preprocessor. ACM SIGPLAN Notices, 10(10):17–24, October 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lambird:1975:EIL

[595] Robert J. Lambird and Lawrence E. Cornish. The EXEC 8 implementation of List Processor-N. ACM SIG-PLAN Notices, 10(10):25–36, October 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1975:SD

[596] Daniel M. Berry. Structured documentation. ACM SIGPLAN Notices, 10(11):7–12, November 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Coakley:1975:EUC

[597] F. P. Coakley. Experience with the use of the case statement within a M.O.L. ACM SIGPLAN Notices, 10(11):13– 17, November 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dailey:1975:TAP

[598] W. H. Dailey. A tabular approach to program optimization. ACM SIG-PLAN Notices, 10(11):17–22, November 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leibowitz:1975:MOU

[599] Stephen Leibowitz. Memory optimization using block structures. ACM SIG-PLAN Notices, 10(11):22–25, November 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Middleton:1975:CFA

[600] Anthony G. Middleton. Connector functions: another view of the GOTO? ACM SIGPLAN Notices, 10(11):25– 37, November 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Williams:1975:NAC

[601] M. H. Williams. A note on the ambiguity of the common statement. ACM SIG-PLAN Notices, 10(11):38–40, November 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dailey:1975:GBD

[602] W. H. Dailey. On generating binary decision trees with minimum nodes. ACM SIGPLAN Notices, 10(12):14–21, December 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holton:1975:TEP

[603] John B. Holton. Toward efficient programming in virtual systems. ACM SIGPLAN Notices, 10(12):22–25, December 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lassagne:1975:RDI

[604] Theodore D. Lassagne. Recursive definitions for iterative algorithms. ACM SIGPLAN Notices, 10(12):26–38, December 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meissner:1975:P

[605] L. P. Meissner. On preprocessors. ACM SIGPLAN Notices, 10(12):39–??, December 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yuval:1975:SEP

[606] G. Yuval. A SNOBOL 4 execution profile. ACM SIGPLAN Notices, 10(12): 40–??, December 1975. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meissner:1976:PCS

[607] Loren P. Meissner. Proposed control structures for extended FORTRAN. ACM SIGPLAN Notices, 11(1):16–21, January 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Embley:1976:KSU

[608] David W. Embley and Wilfred J. Hansen. The KAIL selector — a unified control construct. ACM SIG-PLAN Notices, 11(1):22–29, January 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Landau:1976:HLL

[609] Charles Landau. On high-level languages for system programming. ACM SIGPLAN Notices, 11(1):30–31, January 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Winiger:1976:NOP

[610] G. Winiger. A note on one-pass case statement compilation. ACM SIG-PLAN Notices, 11(1):32–36, January 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirth:1976:CND

[611] N. Wirth. Comment on a note on dynamic arrays in PASCAL. ACM SIG-PLAN Notices, 11(1):37–38, January 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Worona:1976:PCA

[612] Steven Worona. PL/CT, another approach to two problems in interactive PL/I. ACM SIGPLAN Notices, 11(1): 39–44, January 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahams:1976:RPE

[613] Paul W. Abrahams. On realism in programming examples. ACM SIG-PLAN Notices, 11(2):17–19, February 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lawson:1976:TMA

[614] Harold W. Lawson, Jr. and David R. Doucette. A translation machine with automated top-down parsing. ACM SIGPLAN Notices, 11(2):20–29, February 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reifer:1976:SFD

[615] Donald J. Reifer. The structured Fortran dilemma. ACM SIGPLAN Notices, 11(2):30–32, February 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1976:DPA

[616] ACM. Draft proposed ANS FOR-TRAN. ACM SIGPLAN Notices, 11 (3):1–223, March 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ANSI:1976:DPA

[617] ANSI Subcommittee X3J3. Draft proposed ANS Fortran. ACM SIGPLAN Notices, 11(3):various, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See also final standard [?].

Nordstrom:1976:OMM

[618] Bengt Nordström. An outline of a mathematical model for the definition and manipulation of data. ACM SIGPLAN Notices, 11(3S):1–11, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Linden:1976:UAD

[619] Theodore A. Linden. The use of abstract data types to simplify program modifications. ACM SIGPLAN Notices, 11(3S):12–23, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zloof:1976:SBA

[620] Moshe M. Zloof and S. Peter de Jong. The System for Business Automation (SBA): Programming language. ACM SIGPLAN Notices, 11(3S):24, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allman:1976:ERD

[621] Eric Allman, Michael Stonebraker, and Gerald Held. Embedding a relational data sublanguage in a general purpose programming language. ACM SIG-PLAN Notices, 11(3S):25–35, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1976:ADT

[622] Robert T. Johnson and James B. Morris. Abstract data types in the Model programming language. ACM SIG-PLAN Notices, 11(3S):36–46, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McLeod:1976:HLD

[623] Dennis J. McLeod. High level domain definition in a relational data base system. ACM SIGPLAN Notices, 11 (3S):47–57, March 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hammer:1976:DAD

[624] Michael Hammer. Data abstractions for data bases. ACM SIGPLAN Notices, 11(3S):58–59, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Horning:1976:SDP

[625] J. J. Horning. Some desirable properties of data abstraction facilities. ACM SIGPLAN Notices, 11(3S):60–62, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ross:1976:TFU

[626] Douglas T. Ross. Toward foundations for the understanding of type. ACM SIGPLAN Notices, 11(3S):63–65, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1976:RDA

[627] Mary Shaw. Research directions in abstract data structures. ACM SIG-PLAN Notices, 11(3S):66–68, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mitchell:1976:NSD

[628] Jim Mitchell and Ben Wegbreit. A next step in data structuring for programming languages. ACM SIGPLAN Notices, 11(3S):69–70, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1976:DBA

[629] John M. Smith and Diane C. P. Smith. Data base abstraction. ACM SIG-PLAN Notices, 11(3S):71, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guttag:1976:ADT

[630] John Guttag. Abstract data types and the development of data structures. ACM SIGPLAN Notices, 11 (3S):72, March 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bubenko:1976:IRD

[631] J. A. Bubenko, Jr., S. Berild, E. Lindencrona-Ohlin, and S. Nachmens. From information requirements to DBTG-data structures. ACM SIG-PLAN Notices, 11(3S):73–85, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hardgrave:1976:TIS

[632] W. T. Hardgrave. A technique for implementing a set processor. ACM SIG-PLAN Notices, 11(3S):86–94, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kieburtz:1976:PPV

[633] Richard B. Kieburtz. Programming without pointer variables. ACM SIG-PLAN Notices, 11(3S):95–107, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vallarino:1976:UBM

[634] Oscar Vallarino. On the use of bit maps for multiple key retrieval. ACM SIG-PLAN Notices, 11(3S):108–114, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1976:CDR

[635] D. M. Berry, Z. Erlich, and C. J. Lucena. Correctness of data representations (extended abstract): Pointers in high level languages. ACM SIG-PLAN Notices, 11(3S):115–119, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gries:1976:SID

[636] David Gries and Narain Gehani. Some ideas on data types in high level languages. ACM SIGPLAN Notices, 11 (3S):120, March 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Senko:1976:DIB

[637] Michael E. Senko. DIAM II: The binary infological level and its database language — FORAL. ACM SIGPLAN Notices, 11(3S):121–140, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Flon:1976:TCV

[638] L. Flon and A. N. Habermann. Towards the construction of verifiable software systems. ACM SIGPLAN Notices, 11(3S):141–148, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parnas:1976:ATD

[639] D. L. Parnas, John E. Shore, and David Weiss. Abstract types defined as classes of variables. ACM SIGPLAN Notices, 11(3S):149–154, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Housel:1976:HLD

[640] Barron C. Housel and Nan C. Shu. A high-level data manipulation language for hierarchical data structures. ACM SIGPLAN Notices, 11(3S):155– 169, March 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:1976:DTE

[641] Mark B. Wells and Fred L. Cornwell. A data type encapsulation scheme utilizing base language operators. ACM SIGPLAN Notices, 11(3S): 170–178, March 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koster:1976:VT

[642] Cornelis H. A. Koster. Visibility and types. ACM SIGPLAN Notices, 11 (3S):179–190, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gorman:1976:MPL

[643] Walter Gorman and Michael Broussard. Minicomputer programming languages. ACM SIGPLAN Notices, 11 (4):4–15, April 1, 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wasserman:1976:BVM

[644] Anthony I. Wasserman and David D. Sherertz. A balanced view of MUMPS. ACM SIGPLAN Notices, 11(4):16–26, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Russell:1976:EDI

[645] Robert D. Russell. Experience in the design, implementation and use of PL-11, a programming language for the PDP-11. ACM SIGPLAN Notices, 11 (4):27–34, April 1, 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herrick:1976:EMC

[646] Steven S. Herrick. Experiences moving COBOL onto a minicomputer. ACM SIGPLAN Notices, 11(4):35–39, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bennett-Novak:1976:OMI

[647] Gordon Bennett-Novak. Optimized machine-independent extended FOR-TRAN for minicomputer programming. ACM SIGPLAN Notices, 11(4): 40–44, April 1, 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Johnson:1976:AGC

[648] Gearold R. Johnson and Robert A. Mueller. The automated generation of cross-system software for supporting micro/mini computer systems. ACM SIGPLAN Notices, 11(4):45–56, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Braun:1976:IMS

[649] Christine L. Braun. An integrated microprocessor support software system. ACM SIGPLAN Notices, 11(4):57–65, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Blake-Knox:1976:MSD

[650] M. W. Blake-Knox and G. C. Morris. Minicomputer software development using a maxicomputer. ACM SIGPLAN Notices, 11(4):66–69, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Oliver:1976:PCT

[651] S. Ron Oliver and Nell D. Jones. Program control via transition matrices — a novel application of microprogramming. ACM SIGPLAN Notices, 11(4):70–77, April 1, 1976. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Adix:1976:IER

[652] Mary S. Adix and Henrik A. Schutz. Interpretive execution of real-time control applications. ACM SIGPLAN Notices, 11(4):78–87, April 1, 1976.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cary:1976:SBP

[653] Ted Cary. Structuring BASIC programs for managing overlays in a small computer system. ACM SIGPLAN Notices, 11(4):88–93, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lozier:1976:NAM

[654] Daniel W. Lozier. Numerical applications of minicomputers at the National Bureau of Standards. ACM SIG-PLAN Notices, 11(4):94–101, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nash:1976:EDM

[655] J. C. Nash. Experiences in the development of mathematical software for small computers. ACM SIG-PLAN Notices, 11(4):102, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maurer:1976:PCF

[656] W. D. Maurer. Proving the correctness of a flight-director program for an airborne minicomputer. ACM SIG-PLAN Notices, 11(4):103–108, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cowan:1976:DES

[657] D. D. Cowan, P. H. Dirksen, J. W. Graham, and J. W. Welch. Development of educational software using the DEC PDP-11. ACM SIGPLAN Notices, 11(4):109–112, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vickery:1976:MDL

[658] Christopher Vickery. A microprogramming design laboratory. ACM SIG-PLAN Notices, 11(4):113–117, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schreier:1976:EML

[659] Thomas Schreier. ETOS: multi-lingual Time Sharing System for the PDP8. ACM SIGPLAN Notices, 11(4):118, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fulton:1976:AXI

[660] David L. Fulton and Richard T. Thomas. ABACUS/X an incremental compiler for minicomputer use. ACM SIGPLAN Notices, 11(4):119– 126, April 1, 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kieburtz:1976:THL

[661] R. B. Kieburtz and J. L. Hennessy. TOMAL: a high-level programming language for microprocessor process control applications. ACM SIGPLAN Notices, 11(4):127–134, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Plauger:1976:LIL

[662] P. J. Plauger. A Little Implementation Language. ACM SIGPLAN Notices, 11(4):135–137, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stocks:1976:THL

[663] A. Ian Stocks and Jayant Krishnaswamy. On a transportable high level language for minicomputers. ACM SIGPLAN Notices, 11(4):138– 143, April 1, 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Stuart:1976:ALS

[664] Thomas Stuart. Adapting large systems to small machines. ACM SIG-PLAN Notices, 11(4):144–150, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McKeeman:1976:CIT

[665] William M. McKeeman. Corporate integration through systems software. ACM SIGPLAN Notices, 11(4):151– 155, April 1, 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lennon:1976:UDM

[666] William J. Lennon, Eric L. Baatz, Sandra L. Colussi, and Kenneth E. Kinnear, Jr. Using a distributed minicomputer network to automate a biochemical laboratory. ACM SIGPLAN Notices, 11(4):156–164, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:1976:ITT

[667] Reford Bond. An intuitive translation technique. ACM SIGPLAN Notices, 11 (5):33–36, May 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buchegger:1976:IET

[668] O. Buchegger and N. E. Fuchs. If ‘ELSE’ then ‘CASE’. ACM SIG-PLAN Notices, 11(5):37, May 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elshoff:1976:MCP

[669] James L. Elshoff. Measuring commercial PL/I programs using Halstead’s criteria. ACM SIGPLAN Notices, 11 (5):38–46, May 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hamlet:1976:PBT

[670] Richard Hamlet. The PDP-11 as B5500 in teaching systems programming. ACM SIGPLAN Notices, 11 (5):47–52, May 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hardgrave:1976:PVK

[671] W. T. Hardgrave. Positional versus keyword parameter communication in programming languages. ACM SIG-PLAN Notices, 11(5):52–58, May 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Palme:1976:NFM

[672] Jacob Palme. New feature for module protection in SIMULA. ACM SIG-PLAN Notices, 11(5):59–62, May 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steensgaard-Madsen:1976:MDA

[673] J. Steensgaard-Madsen. More on dynamic arrays in PASCAL. ACM SIG-PLAN Notices, 11(5):63–64, May 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tanenbaum:1976:DPT

[674] Andrew S. Tanenbaum. In defense of program testing or correctness proofs considered harmful. ACM SIGPLAN Notices, 11(5):64–68, May 1976. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pfister:1976:HLL

[675] Gregory F. Pfister. A high level language extension for creating and controlling dynamic pictures. ACM SIG-PLAN Notices, 11(6):1–9, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schrack:1976:DIE

[676] Guenther Schrack. Design, implementation and experiences with a higherlevel graphics language for interactive computer-aided design purposes. ACM SIGPLAN Notices, 11(6):10–17, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1976:EAS

[677] Ben Jones. An extended Algol-60 for Shaded Computer Graphics. ACM SIGPLAN Notices, 11(6):18–23, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guthery:1976:DIE

[678] Scott B. Guthery. DDA: An interactive and extensible language for data display and analysis. ACM SIG-PLAN Notices, 11(6):24–31, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wardle:1976:AGL

[679] Caroline Wardle. ARTSPEAK: A graphics language for artists. ACM SIGPLAN Notices, 11(6):32–39, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kriger:1976:SHL

[680] Mark P. Kriger. SUGAR: A high-level programming language for geographical analysis and mapping. ACM SIG-PLAN Notices, 11(6):40–48, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foley:1976:PNM

[681] James D. Foley. Picture naming and modification: An overview. ACM SIG-PLAN Notices, 11(6):49–53, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thomas:1976:MSD

[682] Elaine L. Thomas. Methods for specifying display parameters in graphics programming languages. ACM SIG-PLAN Notices, 11(6):54–56, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergeron:1976:PPD

[683] R. Daniel Bergeron. Picture primitives in device independent graphics systems. ACM SIGPLAN Notices, 11 (6):57–60, June 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wallace:1976:SGI

[684] Victor L. Wallace. The semantics of graphic input devices. ACM SIG-PLAN Notices, 11(6):61–65, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leiberman:1976:TTL

[685] H. Leiberman. The TV turtle, a Logo graphics system for raster displays. ACM SIGPLAN Notices, 11(6):66–72, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lieberman:1976:TTL

[686] Henry Lieberman. The TV Turtle: a Logo graphics system for raster displays. ACM SIGPLAN Notices, 11 (6):66–72, June 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Katz:1976:PST

[687] Lou Katz and Bill Etra. A preliminary step towards a language for picture control in a real-time mode. ACM SIGPLAN Notices, 11(6):73–76, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeFanti:1976:CSP

[688] T. A. DeFanti, D. J. Sandin, and R. A. Ainsworth. Control structures for performance graphics. ACM SIG-PLAN Notices, 11(6):79–84, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gujar:1976:DIC

[689] Uday G. Gujar. A device independent computer plotting system. ACM SIGPLAN Notices, 11(6):85–100, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Garrett:1976:IIS

[690] Michael T. Garrett. An interpretive/ interactive subroutine system for raster graphics. ACM SIGPLAN Notices, 11 (6):101–105, June 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shen-PeiWang:1976:LTD

[691] P. Shen-Pei Wang and W. I. Grosky. A language for two-dimensional digital picture processing. ACM SIG-PLAN Notices, 11(6):106–112, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wang:1976:LTD

[692] Patrick Shen-Pei Wang and William I. Grosky. A language for twodimensional digital picture processing. ACM SIGPLAN Notices, 11(6): 106–112, June 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Suski:1976:AGC

[693] Gregory J. Suski. Automatic generation of computer graphics languages. ACM SIGPLAN Notices, 11(6):113– 122, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herring:1976:DUG

[694] Travis L. Herring. On the design and use of graphics-oriented continuous system simulation languages (CSSL). ACM SIGPLAN Notices, 11 (6):123–128, June 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Archibald:1976:TDF

[695] Julius A. Archibald, Jr. Towards a dynamic (FORTRAN) programming system. ACM SIGPLAN Notices, 11 (7):17–24, July 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allison:1976:DNT

[696] Dennis Allison. Design notes for TINY BASIC. ACM SIGPLAN Notices, 11 (7):25–33, July 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bennett:1976:BPA

[697] Richard K. Bennett. BUILD: a primitive approach to the design of computer languages and their translators. ACM SIGPLAN Notices, 11(7):34–40, July 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brown:1976:TBL

[698] Robert E. Brown. Toward a better language for structured programming. ACM SIGPLAN Notices, 11(7):41–54, July 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ward:1976:NIS

[699] Ronnie G. Ward. A non-iterative situation case statement. ACM SIG-PLAN Notices, 11(7):55–62, July 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Williams:1976:QAS

[700] M. H. Williams. A question-answering system for automatic program synthesis. ACM SIGPLAN Notices, 11(7):63– 68, July 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feurman:1976:SPC

[701] Melvyn Feurman, Mary R. Dallal, and Lillian Liebling. Structured programming in COBOL under IBM 360/370 OS. ACM SIGPLAN Notices, 11(8): 16–30, August 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barth:1976:SES

[702] C. Wrandle Barth. STRCMACS: an extensive set of macros to aid in structured programming in 360/ 370 assembly language. ACM SIG-PLAN Notices, 11(8):31–35, August 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matuszek:1976:CAS

[703] David Matuszek. The case for the assert statement. ACM SIGPLAN Notices, 11(8):36–37, August 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Misuri:1976:SEP

[704] Giorgio Misuri. Survey of existing programming aids. ACM SIGPLAN Notices, 11(8):38–41, August 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Matwin:1976:SMI

[705] S. Matwin and M. Missala. A simple, machine independent tool for obtaining rough measures of PASCAL programs. ACM SIGPLAN Notices, 11 (8):42–45, August 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McClure:1976:OCD

[706] C. L. McClure. Ordering control data in structured COBOL. ACM SIG-PLAN Notices, 11(8):46–51, August 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bridge:1976:BTI

[707] R. F. Bridge and E. W. Thompson. Bridges: a tool for increasing the reliability of references to Fortran variables. ACM SIGPLAN Notices, 11(9): 2–9, September 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cook:1976:EEP

[708] A. James Cook. Experience with extensible, portable Fortran extensions. ACM SIGPLAN Notices, 11(9):10– 17, September 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Musstoph:1976:PSP

[709] Guenter Musstoph. The programming system POLYP: structure and experience. ACM SIGPLAN Notices, 11(9): 18–24, September 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wells:1976:PTT

[710] Mark B. Wells. Preprocessing of typed two-dimensional mathematical expressions. ACM SIGPLAN Notices, 11(9): 25–37, September 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wolley:1976:LEL

[711] J. D. Wolley, L. R. Miller, and C. M. Bernstein. Linus: an experiment in language preprocessing. ACM SIGPLAN Notices, 11(9):38–48, September 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Woolley:1976:LEL

[712] John D. Woolley, Leland R. Miller, and Charles M. Bernstein. LINUS: an experiment in language preprocessing. ACM SIGPLAN Notices, 11(9): 38–48, September 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeMillo:1976:CSP

[713] R. A. DeMillo, S. C. Eisenstat, and R. J. Lipton. Can structured programs be efficient? ACM SIGPLAN Notices, 11(10):10–18, October 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kessels:1976:PPC

[714] J. L. W. Kessels. Parallel programming concepts in a definitional language. ACM SIGPLAN Notices, 11(10):19– 31, October 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kittlitz:1976:BSS

[715] E. N. Kittlitz. Block statements and synonyms for PASCAL. ACM SIG-PLAN Notices, 11(10):32–35, October 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pokrovsky:1976:FTT

[716] Sergei Pokrovsky. Formal types and their application to dynamic arrays in PASCAL. ACM SIGPLAN Notices, 11(10):36–42, October 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rowland:1976:SCS

[717] Stuart W. Rowland. Some comments on structured FORTRAN. ACM SIG-PLAN Notices, 11(10):43–48, October 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simon:1976:CDT

[718] Janos Simon. A comment on DO Traces. ACM SIGPLAN Notices, 11 (10):49–52, October 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Conradi:1976:FCC

[719] Reidar Conradi. ‘further critical comments on PASCAL, particularly as a systems programming language’. ACM SIGPLAN Notices, 11(11):8– 25, November 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Goodell:1976:GAI

[720] Ross W. Goodell. A generalized approach to interpretation machines. ACM SIGPLAN Notices, 11(11):26– 37, November 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fleck:1976:ICE

[721] A. C. Fleck. On the impossibility of content exchange through the byname parameter transmission mechanism. ACM SIGPLAN Notices, 11(11): 38–41, November 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Love:1976:ITT

[722] L. T. Love and A. B. Bowman. An independent test of the theory of software physics. ACM SIGPLAN Notices, 11(11):42–49, November 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marlin:1976:EES

[723] C. D. Marlin. An experiment with the extensibility of SIMULA. ACM SIG-PLAN Notices, 11(11):50–57, November 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Roy:1976:LFG

[724] Patrick Roy and Richard St-Denis. Linear flowchart generator for a structured language. ACM SIGPLAN Notices, 11(11):58–64, November 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Conway:1976:PCH

[725] R. Conway. PL/CS: a highlydisciplined subset of PL/C. ACM SIG-PLAN Notices, 11(12):21–24, December 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldman:1976:FLC

[726] S. I. Feldman. A Fortranner’s lament: Comments on the draft proposed ANS FORTRAN standard. ACM SIGPLAN Notices, 11(12):25–34, December 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gottlieb:1976:FP

[727] Allan Gottlieb. A flowcharting proposal. ACM SIGPLAN Notices, 11(12): 35–37, December 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kahn:1976:PSD

[728] Lars Kahn. Program structure definition using binary relations. ACM SIG-PLAN Notices, 11(12):38–55, December 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Korncoff:1976:SSA

[729] Alan R. Korncoff and Charles H. Goodspeed. SAGES: System Aiding the Generation of Engineering Software. ACM SIGPLAN Notices, 11(12):56– 73, December 1976. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nylin:1976:MTC

[730] William C. Nylin, Jr. and John B. Harvill. Multiple tense computer programming. ACM SIGPLAN Notices, 11(12):74–93, December 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schwabe:1976:SUR

[731] D. Schwabe and C. J. Lucena. Specification and uniform reference to data structures in PL/1. ACM SIG-PLAN Notices, 11(12):94–104, December 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sehwabe:1976:SUR

[732] Daniel Sehwabe and Carlos J. Lucena. Specification and uniform reference to data structures in PL/I. ACM SIG-PLAN Notices, 11(12):94–104, December 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chevance:1977:DHLa

[733] R. J. Chevance. Design of high level language oriented processors. ACM SIGPLAN Notices, 12(1):40–51, January 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeMorgan:1977:SAR

[734] R. M. De Morgan, I. D. Hill, and B. A. Wichmann. A supplement to the AL-GOL 60 revised report. ACM SIG-PLAN Notices, 12(1):52–66, January 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldman:1977:EEA

[735] Michael B. Feldman. Embedding extended arithmetic in SNOBOL4. ACM SIGPLAN Notices, 12(1):67–72, January 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kittlitz:1977:APV

[736] Edward N. Kittlitz. Another proposal for variable size arrays in PASCAL. ACM SIGPLAN Notices, 12(1):82–86, January 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ludemann:1977:LDP

[737] Peter Ludemann. A language design for program structure. ACM SIG-PLAN Notices, 12(1):87–92, January 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Meissner:1977:F

[738] L. P. Meissner. FORTRAN 77. ACM SIGPLAN Notices, 12(1):93–94, January 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salomon:1977:DFF

[739] David Salomon. A design for Fortran to facilitate structured programming. ACM SIGPLAN Notices, 12(1):95–100, January 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lampson:1977:RPL

[740] B. W. Lampson, J. J. Horning, R. L. London, J. G. Mitchell, and G. J. Popek. Report on the programming language Euclid. ACM SIGPLAN Notices, 12(2):1–77, February 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ambler:1977:GLS

[741] Allen L. Ambler, Donald I. Good, James C. Browne, Wilhelm F. Burger, Richard M. Cohen, Charles G. Hoch, and Robert E. Wells. Gypsy: A language for specification and implementation of verifiable programs. ACM SIGPLAN Notices, 12(3):1–10, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Popek:1977:NDE

[742] G. J. Popek, J. J. Horning, B. W. Lampson, J. G. Mitchell, and R. L. London. Notes on the design of Euclid. ACM SIGPLAN Notices, 12(3):11–18, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fischer:1977:EIO

[743] Charles N. Fischer and Richard J. LeBlanc. Efficient implementation and optimization of run-time checking in PASCAL. ACM SIGPLAN Notices, 12 (3):19–24, March 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ambler:1977:SPP

[744] Allen L. Ambler and Charles G. Hoch. A study of protection in programming languages. ACM SIGPLAN Notices, 12 (3):25–40, March 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Friedman:1977:AAP

[745] Daniel P. Friedman and David S. Wise. Aspects of applicative programming for file systems (preliminary version). ACM SIGPLAN Notices, 12(3):41–55, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Herriot:1977:TIP

[746] Robert G. Herriot. Towards the ideal programming language. ACM SIG-PLAN Notices, 12(3):56–62, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Guttag:1977:SEA

[747] Jhon V. Guttag, Ellis Horowitz, and David R. Musser. Some extensions to algebraic specifications. ACM SIG-PLAN Notices, 12(3):63–67, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Buckle:1977:RDT

[748] Normand Buckle. Restricted data types, specification and enforcement of invariant properties of variables. ACM SIGPLAN Notices, 12(3):68–76, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Causot:1977:SDD

[749] Patrick Causot and Radhia Cousot. Static determination of dynamic properties of generalized type unions. ACM SIGPLAN Notices, 12(3):77–94, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cousot:1977:SDD

[750] P. Cousot and R. Cousot. Static determination of dynamic properties of generalized type unions. ACM SIG-PLAN Notices, 12(3):77–94, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Melliar-Smith:1977:SRR

[751] P. M. Melliar-Smith and B. Randell. Software reliability: The role of programmed exception handling. ACM SIGPLAN Notices, 12(3):95–100, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLaren:1977:EHP

[752] M. Donald MacLaren. Exception handling in PL/I. ACM SIGPLAN Notices, 12(3):101–104, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Love:1977:EIE

[753] Tom Love. An experimental investigation of the effect of program structure on program understanding. ACM SIG-PLAN Notices, 12(3):105–113, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Andrews:1977:LFP

[754] Gregory R. Andrews and James R. McGraw. Language features for process interaction. ACM SIGPLAN Notices, 12(3):114–127, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lomet:1977:PSS

[755] D. B. Lomet. Process structuring, synchronization, and recovery using atomic actions. ACM SIGPLAN Notices, 12(3):128–137, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Geschke:1977:EEM

[756] Charles M. Geschke, James H. Morris, Jr., and Edwin H. Satterthwaite. Early experience with Mesa. ACM SIGPLAN Notices, 12(3):138, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1977:AVA

[757] Mary Shaw, Wm A. Wulf, and Ralph L. London. Abstraction and verification in Alphard: Defining and specifying iteration and generators. ACM SIG-PLAN Notices, 12(3):139, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Liskov:1977:AMC

[758] Barbara Liskov, Alan Snyder, Russell Atkinson, and Craig Schaffert. Abstraction mechanisms in CLU. ACM SIGPLAN Notices, 12(3):140, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gannon:1977:EED

[759] J. D. Gannon. An experimental evaluation of data types on programming reliability. ACM SIGPLAN Notices, 12 (3):141, March 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wirth:1977:TDR

[760] N. Wirth. Towards a discipline of real-time programming. ACM SIG-PLAN Notices, 12(3):142, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1977:FDN

[761] Anonymous. Fortran Development Newsletter: Volume 3, Number 1, January 1977, pp. 1–10: X3J3 expects to complete Fortran 77 revised at March meeting. ACM SIGPLAN Notices, 12 (4):21–30, April 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Biedl:1977:EPL

[762] Albrecht Biedl. An extension of programming languages for clerical computation in science and engineering with special reference to PASCAL. ACM SIGPLAN Notices, 12(4):31–33, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Boddy:1977:SFP

[763] David E. Boddy. Structured Fortran: with or without a preprocessor. ACM SIGPLAN Notices, 12(4):34–39, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griswold:1977:OS

[764] Ralph E. Griswold and David R. Hanson. An overview of SL5. ACM SIGPLAN Notices, 12(4):40–50, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hamlet:1977:IAC

[765] Richard Hamlet. Ignorance of AL-GOL 68 considered harmful. ACM SIGPLAN Notices, 12(4):51–56, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hartnett:1977:SFO

[766] Edward T. Hartnett. Star Fortran: an overview of essential characteristics. ACM SIGPLAN Notices, 12(4):57–66, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Huybrechts:1977:DSS

[767] M. Huybrechts. DYNOSOR: a set of subroutines for dynamic memory organization in Fortran programs. ACM SIGPLAN Notices, 12(4):67–74, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Middleton:1977:MAA

[768] Tony Middleton. A macro approach to abstractions of data structures. ACM SIGPLAN Notices, 12(4):75–79, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Robinson:1977:AMR

[769] S. K. Robinson and I. S. Torsun. The automatic measurement of the relative merits of student programs. ACM SIGPLAN Notices, 12(4):80–93, April 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Beckman:1977:CCH

[770] Anders Beckman. Comments considered harmful. ACM SIGPLAN Notices, 12(4):94–96, April 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanWijngaarden:1977:RRA

[771] A. van Wijngaarden, B. J. Mailloux, J. E. L. Peck, Koster, C. H. A., M. Sintzoff, C. H. Lindsey, L. G. L. T. Meertens, and R. G. Fisker. Revised report on the algorithmic language AL-GOL 68. ACM SIGPLAN Notices, 12 (5):1–70, May 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hibbard:1977:SA

[772] P. G. Hibbard. A sublanguage of AL-GOL 68. ACM SIGPLAN Notices, 12 (5):71–79, May 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1977:RSH

[773] W. J. Hansen and H. Boom. The report on the standard hardware representation for ALGOL 68. ACM SIG-PLAN Notices, 12(5):80–87, May 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanderMeulen:1977:AMB

[774] S. G. van der Meulen. ALGOL 68 might-have-beens. ACM SIG-PLAN Notices, 12(6):1–18, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Craven:1977:DTA

[775] P. G. Craven. Derivatives without tears in Algol 68. ACM SIGPLAN Notices, 12(6):19–26, June 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brailsford:1977:RTP

[776] D. F. Brailsford, E. Foxley, K. C. Mander, and D. J. Morgan. Run-time profiling of Algol 68-R programs using DIDYMUS and SCAMP. ACM SIGPLAN Notices, 12(6):27–35, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hennell:1977:QTE

[777] M. A. Hennell, D. Hedley, and M. R. Woodward. Quantifying the test effectiveness of Algol 68 programs. ACM SIGPLAN Notices, 12(6):36–41, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Szoke:1977:SRN

[778] Péter Szöke. Some remarks on new instances and garbage collection. ACM SIGPLAN Notices, 12(6):42–48, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hunter:1977:LVL

[779] R. B. Hunter, A. D. McGettrick, and R. Patel. LL versus LR parsing with illustrations from ALGOL 68. ACM SIGPLAN Notices, 12(6):49–53, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pagan:1977:AIL

[780] Frank G. Pagan. ALGOL 68 as an implementation language for portable interpreters. ACM SIGPLAN Notices, 12 (6):54–62, June 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Vansina:1977:DDM

[781] C. F. Vansina. Description of dynamic manipulation of ALGOL 68 values using a generative grammar. ACM SIG-PLAN Notices, 12(6):63–70, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

vanVliet:1977:TMI

[782] J. C. van Vliet. Towards a machineindependent transput section. ACM SIGPLAN Notices, 12(6):71–77, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grune:1977:TDS

[783] Dick Grune. Towards the design of a super-language of ALGOL 68 for the Standard Prelude (excerpt). ACM SIGPLAN Notices, 12(6):78–81, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Birrell:1977:SMA

[784] Andrew D. Birrell. Storage management for ALGOL 68. ACM SIG-PLAN Notices, 12(6):82–94, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gardner:1977:TA

[785] P. J. Gardner. A transportation of AL-GOL 68C. ACM SIGPLAN Notices, 12 (6):95–101, June 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Koch:1977:BAI

[786] Wilfried Koch and Christoph Oeters. The Berlin ALGOL 68 implementation. ACM SIGPLAN Notices, 12(6): 102–108, June 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Taupin:1977:ACP

[787] Daniel Taupin. The ALGOL 68 compiler of Paris-XI University (Orsay). ACM SIGPLAN Notices, 12(6): 109–116, June 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Houssais:1977:VAI

[788] Bernard Houssais. Verification of an Algol 68 implementation. ACM SIG-PLAN Notices, 12(6):117–128, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Simonet:1977:ADS

[789] M. Simonet. An attribute description of a subset of Algol 68. ACM SIG-PLAN Notices, 12(6):129–137, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prudom:1977:SPC

[790] A. Prudom and M. A. Hennell. Some problems concerning the automatic translation of Fortran to Algol 68. ACM SIGPLAN Notices, 12(6):138– 143, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Housden:1977:CSP

[791] R. J. W. Housden and N. Kotarski. Character string pattern matching in Algol 68. ACM SIGPLAN Notices, 12 (6):144–152, June 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bennett:1977:IPI

[792] M. W. Bennett. Implementation of a PDP11/ICL1900 cross assembler in Algol 68R. ACM SIGPLAN Notices, 12 (6):153–156, June 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ratcliff:1977:ASP

[793] Bryan Ratcliff. ALGOL 68 and structured programming for learnerprogrammers. ACM SIGPLAN Notices, 12(6):157–160, June 1977. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Foster:1977:ADF

[794] J. M. Foster and P. D. Foster. Abstract data and functors. ACM SIG-PLAN Notices, 12(6):161–167, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Braid:1977:GMA

[795] I. C. Braid and R. C. Hillyard. Geometric modelling in ALGOL 68. ACM SIG-PLAN Notices, 12(6):168–174, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

West:1977:MRT

[796] P. R. West. A Multi Radar Tracking simulation using ALGOL 68’R. ACM SIGPLAN Notices, 12(6):175– 178, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rayward-Smith:1977:UPL

[797] V. J. Rayward-Smith. Using procedures in list processing. ACM SIG-PLAN Notices, 12(6):179–183, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Edwards:1977:PLS

[798] Roy Edwards. Is Pascal a logical subset of Algol 68 or not? Part One. ACM SIGPLAN Notices, 12(6): 184–191, June 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Freiburghouse:1977:PEP

[799] R. A. Freiburghouse. Proposed extensions to PL/I for real-time applications. ACM SIGPLAN Notices, 12(7):26–42, July 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comment [904].

Atkinson:1977:NDS

[800] George Atkinson. The non-desirability of structured programming in user languages. ACM SIGPLAN Notices, 12 (7):43–50, July 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cichelli:1977:GDP

[801] Richard J. Cichelli and Martha J. Cichelli. Goal directed programming. ACM SIGPLAN Notices, 12(7):51–59, July 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dembinski:1977:TP

[802] Piotr Dembinski and Richard Schwartz. The taming of the pointer. ACM SIG-PLAN Notices, 12(7):60–74, July 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grune:1977:VC

[803] Dick Grune. A view of coroutines. ACM SIGPLAN Notices, 12(7):75–81, July 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hueras:1977:AFP

[804] Jon Hueras and Henry Ledgard. An automatic formatting program for PAS-CAL. ACM SIGPLAN Notices, 12 (7):82–84, July 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karasz:1977:ISL

[805] Peter Karasz. An intermediate systems language for the PDP-11. ACM SIGPLAN Notices, 12(7):85–100, July 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ledgard:1977:BEP

[806] Henry Ledgard, Andrew Singer, and Jon Hueras. A basis for executing PASCAL programmers. ACM SIG-PLAN Notices, 12(7):101–105, July 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Thomson:1977:ECT

[807] C. M. Thomson. Error checking, tracing, and dumping in an ALGOL 68 checkout compiler. ACM SIG-PLAN Notices, 12(7):106–111, July 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Woolley:1977:FCN

[808] John D. Woolley. FORTRAN: A comparison of the new proposed language (1976) to the old standard (1966). ACM SIGPLAN Notices, 12 (7):112–125, July 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Levy:1977:SRA

[809] Michael R. Levy. Some remarks on abstract data types. ACM SIG-PLAN Notices, 12(7):126–128, July 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cousot:1977:ASO

[810] Patrick Cousot and Radhia Cousot. Automatic synthesis of optimal invariant assertions: Mathematical foundations. ACM SIGPLAN Notices, 12 (8):1–12, August 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yelowitz:1977:AIP

[811] Lawrence Yelowitz and Arthur G. Duncan. Abstractions, instantiations, and proofs of marking algorithms. ACM SIGPLAN Notices, 12(8):13–21, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ulrich:1977:PSA

[812] John Wade Ulrich and Robert Moll. Program synthesis by analogy. ACM SIGPLAN Notices, 12(8):22–28, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Manna:1977:ASR

[813] Zohar Manna and Richard Waldinger. The automatic synthesis of recursive programs. ACM SIGPLAN Notices, 12 (8):29–36, August 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rychener:1977:CRD

[814] Michael D. Rychener. Control requirements for the design of production system architectures. ACM SIG-PLAN Notices, 12(8):37–44, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Davis:1977:GPC

[815] Randall Davis. Generalized procedure calling and content-directed invocation. ACM SIGPLAN Notices, 12 (8):45–54, August 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1977:IGC

[816] Henry C. Baker, Jr. and Carl Hewitt. The incremental garbage collection of processes. ACM SIGPLAN Notices, 12 (8):55–59, August 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steele:1977:MBT

[817] Guy Lewis Steele, Jr. Macaroni is better than spaghetti (stack implementation). ACM SIGPLAN Notices, 12(8): 60–66, August 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Travis:1977:DRT

[818] Larry Travis, Masahiro Honda, Richard LeBlanc, and Stephen Zeigler. Design rationale for TELOS, a PASCALbased AI language. ACM SIGPLAN Notices, 12(8):67–76, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wile:1977:ADP

[819] David Wile, Robert Balzer, and Neil Goldman. Automated derivation of program control structure from natural language program descriptions. ACM SIGPLAN Notices, 12(8):77–84, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hobbs:1977:WNN

[820] Jerry R. Hobbs. What the nature of natural language tells us about how to make natural-language-like programming languages more natural. ACM SIGPLAN Notices, 12(8):85–93, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griswold:1977:LFP

[821] Ralph E. Griswold and David R. Hanson. Language facilities for programmable backtracking. ACM SIG-PLAN Notices, 12(8):94–99, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bohm:1977:CLC

[822] H. P. Böhm, H. L. Fischer, and P. Raulefs. CSSA: Language concepts and programming methodology. ACM SIGPLAN Notices, 12(8):100–108, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Warren:1977:PLI

[823] David H. D. Warren, Luis M. Pereira, and Fernando C. N. Pereira. Prolog: The language and its implementation compared with Lisp. ACM SIGPLAN Notices, 12(8):109–115, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). Also published in SIGART NEWSLETTER, ACM Symposium on A.I. and Programming Languages.

deKleer:1977:AEC

[824] Johan de Kleer, Jon Doyle, Guy L. Steele, Jr., and Gerald Jay Sussman. AMORD explicit control of reasoning. ACM SIGPLAN Notices, 12(8): 116–125, August 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraser:1977:KBC

[825] Christopher W. Fraser. A knowledgebased code generator generator. ACM SIGPLAN Notices, 12(8):126–129, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCune:1977:PPM

[826] Brian P. McCune. The PSI Program Model Builder — synthesis of very high-level programs. ACM SIGPLAN Notices, 12(8):130–139, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kant:1977:SEI

[827] Elaine Kant. The selection of efficient implementations for a high-level language. ACM SIGPLAN Notices, 12(8): 140–146, August 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosenschein:1977:SRD

[828] Stanley J. Rosenschein and Shmuel M. Katz. Selection of representations for data structures. ACM SIGPLAN Notices, 12(8):147–154, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Samet:1977:NFC

[829] Hanan Samet. A normal form for compiler testing. ACM SIGPLAN Notices, 12(8):155–162, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kibler:1977:PME

[830] D. F. Kibler, J. M. Neighbors, and T. A. Standish. Program manipulation via an efficient production system. ACM SIGPLAN Notices, 12(8): 163–173, August 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chatelin:1977:SRP

[831] Philippe Chatelin. Self-redefinition as a program manipulation strategy. ACM SIGPLAN Notices, 12(8):174–179, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dowsing:1977:SPC

[832] R. D. Dowsing. Structured programming constructs for concurrency in program optimisation. ACM SIGPLAN Notices, 12(9):31–35, September 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Keller:1977:CCC

[833] Roy F. Keller. On control constructs for constructing programs. ACM SIG-PLAN Notices, 12(9):36–44, September 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Malkosh:1977:IPP

[834] Menachem Malkosh. Internal procedure parameters in structured FOR-TRAN precompilers. ACM SIGPLAN Notices, 12(9):45–51, September 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rechenberg:1977:MMI

[835] Peter Rechenberg. MUMS: a machine independent programming language consisting of 360 assembler macro calls. ACM SIGPLAN Notices, 12(9):52–59, September 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ris:1977:UDF

[836] Frederic N. Ris. A unified decimal floating-point architecture for the support of high-level languages. ACM SIG-PLAN Notices, 12(9):60–70, September 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rutter:1977:UHE

[837] Paul Rutter. Uniform handling of exceptions in a stack based language. ACM SIGPLAN Notices, 12(9):71– 76, September 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silverston:1977:ESS

[838] Stefan M. Silverston. Extensions to SNOBOL4 in the SNOBAT implementation. ACM SIGPLAN Notices, 12(9): 77–84, September 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smith:1977:PFD

[839] David N. Smith. Proposals for Fortran data structures. ACM SIGPLAN Notices, 12(9):85–102, September 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Terashima:1977:HLS

[840] Nobuyoshi Terashima. The hierarchical language system. ACM SIGPLAN Notices, 12(9):103–113, September 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Witty:1977:SRP

[841] Robert W. Witty. The switching reverse Polish algorithm. ACM SIG-PLAN Notices, 12(9):114–123, September 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Yasumura:1977:ELS

[842] Michiaki Yasumura. Evolution of loop statements. ACM SIGPLAN Notices, 12(9):124–129, September 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fraley:1977:RF

[843] R. A. Fraley. On replacing Fortran. ACM SIGPLAN Notices, 12(9):130– 132, September 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chevance:1977:DHLb

[844] R. J. Chevance. Design of high level language oriented processors. ACM SIGPLAN Notices, 12(10):10–32, October 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Leibowitz:1977:IC

[845] Stephen Leibowitz. Interactive computing. ACM SIGPLAN Notices, 12 (10):33–36, October 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Majster:1977:LAS

[846] Mila E. . Majster. Limits of the “algebraic” specification of abstract data types. ACM SIGPLAN Notices, 12 (10):37–42, October 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Middleton:1977:SPS

[847] Tony Middleton. Specifying program structure through sequence relationships. ACM SIGPLAN Notices, 12(10): 43–47, October 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Musinski:1977:LRE

[848] Jean E. Musinski. Lookahead recall error recovery for LALR parsers. ACM SIGPLAN Notices, 12(10):48–60, October 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Myers:1977:ECM

[849] Glenford J. Myers. An extension to the cyclomatic measure of program complexity. ACM SIGPLAN Notices, 12(10):61–64, October 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shapiro:1977:FIO

[850] Michael D. Shapiro. Fortran 77 inputoutput seems out of touch. ACM SIG-PLAN Notices, 12(10):65–69, October 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Silverston:1977:NPM

[851] Stefan M. Silverston. A note on pattern matching under Quickscan in SNOBOL4. ACM SIGPLAN Notices, 12(10):70–74, October 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Filipski:1977:CRM

[852] Alan Filipski. Call by restricted memory reference. ACM SIGPLAN Notices, 12(10):75–77, October 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lindamood:1977:WN

[853] George E. Lindamood. What’s in a name? ACM SIGPLAN Notices, 12 (10):78–80, October 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bates:1977:EPC

[854] D. Bates and R. Cailliau. Experience with Pascal compilers on minicomputers. ACM SIGPLAN Notices, 12(11):10–22, November 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Condict:1977:PDA

[855] Michael N. Condict. The Pascal dynamic array controversy and a method for enforcing global assertions.

ACM SIGPLAN Notices, 12(11):23– 27, November 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ditzel:1977:MFO

[856] D. R. Ditzel and P. C. Hutchison. MASK and FORMAT: operators for editing and formatting. ACM SIG-PLAN Notices, 12(11):28–35, November 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lindsey:1977:SCS

[857] C. H. Lindsey. Structure charts — a structured alternative to flowcharts. ACM SIGPLAN Notices, 12(11):36– 49, November 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Radhakrishnan:1977:TPL

[858] T. Radhakrishnan and I. A. Syed. Towards a programming language for the implementation of information storage and retrieval systems. ACM SIGPLAN Notices, 12(11):50–54, November 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sedgwick:1977:SCF

[859] Arthur Sedgwick. Structuring control in Fortran. ACM SIGPLAN Notices, 12(11):55–60, November 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Steele:1977:ASC

[860] Guy L. Steele, Jr. Arithmetic shifting considered harmful. ACM SIGPLAN Notices, 12(11):61–69, November 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wilcox:1977:ALS

[861] Thomas R. Wilcox. Aspen language specifications. ACM SIGPLAN Notices, 12(11):70–87, November 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arisawa:1977:PLH

[862] Makoto Arisawa. Programming languages with hierarchical structure. ACM SIGPLAN Notices, 12(12):36– 38, December 1977. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1977:DDR

[863] Anonymous. Department of Defense requirements for high order computer programming languages. revised ‘IRONMAN’ July, 1977. the technical requirements. ACM SIGPLAN Notices, 12(12):39–54, December 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abramson:1977:OTG

[864] Harvey Abramson. Ordered types and a generalized \underline{for} statement. ACM SIG-PLAN Notices, 12(12):55–59, December 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brainerd:1977:PFL

[865] Walt Brainerd. A proposal for a Fortran loop construct. ACM SIGPLAN Notices, 12(12):60–67, December 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chroust:1977:SCE

[866] G. Chroust. Scope conserving expression evaluation: a class of storage optimal evaluation strategies for arithmetic expressions. ACM SIGPLAN Notices, 12(12):68–72, December 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Richard:1977:RLD

[867] Frederic Richard and Henry F. Ledgard. A reminder for language designers. ACM SIGPLAN Notices, 12 (12):73–82, December 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Peterson:1977:FPP

[868] James L. Peterson. On the formatting of Pascal programs. ACM SIGPLAN Notices, 12(12):83–86, December 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Richardson:1977:UNI

[869] M. G. Richardson. The use of names to indicate the scope of structured language constructs. ACM SIGPLAN Notices, 12(12):87, December 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Benediktsson:1978:NAP

[870] Oddur Benediktsson. Notes on argument-parameter association in Fortran. ACM SIGPLAN Notices, 13 (1):16–20, January 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Brand:1978:NDA

[871] D. Brand. A note on data abstractions. ACM SIGPLAN Notices, 13(1): 21–24, January 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1978:PCS

[872] Anonymous. Parallelism, control and synchronization expression in a single assignment language. ACM SIG-PLAN Notices, 13(1):25–33, January 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Comte:1978:PCS

[873] D. Comte, G. Durrieu, O. Gelly, A. Plas, and J. C. Syre. Parallelism, control and synchronization expression in a single assignment language. ACM SIGPLAN Notices, 13(1):25–33, January 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ferstl:1978:FSR

[874] O. Ferstl. Flowcharting by stepwise refinement. ACM SIGPLAN Notices, 13 (1):34–42, January 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ceorges:1978:PAE

[875] J. Ceorges and F. Heymans. On parametric array expressions. ACM SIG-PLAN Notices, 13(1):43–47, January 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Georges:1978:PAE

[876] J. Georges and F. Heymans. On parametric array expressions. ACM SIG-PLAN Notices, 13(1):43–47, January 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Iglewski:1978:CIP

[877] M. Iglewski, J. Madey, and S. Matwin. A contribution to an improvement of Pascal. ACM SIGPLAN Notices, 13 (1):48–58, January 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Payne:1978:FTE

[878] A. J. Payne. A formalised technique for expressing compiler exercisers. ACM SIGPLAN Notices, 13(1):59–69, January 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wortman:1978:RXI

[879] David B. Wortman. A rooster of XPL implementations. ACM SIG-PLAN Notices, 13(1):70–74, January 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Whitaker:1978:UDD

[880] W. A. Whitaker. The US Department of Defense common higher order language effort. ACM SIGPLAN Notices, 13(2):19–29, February 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elshoff:1978:IEC

[881] James L. Elshoff. An investigation into the effects of the counting method used on software science measurements. ACM SIGPLAN Notices, 13(2):30– 45, February 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grouse:1978:FTS

[882] Phil Grouse. “flowblocks” — a technique for structured programming. ACM SIGPLAN Notices, 13(2):46– 56, February 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nutt:1978:CPF

[883] Gary J. Nutt. A comparison of PAS-CAL and FORTRAN as introductory programming languages. ACM SIG-PLAN Notices, 13(2):57–62, February 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pflasterer:1978:AEF

[884] David C. Pflasterer. An approach to efficient, fault tolerant programming. ACM SIGPLAN Notices, 13(2): 63–70, February 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wood:1978:BTD

[885] D. Wood. A bibliography of top-down deterministic parsing. ACM SIGPLAN Notices, 13(2):71–76, February 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1978:NPE

[886] Anonymous. News of past events. ACM SIGPLAN Notices, 13(3):9–10, March 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Glass:1978:PAL

[887] Robert L. Glass. Patching is alive and, lamentably, thriving in the real-time world. ACM SIGPLAN Notices, 13 (3):25–28, March 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1978:MPC

[888] Wilfred J. Hansen. Measurement of program complexity by the pair: (cyclomatic number, operator count). ACM SIGPLAN Notices, 13(3):29–33, March 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chang:1978:ADT

[889] Ernest Chang, Neil E. Kaden, and W. David Elliott. Abstract data types in Euclid. ACM SIGPLAN Notices, 13 (3):34–42, March 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Aseltine:1978:IMD

[890] Edward G. Aseltine and Henry Spencer. Isolation of machine dependencies in Euclid. ACM SIG-PLAN Notices, 13(3):43–48, March 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

desRivieres:1978:RWE

[891] Jim des Rivieres and Henry Spencer. Readability and writability in Euclid. ACM SIGPLAN Notices, 13(3):49–56, March 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Venema:1978:EP

[892] Ted Venema and Jim des Rivieres. Euclid and Pascal. ACM SIGPLAN Notices, 13(3):57–69, March 1978. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barnard:1978:EM

[893] David T. Barnard, W. David Elliott, and David H. Thompson. Euclid and Modula. ACM SIGPLAN Notices, 13 (3):70–84, March 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elliott:1978:IER

[894] W. David Elliott. Index to the Euclid report. ACM SIGPLAN Notices, 13 (3):85–89, March 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Estell:1978:CHD

[895] Robert G. Estell. A chapter in the history of DOD-1. ACM SIG-PLAN Notices, 13(3):90–92, March 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baskett:1978:BSC

[896] Forest Baskett. The best simple code generation technique for WHILE, FOR, and DO loops. ACM SIGPLAN Notices, 13(4):31–32, April 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Burda:1978:TBE

[897] John C. Burda. The translation of Boolean: expressions into optimal transfer logic code for the OLIVER data base management system compiler. ACM SIGPLAN Notices, 13 (4):33–43, April 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chevance:1978:SPD

[898] R. J. Chevance and T. Heidet. Static profile and dynamic behavior of COBOL programs. ACM SIGPLAN Notices, 13(4):44–57, April 1978. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clifton:1978:TMS

[899] Mitchell H. Clifton. A technique for making structured programs more readable. ACM SIGPLAN Notices, 13 (4):58–63, April 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jones:1978:NSL

[900] Douglas W. Jones. A note on some limits of the algebraic specification method. ACM SIGPLAN Notices, 13 (4):64–67, April 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jacobsen:1978:AVC

[901] Tom Jacobsen. Another view of coroutines. ACM SIGPLAN Notices, 13 (4):68–75, April 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Slater:1978:SPC

[902] William Slater and Howard Modell. Structured programming considered harmful. ACM SIGPLAN Notices, 13(4):76–79, April 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Subrahmanyam:1978:FAD

[903] P. A. Subrahmanyam. On a finite axiomatization of the data type L. ACM SIGPLAN Notices, 13(4):80–84, April 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Peschke:1978:CPR

[904] J. V. Peschke. Comment on a paper of R. A. Freiburghouse. ACM SIGPLAN Notices, 13(4):85–88, April 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [799].

Wexelblat:1978:NBa

[905] Richard L. Wexelblat and Debra D. Child. New books. ACM SIG-PLAN Notices, 13(5):7–8, May 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Battarel:1978:RSP

[906] G. J. Battarel and R. J. Chevance. Requirements for a safe PL/1 implementation. ACM SIGPLAN Notices, 13 (5):12–22, May 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bezivin:1978:AVC

[907] Jean Bezivin, Jean-Louis Nebut, and Robert Rannou. Another view of coroutines. ACM SIGPLAN Notices, 13(5):23–35, May 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chapin:1978:CSP

[908] Ned Chapin and Susan P. Denniston. Characteristics of a structured program. ACM SIGPLAN Notices, 13 (5):36–45, May 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ditzel:1978:PMH

[909] D. R. Ditzel. Pattern matching for high level languages. ACM SIG-PLAN Notices, 13(5):46–55, May 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldman:1978:ESM

[910] Michael B. Feldman. Embedding sparse matrix operations in SNOBOL4. ACM SIGPLAN Notices, 13(5):56–64, May 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holt:1978:ACS

[911] R. C. Holt, G. S. Graham, E. D. Lazowska, and M. A. Scott. Announcing CONCURRENT SP/k. ACM SIG-PLAN Notices, 13(5):65–68, May 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Alleman:1978:PVS

[912] Glen B. Alleman. Parameter validation for software reliability. ACM SIG-PLAN Notices, 13(6):18–28, June 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elshoff:1978:SSC

[913] James L. Elshoff. A study of the structural composition of PL/I programs. ACM SIGPLAN Notices, 13(6):29–37, June 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gorski:1978:SVM

[914] Janusz Górski. Sticky variables and memory optimization. ACM SIG-PLAN Notices, 13(6):38–48, June 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schneider:1978:PSE

[915] Victor Schneider. Prediction of software effort and project duration: four new formulas. ACM SIGPLAN Notices, 13(6):49–59, June 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1978:NBb

[916] R. L. Wexelblat and Debbie Child. New books. ACM SIGPLAN Notices, 13 (7):14–17, July 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Allison:1978:NCS

[917] Lloyd Allison. On nondeterministic control structures. ACM SIG-PLAN Notices, 13(7):18–20, July 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1978:DS

[918] Edsger W. Dijkstra. DoD-I: the summing up. ACM SIGPLAN Notices, 13 (7):21–26, July 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ford:1978:PET

[919] B. Ford. Parameterisation of the environment for transportable numerical software. ACM SIGPLAN Notices, 13 (7):27–30, July 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lewis:1978:FCV

[920] Brian Lewis. Further comments on “A view of coroutines”. ACM SIG-PLAN Notices, 13(7):31–33, July 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [803, 907, 901].

Mohilner:1978:PPP

[921] Patricia R. Mohilner. Prettyprinting PASCAL programs. ACM SIG-PLAN Notices, 13(7):34–40, July 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Parkin:1978:UKP

[922] Rodney Parkin. On the use of keywords for passing procedure parameters. ACM SIGPLAN Notices, 13 (7):41–42, July 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Postma:1978:FFE

[923] Stef. W. Postma. FEA: a formal English subset for Algebra/Assertions. ACM SIGPLAN Notices, 13(7):43–59, July 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gries:1978:ASH

[924] David Gries. ACM SIGPLAN history of programming languages conference ALGOL 60 language summary. ACM SIGPLAN Notices, 13(8):1, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perlis:1978:ASD

[925] Alan J. Perlis. The American side of the development of Algol. ACM SIGPLAN Notices, 13(8):3–14, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Naur:1978:ESL

[926] Peter Naur. The European side of the last phase of the development of AL-GOL 60. ACM SIGPLAN Notices, 13 (8):15–44, August 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1978:ALS

[927] Anonymous. APL language summary. ACM SIGPLAN Notices, 13(8):45, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Falkoff:1978:EA

[928] Adin D. Falkoff and Kenneth E. Iverson. The evolution of APL. ACM SIGPLAN Notices, 13(8):47–57, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1978:ASHa

[929] Anonymous. ACM SIGPLAN history of programming languages conference APT language summary. ACM SIG-PLAN Notices, 13(8):59–60, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ross:1978:OAL

[930] Douglas T. Ross. Origins of the APT language for automatically programmed tools. ACM SIGPLAN Notices, 13(8):61–99, August 1978. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lewis:1978:ASH

[931] Ted G. Lewis. ACM SIGPLAN history of programming languages conference BASIC language summary. ACM SIG-PLAN Notices, 13(8):101–102, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kurtz:1978:B

[932] Thomas E. Kurtz. BASIC. ACM SIG-PLAN Notices, 13(8):103–118, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1978:ASHb

[933] Anonymous. ACM SIGPLAN history of programming languages conference COBOL 60 language summary. ACM SIGPLAN Notices, 13(8):119, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sammet:1978:EHC

[934] Jean E. Sammet. The early history of COBOL. ACM SIGPLAN Notices, 13(8):121–161, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Galler:1978:FL

[935] Bernard A. Galler. The FORTRAN Language. ACM SIGPLAN Notices, 13(8):163–164, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Backus:1978:HFI

[936] John Backus. The history of FOR-TRAN I, II, and III. ACM SIG-PLAN Notices, 13(8):165–180, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). URL http:

//community.computerhistory.org/
scc/projects/FORTRAN/paper/p165-
backus.pdf.

Anonymous:1978:ASHc

[937] Anonymous. ACM SIGPLAN history of programming languages conference GPSS language summary. ACM SIG-PLAN Notices, 13(8):181–182, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gordon:1978:DGP

[938] Geoffrey Gordon. The development of the General Purpose Simulation System (GPSS). ACM SIGPLAN Notices, 13(8):183–198, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Baker:1978:ASH

[939] Charles L. Baker. ACM SIGPLAN history of programming languages conference JOSS language summary. ACM SIGPLAN Notices, 13(8):199–200, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cheatham:1978:BDJ

[940] T. E. Cheatham, Jr. A brief description of JOVIAL. ACM SIGPLAN Notices, 13(8):201–202, August 1978.

CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schwartz:1978:DJ

[941] Jules I. Schwartz. The development of JOVIAL. ACM SIGPLAN Notices, 13(8):203–214, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCarthy:1978:MML

[942] John McCarthy. A micro-manual for LISP — not the whole truth. ACM SIGPLAN Notices, 13(8):215–216, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McCarthy:1978:HL

[943] John McCarthy. History of LISP. ACM SIGPLAN Notices, 13(8):217–223, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosin:1978:ASH

[944] Robert F. Rosin. ACM SIGPLAN history of programming languages conference PL/I language summary. ACM SIGPLAN Notices, 13(8):225–226, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Radin:1978:EHC

[945] George Radin. The early history and characteristics of PL/I. ACM SIG-PLAN Notices, 13(8):227–241, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Franta:1978:ASH

[946] W. R. Franta. ACM SIGPLAN history of programming languages conference SIMULA language summary. ACM SIGPLAN Notices, 13(8):243–244, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nygaard:1978:DSL

[947] Kristen Nygaard and Ole-Johan Dahl. The development of the SIMULA languages. ACM SIGPLAN Notices, 13 (8):245–272, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shapiro:1978:ASH

[948] Michael D. Shapiro. ACM SIGPLAN history of programming languages conference SNOBOL language summary. ACM SIGPLAN Notices, 13(8):273– 274, August 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griswold:1978:HSP

[949] Ralph E. Griswold. A history of the SNOBOL programming languages. ACM SIGPLAN Notices, 13(8):275– 308, August 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1978:A

[950] R. L. Wexelblat. Afterword. ACM SIG-PLAN Notices, 13(8):309–310, August 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahams:1978:TPP

[951] Paul W. Abrahams. Two programming problems. ACM SIGPLAN Notices, 13(9):28–29, September 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LeBlanc:1978:EPS

[952] Richard J. LeBlanc. Extensions to PASCAL for separate compilation. ACM SIGPLAN Notices, 13(9):30– 33, September 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Maurer:1978:RTB

[953] W. D. Maurer. Register type bits — a proposal for efficient instruction repertoire extension. ACM SIGPLAN Notices, 13(9):34–35, September 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1978:LDIa

[954] Mary Shaw, Paul Hilfinger, and Wm. A. Wulf. Language design for the Ironman requirement: reference manual. ACM SIGPLAN Notices, 13(9): 36–58, September 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1978:TLDa

[955] M. Shaw, P. Hilfinger, and W. A. Wulf. TARTAN-language design for the Ironman requirement: reference manual. ACM SIGPLAN Notices, 13(9):36– 58, September 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1978:LDIb

[956] Mary Shaw, Paul Hilfinger, and Wm. A. Wulf. Language design for the Ironman requirement: notes and examples. ACM SIGPLAN Notices, 13(9): 59–75, September 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shaw:1978:TLDb

[957] M. Shaw, P. Hilfinger, and W. A. Wulf. TARTAN-language design for the Ironman requirement: notes and examples. ACM SIGPLAN Notices, 13(9): 59–75, September 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Skordalakis:1978:CF

[958] E. Skordalakis and G. Papakonstantinou. Coroutines in FORTRAN. ACM SIGPLAN Notices, 13(9):76– 84, September 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1978:BLS

[959] E. W. Dijkstra. On the BLUE language submitted to the DoD. ACM SIG-PLAN Notices, 13(10):10–15, October 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1978:GLS

[960] E. W. Dijkstra. On the GREEN language submitted to the DoD. ACM SIGPLAN Notices, 13(10):16–21, October 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1978:YLS

[961] E. W. Dijkstra. On the YELLOW language submitted to the DoD. ACM SIGPLAN Notices, 13(10):22–26, October 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dijkstra:1978:RLS

[962] E. W. Dijkstra. On the RED language submitted to the DoD. ACM SIG-PLAN Notices, 13(10):27–32, October 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

James:1978:FM

[963] John S. James. FORTH for microcomputers. ACM SIGPLAN Notices, 13(10):33–39, October 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lindamood:1978:SDC

[964] George E. Lindamood. Some delitescent concepts in computer architecture. ACM SIGPLAN Notices, 13(10): 40–41, October 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marti:1978:MRT

[965] Jed Marti. The META/REDUCE translator writing system. ACM SIG-PLAN Notices, 13(10):42–49, October 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zucconi:1978:CCD

[966] Linda Zucconi. A comment on Chapin’s and Denniston’s ‘Characteristics of a structured program’. ACM SIGPLAN Notices, 13(10):50–51, October 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1978:NBI

[967] R. L. Wexelblat. New books III. ACM SIGPLAN Notices, 13(11):8– 10, November 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crider:1978:SFP

[968] John E. Crider. Structured formatting of Pascal programs. ACM SIGPLAN Notices, 13(11):15–22, November 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elliott:1978:OCS

[969] Irving B. Elliott. Optimization of the compiling of subscripted array references. ACM SIGPLAN Notices, 13(11): 23–29, November 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kovats:1978:PRC

[970] T. A. Kovats. Program readability. closing keywords and prefix-style intermediate keywords. ACM SIGPLAN Notices, 13(11):30–42, November 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pflasterer:1978:CPL

[971] David C. Pflasterer. The CAMIL programming language. ACM SIGPLAN Notices, 13(11):43–55, November 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sammet:1978:RPL

[972] Jean E. Sammet. Roster of programming languages for 1976–77. ACM SIGPLAN Notices, 13(11):56– 85, November 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bauer:1978:TWS

[973] F. L. Bauer, M. Broy, R. Gnatz, W. Hesse, B. Krieg-Brückner, H. Partsch, P. Pepper, and H. Wössner. Towards a wide spectrum language to support program specification and program development. ACM SIGPLAN Notices, 13(12):15–24, December 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chaudhary:1978:SAS

[974] B. D. Chaudhary and H. V. Sahasrabuddhe. Suggestions about a specification technique (program). ACM SIGPLAN Notices, 13(12):25– 28, December 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elshoff:1978:UCN

[975] James L. Elshoff and Michael Marcotty. On the use of the cyclomatic number to measure program complexity. ACM SIGPLAN Notices, 13(12):29–40, December 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lauer:1978:ASR

[976] P. E. Lauer and M. W. Shields. Abstract specification of resource accessing disciplines: adequacy, starvation, priority and interrupts (programming theory). ACM SIGPLAN Notices, 13(12):41–59, December 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Montgomery:1978:EBE

[977] Peter L. Montgomery. Evaluation of Boolean expressions on one’s complement machines. ACM SIGPLAN Notices, 13(12):60–72, December 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nelson:1978:XLS

[978] R. A. Nelson. XL: a language structure. ACM SIGPLAN Notices, 13(12): 73–87, December 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sinha:1978:EEB

[979] V. Sinha and K. Vinod Kumar. Efficient evaluation of Boolean expressions. ACM SIGPLAN Notices, 13(12): 88–97, December 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sites:1978:PTS

[980] Richard L. Sites. Programming tools: statement counts and procedure timings. ACM SIGPLAN Notices, 13(12): 98–101, December 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1978:SFL

[981] Peter Wegner. Supporting a flourishing language culture. ACM SIGPLAN Notices, 13(12):102–104, December 1978. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Arisawa:1979:FPU

[982] Makoto Arisawa and Minoru Iuchi. Fortran + preprocessor = Utopia 84. ACM SIGPLAN Notices, 14(1):12–15, January 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ciesinger:1979:BEH

[983] J. Ciesinger. A bibliography of errorhandling. ACM SIGPLAN Notices, 14 (1):16–26, January 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hunt:1979:MTL

[984] J. G. Hunt. Messages in typed languages. ACM SIGPLAN Notices, 14 (1):27–45, January 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LaLonde:1979:FSA

[985] Wilf R. LaLonde. A finite state approach to operator classification. ACM SIGPLAN Notices, 14(1):46–54, January 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Low:1979:SNS

[986] James R. Low. A short note on scanning signed integers. ACM SIG-PLAN Notices, 14(1):55–56, January 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Metzner:1979:GBM

[987] John R. Metzner. A graded bibliography on macro systems and extensible languages. ACM SIGPLAN Notices, 14 (1):57–68, January 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moya:1979:SGC

[988] R. Moya and I. Ramos. SP-grammars (context-free grammars for structured programming). ACM SIGPLAN Notices, 14(1):69–76, January 1979. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sakoda:1979:DGP

[989] James M. Sakoda. DYSTAL 2: a general purpose extension of FORTRAN. ACM SIGPLAN Notices, 14(1):77–90, January 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sommerville:1979:SSS

[990] Ian Sommerville. S-SNOBOL: structured SNOBOL4. ACM SIGPLAN Notices, 14(1):91–99, January 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Tai:1979:IPT

[991] Kuo-Chung Tai. On the implementation of parsing tables. ACM SIG-PLAN Notices, 14(1):100–101, January 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Backhouse:1979:LCR

[992] R. C. Backhouse and S. Anderson. Least-cost repair of syntax errors. ACM SIGPLAN Notices, 14(1):102–104, January 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clark:1979:IIP

[993] Robert G. Clark. Interactive input in PASCAL. ACM SIGPLAN Notices, 14 (2):9–13, February 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elliott:1979:DDL

[994] Irving B. Elliott. DLP: a design language pre-processor. ACM SIGPLAN Notices, 14(2):14–20, February 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Jackson:1979:TDG

[995] Glenn A. Jackson. Two-dimensional grammars and structured programming languages. ACM SIGPLAN Notices, 14(2):21–28, February 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kling:1979:DCH

[996] Rob Kling and Walter Scacchi. The DoD common high order programming language effort (DoD-1): What will the impacts be? ACM SIGPLAN Notices, 14(2):29–43, February 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosen:1979:SL

[997] Barry K. Rosen. Stickiness and liveness. ACM SIGPLAN Notices, 14(2): 44–45, February 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turchin:1979:SSB

[998] V. F. Turchin. A supercompiler system based on the language REFAL. ACM SIGPLAN Notices, 14(2):46–54, February 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Veloso:1979:TSFa

[999] Paulo A. S. Veloso. Traversable stack with fewer errors. ACM SIGPLAN Notices, 14(2):55–59, February 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Watt:1979:EAG

[1000] David A. Watt. An extended attribute grammar for Pascal. ACM SIGPLAN Notices, 14(2):60–74, February 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Axford:1979:ASL

[1001] T. H. Axford, Diana Burkhardt, W. P. Dodd, Susan Laflin, D. G. Parkyn, and P. Ramsay. ATOL: a simple language with powerful data structuring facilities. ACM SIGPLAN Notices, 14 (3):5–15, March 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Elliot:1979:DDL

[1002] I. B. Elliot. DLP, a design language pre-processor. ACM SIGPLAN Notices, 14(3):14–20, February 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chapin:1979:FRF

[1003] Ned Chapin. Full report of the Flowchart Committee on ANS Standard X3.5-1970. ACM SIGPLAN Notices, 14(3):16–27, March 1979. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kamin:1979:SDA

[1004] Sam Kamin. Some definitions for algebraic data type specifications. ACM SIGPLAN Notices, 14(3):28–37, March 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Lamport:1979:GCE

[1005] L. Lamport. A general construction for expressing repetition. ACM SIGPLAN Notices, 14(3):38–42, March 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rosen:1979:SLP

[1006] B. K. Rosen. Stickiness and liveness (program variables). ACM SIGPLAN Notices, 14(3):44–45, February 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahams:1979:MCP

[1007] Paul Abrahams. The map coloring problem. ACM SIGPLAN Notices, 14 (4):10–11, April 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Culik:1979:CNN

[1008] Karel Culik. The cyclomatic number and the normal number of programs. ACM SIGPLAN Notices, 14(4):12–17, April 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Griswold:1979:IPL

[1009] Ralph E. Griswold, David R. Hanson, and John T. Korb. The Icon programming language: an overview. ACM SIGPLAN Notices, 14(4):18–31, April 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Glanc:1979:ECD

[1010] Alois Glanc. On expressing control and data structures in predicate logic language ALGOL M. ACM SIGPLAN Notices, 14(4):32–34, April 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grogono:1979:LIS

[1011] Peter Grogono. On layout, identifiers and semicolons in Pascal programs. ACM SIGPLAN Notices, 14(4):35–40, April 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Newton:1979:EHI

[1012] Roy Newton. On exception handling with the ‘IRONMAN’ languages, GREEN and REDL. ACM SIGPLAN Notices, 14(4):41–54, April 1979. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Newton:1979:SEH

[1013] Roy Newton. Some exception handling problems in language systems displaying a multi-path capability. ACM SIGPLAN Notices, 14(4):55–63, April 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schulz:1979:IFC

[1014] Dean Schulz. The Intel Fortran-80 compiler for the 8080 and 8085 microprocessors. ACM SIGPLAN Notices, 14 (4):64–76, April 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zabek:1979:DCS

[1015] Swiatomir G. Zabek. Decomposition of complex selection criteria. (programming). ACM SIGPLAN Notices, 14 (4):77–80, April 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gini:1979:MTS

[1016] Giuseppina Gini, Maria Gini, Renato Gini, and Dario Giuse. A multi-task system for robot programming. ACM SIGPLAN Notices, 14(5):11–18, May 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Anonymous:1979:ACC

[1017] Anonymous. APL comparison chart.

ACM SIGPLAN Notices, 14(5):19–20, May 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chinnaswamy:1979:TDT

[1018] V. Chinnaswamy. Translation of decision tables into computer programs with the use of Carnaugh maps. ACM SIGPLAN Notices, 14(5):21–23, May 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Deransart:1979:LLD

[1019] P. Deransart. The language LISP does not exist? ACM SIGPLAN Notices, 14(5):24–27, May 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeRemer:1979:SCS

[1020] Frank DeRemer, Philip Levy, Steve Hanson, Philip Jackson, Richard Jullig, and Tom Pittman. Summary of the characteristics of several ‘modern’ programming languages. ACM SIG-PLAN Notices, 14(5):28–45, May 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hanson:1979:SCS

[1021] S. Hanson, P. Jackson, R. Jullig, P. Levy, and T. Pittman. Summary of the characteristics of several ‘modern’ programming languages. ACM SIG-PLAN Notices, 14(5):28–45, May 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kapur:1979:SMT

[1022] Deepak Kapur. Specifications of Majster’s traversable stack and Veloso’s traversable stack. ACM SIGPLAN Notices, 14(5):46–53, May 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Smart:1979:MEP

[1023] Robert K. Smart. Micro Extensible Programming Language. ACM SIG-PLAN Notices, 14(5):54–63, May 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ichbiah:1979:PAR

[1024] Jean D. Ichbiah. Preliminary Ada reference manual. ACM SIGPLAN Notices, 14(6a):1–145, June 1979. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The final standard is [?].

Ichbiah:1979:RDA

[1025] Jean D. Ichbiah, Bernd Krieg-Brueckner, Brian A. Wichmann, John G. P. Barnes, Olivier Roubine, and Jean-Claude Heliard. Rationale for the design of the Ada programming language. ACM SIGPLAN Notices, 14 (6b):1–261, June 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). The final standard is [?].

Denvir:1979:OPL

[1026] B. T. Denvir. On orthogonality in programming languages. ACM SIG-PLAN Notices, 14(7):18–30, July 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fosdick:1979:HLL

[1027] Howard Fosdick. High-level languages for operating systems development — a historical overview. ACM SIG-PLAN Notices, 14(7):31–37, July 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Grune:1979:SSA

[1028] Dick Grune. Some statistics on ALGOL 68 programs. ACM SIG-PLAN Notices, 14(7):38–46, July 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hennell:1979:RPP

[1029] M. A. Hennell. The role of procedural parameters in high order languages. ACM SIGPLAN Notices, 14(7):47–50, July 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

MacLennan:1979:ODB

[1030] B. J. MacLennan. Observations on the differences between formulas and sentences and their application to programming language design. ACM SIG-PLAN Notices, 14(7):51–61, July 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

McGettrick:1979:AAM

[1031] Andrew D. McGettrick. Aspects of the ALGOL 68 mode structure. ACM SIGPLAN Notices, 14(7):62–77, July 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Orejas:1979:PCS

[1032] Fernando Orejas. On the power of conditional specifications. ACM SIG-PLAN Notices, 14(7):78–81, July 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Prakash:1979:IDB

[1033] Naveen Prakash and Y. K. Sharma. An integrated data base language. ACM SIGPLAN Notices, 14(7):82–96, July 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Morris:1979:DAS

[1034] James B. Morris. Data abstraction: A static implementation strategy. ACM SIGPLAN Notices, 14(8):1–7, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Holt:1979:MIE

[1035] Richard C. Holt and David B. Wortman. A model for implementing Euclid modules and type templates. ACM SIGPLAN Notices, 14(8):8–12, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Williams:1979:PC

[1036] Graeme Williams. Program checking.

ACM SIGPLAN Notices, 14(8):13–25, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Raiha:1979:DAS

[1037] Kari-Jouko Räihä. Dynamic allocation of space for attribute instances in multi-pass evaluators of attribute grammars. ACM SIGPLAN Notices, 14 (8):26–38, August 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cordy:1979:SCD

[1038] James R. Cordy, Richard C. Holt, and David B. Wortman. Semantic charts: A diagrammatic approach to semantic processing. ACM SIG-PLAN Notices, 14(8):39–49, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:1979:HST

[1039] Susan L. Graham, William N. Joy, and O. Roubine. Hashed symbol tables for languages with explicit scope control. ACM SIGPLAN Notices, 14(8):50–57, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Donegan:1979:CGG

[1040] Michael K. Donegan, Robert E. Noonan, and Stefan Feyock. A codegenerator generator language. ACM SIGPLAN Notices, 14(8):58–64, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cattell:1979:CGM

[1041] Roderic G. G. Cattell, Joseph M. Newcomer, and Bruce W. Leverett. Code generation in a machine-independent compiler. ACM SIGPLAN Notices, 14 (8):65–75, August 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kahrs:1979:IIP

[1042] Mark Kahrs. Implementation of an Interactive Programming System. ACM SIGPLAN Notices, 14(8):76–82, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fabri:1979:ASO

[1043] Janet Fabri. Automatic storage optimization. ACM SIGPLAN Notices, 14 (8):83–91, August 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bonkowski:1979:PZC

[1044] G. B. Bonkowski, W. M. Gentleman, and M. A. Malcolm. Porting the Zed compiler. ACM SIGPLAN Notices, 14 (8):92–97, August 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feldman:1979:IPF

[1045] Stuart I. Feldman. Implementation of a portable Fortran 77 compiler using modern tools. ACM SIGPLAN Notices, 14(8):98–106, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahams:1979:CPC

[1046] Paul W. Abrahams. The CIMS PL/I compiler. ACM SIGPLAN Notices, 14 (8):107–116, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hansen:1979:CTM

[1047] G. J. Hansen, G. A. Shoults, and J. D. Cointment. Construction of a transportable, multi-pass compiler for extended Pascal. ACM SIGPLAN Notices, 14(8):117–126, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rudmik:1979:CDE

[1048] A. Rudmik and E. S. Lee. Compiler design for efficient code generation and program optimization. ACM SIG-PLAN Notices, 14(8):127–138, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

LeBlanc:1979:ISC

[1049] Richard J. LeBlanc and Charles N. Fischer. On implementing separate compilation in block-structured languages. ACM SIGPLAN Notices, 14(8):139– 143, August 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pleban:1979:UTM

[1050] Uwe F. Pleban. The use of transition matrices in a recursive-descent compiler. ACM SIGPLAN Notices, 14(8):

144–151, August 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Milton:1979:ACG

[1051] D. R. Milton, L. W. Kirchhoff, and B. R. Rowland. An ALL(1) compiler generator. ACM SIGPLAN Notices, 14(8):152–157, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pai:1979:GCR

[1052] Ajit B. Pai and Richard B. Kieburtz. Global context recovery: A new strategy for parser recovery from syntax errors. ACM SIGPLAN Notices, 14(8): 158–167, August 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:1979:PLE

[1053] Susan L. Graham, Charles B. Haley, and William N. Joy. Practical LR error recovery. ACM SIGPLAN Notices, 14(8):168–175, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

DeRemer:1979:ECL

[1054] Frank DeRemer and Thomas J. Pennello. Efficient computation of LALR(1) look-ahead sets. ACM SIG-PLAN Notices, 14(8):176–187, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Frailey:1979:ILS

[1055] Dennis J. Frailey. An intermediate language for source and target independent code optimization. ACM SIG-PLAN Notices, 14(8):188–200, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perkins:1979:MIP

[1056] Daniel R. Perkins and Richard L.

Sites. Machine-independent PASCAL code optimization. ACM SIGPLAN Notices, 14(8):201–207, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nelson:1979:CPI

[1057] Philip A. Nelson. A comparison of PASCAL intermediate languages. ACM SIGPLAN Notices, 14(8):208– 213, August 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ball:1979:PEO

[1058] J. Eugene Ball. Predicting the effects of optimization on a procedure body (program compilers). ACM SIGPLAN Notices, 14(8):214–220, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sites:1979:MIR

[1059] Richard L. Sites. Machine-independent register allocation. ACM SIGPLAN Notices, 14(8):221–225, August 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mintz:1979:DGO

[1060] Ronald J. Mintz, Gerald A. Fisher, Jr., and Micha Sharir. The design of a global optimizer (program compilers). ACM SIGPLAN Notices, 14(8): 226–234, August 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Abrahams:1979:PSA

[1061] Paul Abrahams. The PL/I standard: an apostate’s view. ACM SIGPLAN Notices, 14(9):15–16, September 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ball:1979:PZL

[1062] J. E. Ball, G. J. Williams, and J. R. Low. Preliminary ZENO language description. ACM SIGPLAN Notices, 14(9):17–34, September 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berry:1979:TES

[1063] Daniel M. Berry and Richard L. Schwartz. Type equivalence in strongly typed languages: one more look. ACM SIGPLAN Notices, 14(9):35– 41, September 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gustafson:1979:SPE

[1064] G. G. Gustafson. Some practical experiences formatting Pascal programs. ACM SIGPLAN Notices, 14(9):42– 49, September 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Pflasterer:1979:ARC

[1065] David C. Pflasterer. Application of a remote code generator. ACM SIG-PLAN Notices, 14(9):50–58, September 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Redwine:1979:UCM

[1066] Samuel T. Redwine, Jr. Using COBOL macroprocessing for reliable parameter passing. ACM SIGPLAN Notices, 14(9):59–60, September 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Sale:1979:SP

[1067] Arthur Sale. Scope and Pascal.

ACM SIGPLAN Notices, 14(9):61– 63, September 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Shoch:1979:OPL

[1068] John F. Shoch. An overview of the programming language Smalltalk-72. ACM SIGPLAN Notices, 14(9): 64–73, September 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Ramsdell:1979:PSP

[1069] John Ramsdell. Prettyprinting structured programs with connector lines. ACM SIGPLAN Notices, 14(9):74– 75, September 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wallace:1979:MII

[1070] B. Wallace. More on interactive input in PASCAL. ACM SIGPLAN Notices, 14(9):76–??, September 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Cohen:1979:TDT

[1071] Shimon Cohen. The A-TABLE datatype for LISP systems. ACM SIG-PLAN Notices, 14(10):36–47, October 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Marti:1979:SLR

[1072] J. B. Marti, A. C. Hearn, M. L. Griss, and C. Griss. Standard Lisp Report. ACM SIGPLAN Notices, 14(10): 48–68, October 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schonberg:1979:HDE

[1073] Edmond Schonberg. Hierarchical design and efficient implementation in SETL: A case study. ACM SIGPLAN Notices, 14(10):69–74, October 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Teitelbaum:1979:CPS

[1074] Tim Teitelbaum. The Cornell Program Synthesizer: a syntax-directed programming environment. ACM SIG-PLAN Notices, 14(10):75, October 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Veloso:1979:TSFb

[1075] Paulo A. S. Veloso. Traversable stack with fewer errors: addenda and corrigenda. ACM SIGPLAN Notices, 14 (10):76, October 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Barnes:1979:WDP

[1076] R. Barnes. A working definition of the proposed extensions for PL/I realtime applications. ACM SIGPLAN Notices, 14(10):77–99, October 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bergstra:1979:AFE

[1077] J. A. Bergstra and J. V. Tucker. On the adequacy of finite equational methods for data type specification. ACM SIG-PLAN Notices, 14(11):13–18, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Crowley:1979:NPT

[1078] Charles Crowley and Gary Klimowicz.

A note on procedure timing. ACM SIG-PLAN Notices, 14(11):19–22, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Feign:1979:NLO

[1079] David Feign. A note on loop ‘optimization’. ACM SIGPLAN Notices, 14(11): 23–25, November 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Fisher:1979:LPL

[1080] Gerald A. Fisher, Jr. and Manfred Weber. LALR(1) parsing for languages without reserved words. ACM SIG-PLAN Notices, 14(11):26–30, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Dungan:1979:BDT

[1081] D. M. Dungan. Bibliography on data types. ACM SIGPLAN Notices, 14(11):31–59, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Keller:1979:SRC

[1082] R. F. Keller. On selective and repetitive constructs. ACM SIGPLAN Notices, 14(11):60–63, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laski:1979:ALP

[1083] Janusz W. Laski. An approach to loop programs debugging. ACM SIGPLAN Notices, 14(11):64–72, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Laski:1979:RPL

[1084] Janusz W. Laski. On readability of programs with loops. ACM SIGPLAN Notices, 14(11):73–83, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Nead:1979:SCS

[1085] James M. Nead. On the semantics of control statements. ACM SIGPLAN Notices, 14(11):84–96, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Runciman:1979:SVP

[1086] Colin Runciman. Scarcely variabled programming and Pascal. ACM SIG-PLAN Notices, 14(11):97–106, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Salvadori:1979:SIP

[1087] Antonio Salvadori and Christine Dumont. Some inconsistencies in programming language implementation. ACM SIGPLAN Notices, 14(11):107– 109, November 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Srinivasan:1979:CFC

[1088] B. Srinivasan and V. Gopalakrishna. Control flow complexity and structuredness of programs. ACM SIGPLAN Notices, 14(11):110–115, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See comment [1101].

Zoethout:1979:WGH

[1089] Th. A. Zoethout and E. I. N. Stein.

Why the goto is harmful. ACM SIGPLAN Notices, 14(11):116–122, November 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wegner:1979:PAI

[1090] P. Wegner. Programming with ADA:

an introduction by means of graduated examples. ACM SIGPLAN Notices, 14(12):1–46, December 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bond:1979:ANP

[1091] R. Bond. Another note on Pascal indention. ACM SIGPLAN Notices, 14(12):47–49, December 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hearn:1979:OPP

[1092] Anthony C. Hearn and Arthur C. Norman. A one-pass prettyprinter. ACM SIGPLAN Notices, 14(12):50–58, December 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bron:1979:DPI

[1093] C. Bron and E. J. Dijkstra. A discipline for the programming of interactive I/O in Pascal. ACM SIGPLAN Notices, 14(12):59–61, December 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Soong:1979:OOI

[1094] Norman L. Soong. An overview of an organized industrial research and development activity for language and language processors. ACM SIGPLAN Notices, 14(12):62–63, December 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Gyllstrom:1979:UCS

[1095] H. C. Gyllstrom, R. C. Knippel, L. C. Ragland, and K. E. Spackman. The Universal Compiling System. ACM SIGPLAN Notices, 14(12):64–70, December 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Heacox:1979:RLS

[1096] H. C. Heacox. RDL: a language for software development. ACM SIGPLAN Notices, 14(12):71–79, December 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Greenwood:1979:MPL

[1097] Stephen R. Greenwood. MACRO: a programming language. ACM SIG-PLAN Notices, 14(12):80–91, December 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Turba:1979:GSA

[1098] Thomas N. Turba. General Syntax Analyzer (GSA). ACM SIGPLAN Notices, 14(12):92–109, December 1979. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hart:1979:AID

[1099] Jolene J. Hart. The Advanced Interactive Debugging System (AIDS). ACM SIGPLAN Notices, 14(12):110– 121, December 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Rohde:1979:LTD

[1100] Wm. F. Rohde. Language tools for data access: past, present and future. ACM SIGPLAN Notices, 14(12):122– 141, December 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Zucconi:1980:CSF

[1101] Linda Zucconi. A comment on “Control flow complexity and structuredness of programs” by Srinivasan and Gopalakrishna. ACM SIGPLAN Notices, 15(2):82–84, February 1980. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). See [1088].

ACM:1971:PAS

[1102] ACM, editor. Proceedings of the ACM SIGPLAN symposium on languages for systems implementation. Lafayette, IN, USA, 25–26 October, 1971, volume 6(9) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, October 1971. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1973:PAS

[1103] ACM, editor. Proceedings of ACM SIGPLAN-SIGTOPS Interface Meeting on Programming Languages-Operating Systems. Savannah, GA, USA, 9–12 April, 1973, volume 8(9) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, September 1973. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Graham:1973:PAS

[1104] Robert M. Graham and Michael D. Schroeder, editors. Proceedings of ACM SIGPLAN-SIGOPS Interface Meeting: programming languages, operating systems: Savannah, Georgia, April 9–12, 1973, volume 8(9) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1973. LCCN QA 76.7 A1095 v.8 no.9.

ACM:1974:ASS

[1105] ACM, editor. ACM SIGPLAN Symposium on Very High Level Languages. Santa Monica, CA, USA, 28– 29 March, 1974, volume 9(4) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, April 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7.S53 1992.

ACM:1974:PAS

[1106] ACM, editor. Proceedings of the ACM SIGPLAN-SIGMICRO interface meeting on programming languages-Microprogramming. Harriman, NY, USA, 30 May–1 June, 1973, volume 9(8) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, August 1974. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Wexelblat:1974:PAS

[1107] Richard L. Wexelblat, editor. Proceedings of ACM SIGPLAN-SIGMICRO Interface Meeting: programming languages, microprogramming: Arden House, Harriman, New York, May 30– June 1, 1973, volume 9(8) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1974.

Nance:1974:PAS

[1108] Richard E. Nance, editor. Proceedings of ACM SIGPLAN-SIGIR Interface Meeting: programming languages, information retrieval: Gaithersburg, Maryland, November 4–6, 1973, volume 9(3) of ACM SIGIR forum. ACM Press, New York, NY, USA, 1974. LCCN QA 76.7 A1095 v.10 no.1.

ACM:1975:ICR

[1109] ACM, editor. International Conference on Reliable Software. Los Angeles, CA, USA, 21–23 April, 1975, volume 10(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1975:PAS

[1110] ACM, editor. Proceedings of the ACM SIGPLAN-SIGIR interface meeting on programming languages-information retrieval. Gaithersburg, MD, USA, 4– 6 November 1973, volume 10(1) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, January 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1975:PLC

[1111] ACM, editor. Programming Languages and Compilers for Parallel and Vector Machines. New York, NY, USA, 18–19 March, 1975, volume 10(3) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, March 1975. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1976:PCD

[1112] ACM, editor. Proceedings of Conference on Data: Abstraction, Definition and Structure: March 22–24, 1976, Salt Lake City, Utah, volume 11(3S) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, March 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.9.D35 C65 1976.

ACM:1976:ASS

[1113] ACM, editor. ACM SIGMIN/ SIGPLAN Interface Meeting on Programming Systems in the Small Processor Environment. New Orleans, LA, USA, 4–6 March, 1976, volume 11(4) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, April 1, 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1976:SGL

[1114] ACM, editor. Symposium on Graphic Languages: Miami Beach, FL, USA, 26–27 April, 1976, volume 11(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1976. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Berk:1976:GLP

[1115] Toby Berk, editor. Graphic languages: papers presented at a symposium sponsored by ACM SIGPLAN, ACM SIGGRAPH, Florida International University, April 26–27, 1976, volume 11(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, 1976. LCCN T 385 C739 v.10 no.1 1976. Cover title: Proceedings. Spine title: Proceedings of ACM SIGPLAN/SIGGRAPH Symposium on Graphic Languages, 26–27 Apr. 1976.

ACM:1977:ASA

[1116] ACM, editor. ACM Symposium in Artificial Intelligence and Programming Languages. Rochester, NY, USA, 15– 17 August, 1977, volume 12(8) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, August 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1977:PAC

[1117] ACM, editor. Proceedings of an ACM Conference on Language Design for Reliable Software. Raleigh, NC, USA, 28–30 March, 1977, volume 12(3) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, March 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1977:PSA

[1118] ACM, editor. Proceedings of the Strathclyde ALGOL 68 Conference. Glasgow, UK; 29–31 March, 1977, volume 12(6) of ACM SIGPLAN Notices. ACM Press, New York, NY, USA, June 1977. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1978:ASH

[1119] ACM, editor. ACM SIGPLAN History of Programming Languages Conference. Los Angeles, CA, USA, 1–3 June, 1978, volume 13(8) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, August 1978. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1979:PSS

[1120] ACM, editor. Proceedings of the SIG-PLAN Symposium on Compiler Construction. Denver, CO, USA, 6–10 August, 1979, volume 14(8) of ACM SIG-PLAN Notices. ACM Press, New York, NY, USA, August 1979. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). LCCN QA76.7.S53 1992.